A Complete Bibliography of Publications in Communications of the ACM : 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

01 May 2024 Version 1.49

Title word cross-reference

$10 [327]. 3 [1076, 363]. 5 [88]. _{2} [1072]. L [1094].

-intensive [175]. -Space [1094].

101 [816]. 12 [110]. 19 [871, 461, 123, 294, 474, 487, 621, 319, 622, 246, 98].

2.0 [1242, 570]. 2024 [1231]. 2032 [712]. 21st [730].

4.0 [694].

5.0 [1071].

6G [693, 808].

= [579].

A* [11]. AAC [1203]. Abilities [1093]. ability [825]. Able [1083]. Above [53]. absolutely [580]. Abstractions [646, 985, 945]. absurdity [593]. abuse [120]. Abusing [1257]. Academia [1147]. academic [390, 839, 873]. Accelerating [1040, 658, 781]. Accelerator [592]. Accelerator-level [592]. accelerators [157]. Acceptance [1192]. access [273, 120, 212, 670]. accessed [604]. Accessibility [1144, 1092, 758, 468, 784, 133, 532]. Accountability [1022, 847]. Accuracy [1022, 576]. ACE [919]. Achieve [844, 21].

Achievement [1194]. Achieving [678, 966, 1006, 1055]. ACM [712, 103, 589, 716, 74, 973, 1187, 581, 630, 713, 783, 925, 1274, 738, 96, 99, 402, 653, 714, 988, 1035, 482, 949]. ACM’s [717]. Across [45, 84, 617, 803]. Act [1030]. action [661, 757]. Actionable [923]. Active [932]. Activities [1056]. Actually [1153]. adapting [96]. adaptive [200]. Add [932]. addicted [170]. Adding [953, 64]. additive [354]. Address [962, 838, 151]. Addressing [614, 744]. adoption [131]. ads [151]. Advance [1208, 903]. Advanced [1170, 912]. Advances [796, 87, 944, 292]. Advancing [825, 785, 40, 158]. Advantage [1006]. Adventures [63]. adversarial [263, 522]. adversary [262]. Advertising [1257, 303]. advice [352]. advisors [841]. Afghanistan [1046]. Afraid [1185]. After [949]. Aftermath [947, 398]. Again [1143, 395, 673]. Against [1226, 525, 635, 912, 268, 603]. Age [1149, 1279, 1242, 1102, 192, 668]. agency [402]. Agent [500, 652, 736, 803]. Agile [1154, 221]. Agility [977]. aging [468]. agriculture [869, 599]. ahead [587, 615]. AI [1172, 381, 882, 467, 962, 1061, 1243, 1227, 1220, 272, 1253, 277, 696, 1122, 1189, 1149, 243, 1017, 1221, 1212, 536, 888, 1040, 1262, 1067, 1185, 191, 679, 971, 1076, 574, 442, 102, 1213, 697, 900, 1089, 934, 1259, 324, 806, 1054, 931, 1257, 1112, 809, 1084, 1108, 803, 1057, 1055, 299, 1118, 969, 1242, 1011, 699, 430, 658, 829, 133, 715, 700, 514, 352, 950, 889, 695, 249, 1009, 923, 284, 873, 904, 915, 1258, 150, 1145, 1043, 1239, 1165, 280, 832, 1256, 486, 1053, 489, 684, 1287, 82, 963, 607, 542, 564, 598, 83, 565]. AI [569, 1182]. AI-based [243, 352, 598]. AI-CHD [598]. AI-enabled [381]. AI-Powered [1182]. Aid [1015]. ailing [887]. ails [584]. aim [294, 25]. air [676]. AIs [1045]. alerts [546]. algebra [185, 184]. Algebraic [1226]. Algorithm [95, 140, 500, 433, 141]. Algorithmic [922, 529, 1208, 1228, 659, 1150, 1022, 956, 1284, 1114, 614]. Algorithmics [1051]. Algorithms [1179, 1098, 1013, 1015, 1052, 1050, 1176, 771, 44, 1086, 1029, 646, 424]. alignment [201]. alive [43, 190]. Alleged [1023]. Allen [213]. Allocation [1290, 1270, 1271]. Almost [1179]. Almost-Linear-Time [1179]. Along [972, 475]. AlphaFold [991]. AlphaZero [644]. already [516]. also [344]. alternate [900]. Always [742, 466]. Always-on [466]. Amazon [243]. Ambition [903]. ambitions [389]. Ambitious [1141]. amélioré [328]. Amendment [1190]. America [242, 248, 253, 250, 243, 255, 256, 251, 244, 257, 245, 247, 246, 249, 252, 254]. American [251]. amid [576]. among [893]. Analogy [1190]. Analysis [89, 938, 1066, 1050, 1176, 980, 491, 642, 279, 206, 877, 666, 624, 620, 643]. Analytics [1044, 1170, 1175, 1060, 781, 669]. Analyzing [225, 226]. Angel [401]. Anger [1292]. animating [125]. annual [426]. Anomaly [91]. Answer [94]. Answers [1255, 762]. Anthropomorphism [1223]. anti [329]. anti-patterns [329]. antifraud [25]. Antikythera [93]. antitrust [108]. any [50]. Anymore [73]. anyone [28]. anytime [701]. Anywhere [1175, 701, 370]. Aotearoa [1058]. apes [810]. app [855, 613]. Apple [613, 768]. Application [465, 919]. Application-Centric [919]. Applications [1023, 336, 12, 445, 164, 265, 332]. Applied [832]. Applying [852, 569, 674]. Approach [1057, 868, 171, 116, 246, 666, 483, 176, 158]. Arab [380, 384, 390, 381, 393, 394, 376, 386, 379, 391, 378, 377, 383, 387, 385, 389, 380, 392, 382, 388]. ArabHCI [385]. Arabs [394]. architectural [675]. architecture [305, 876, 310, 449]. Architectures [905, 674]. Archive [1164]. Argentina [255]. arise [316]. Arrival [940]. Art [1287, 898, 36, 111, 507, 504]. Article [1231]. Articles [942, 984, 962, 983, 964, 1026, 1027, 1028, 941, 1050, 982, 1049, 943, 944, 963, 1029]. Artificial [731, 617, 1020, 1045, 721, 1077, 635, 970, 71, 974, 1217, 1104, 47, 886, 610, 705, 233, 775, 380, 558, 154]. Asia [89, 85, 78, 1070]. Asiacrypt [87]. Asian [1069]. ASIC [163, 164]. aspirations [712]. Assessing [840, 550, 321, 524]. assessment [566]. assistance [858]. Assistants [75]. Assisted [1075]. Assistive [1068, 860]. association [763]. Assumptions [1251]. Atlas [812]. attackers [751]. Attacks [982, 211, 162, 603, 445]. Attaining [142]. auctions [496, 497]. Auditing [922, 923]. augment [412]. augmentation [454]. augmented [587]. AuraRing [845]. Australia [86, 83]. authentication [279]. authorship [150, 357]. autism [648, 649, 128]. Autocorrect [1211]. Autoethnography [1093]. Automata [420]. automate [217]. Automated [1134, 75]. automatic [898]. automatically [265]. Automating [101, 671, 38]. automation [575, 101, 690, 7, 744]. Autonomous [1111, 383, 194, 982, 639]. Avalanches [195]. Aversion [435]. aviation [194]. avoided [662]. avoiding [759]. Awakening [364]. awards [426, 630]. Aware [1067, 264]. Awareness [1263]. awry [112]. AZERTY [328].

B.S. [863]. Back [511, 866, 560, 859, 232]. Backbone [1090, 1091]. Backend [1001]. Backup [1275]. backward [851]. bad [8]. bail [803]. bail-bond [803]. baked [314]. bakery [819]. Balance [1187, 312]. Balanced [1284]. Balancing [1030, 734, 735]. balloon [664]. balloons [664]. balls [210]. Bank [960]. bar [144]. Barbershop [1125]. barrier [45]. base [341]. baseboard [32]. Based [1069, 1270, 1271, 1269, 789, 921, 243, 876, 621, 311, 352, 449, 493, 598]. basics [429]. Basing [1008]. basket [815]. Battery [416, 456]. battery-free [456]. Batteryless [1247]. Battle [1235]. BBC [56]. Be [1017, 1241, 1211, 1083, 1284, 582, 414, 174, 344, 790, 240, 300, 115]. beacons [244]. Bear [1124]. Beating [711]. Beautiful [945]. beauty [553]. become [419]. becomes [595]. Becoming [200, 638]. Begin [1255]. Begins [905]. behavior [618, 653]. Behold [425]. Being [1146, 585]. belief [881]. believability [434]. Bell [327]. belongs [179]. below [53]. benchmark [898]. benchmarking [899]. beneath [897]. Benefit [1217, 569]. Best [969, 582, 367, 155, 413]. Better [479, 1013, 495, 582, 405]. between [1187, 730, 1007, 1114]. Beyond [1130, 1190, 1271, 1050, 113, 1110, 461, 1131]. Bias [922, 1188, 1260, 1029, 444, 707]. Biased [923]. Biases [489, 708, 344]. bibliometric [116]. Big [390, 393, 872, 394, 703, 253, 871, 860, 869, 864, 572, 866, 386, 255, 256, 391, 251, 571, 387, 1054, 1056, 257, 1055, 699, 874, 700, 701, 389, 873, 698, 392, 1053, 252, 861, 863, 388, 1193, 569, 570, 254, 568, 702, 807, 596, 169, 21, 483, 517, 608, 844]. biggest [594]. bimodal [19]. Binary [89, 1081]. Bioinformatics [1050]. biological [333]. Biologist [903]. biology [731]. Biomedical [391, 1119]. BioScript [334]. birthdays [341]. bit [56]. Black [938, 371, 790]. BLeak [265, 264]. blind [784]. Bloated [1027]. block [767]. Blockchain [572, 1023, 1063, 1064, 13, 160, 820]. blockchains [320]. BLOG [366, 1097, 1275, 1162, 678, 850, 1254, 1233, 1118, 1141, 3, 232, 550, 1038, 1207, 1183]. Blood [1181, 499, 849]. blooming [389]. Body [1069]. Boeing [295]. Bolster [1121]. bond [803]. book [105]. Boom [1213, 809]. boosted [447]. Boosting [1159]. boot [52]. borders [617]. bot [224]. Both [1262]. Bother [1207]. Bots [1017]. Bottleneck [944]. Bouncing [210]. Bounded [95, 422].

Bounds [1226, 228]. box [790]. Brain [983, 766]. Brand [857]. brave [641]. Brazil [252]. Breakfast [1154]. Breaking [1021]. Breakthrough [1098]. Bridges [1060]. Bridging [1227, 394, 486]. Bringing [866, 337, 232]. broad [336, 697]. broadband [670]. broaden [273]. Broadening [273, 371, 171, 532]. Broken [172]. Brooks [908]. Buffer [1027]. Buffer-Bloated [1027]. bug [491]. bug-free [491]. build [155, 819, 502, 562]. Building [1061, 378, 377, 765, 889, 936, 1288, 861, 375, 183]. Built [989, 543]. built-in [543]. Burrows [756]. Bursting [664]. Buses [986]. Business [1147, 990, 976, 591, 320, 8, 536, 483, 151, 238, 769]. Bust [1213]. Buy [1136]. BWT [755]. Byte [947, 1229, 967, 1116, 1034, 1180, 1252, 1094, 1160, 1292, 1009, 987, 924, 1039, 1139, 1272, 1205, 287, 580, 207, 364, 500, 652, 736, 880, 278, 142, 39, 186, 266, 337, 605, 525, 450, 825, 900, 759, 676, 803, 425, 20, 104, 229, 61, 165, 401, 547, 312, 475, 711, 846, 782, 629].

C [491, 33, 1270]. CACM [366, 1097, 1275, 1162, 678, 850, 1254, 1233, 760, 1118, 1141, 3, 232, 550, 1038, 1207, 1183]. CAD [538]. Cake [95]. California [347]. Call [154, 661, 514]. CALM [202]. Came [1140, 427]. Cameras [1177]. Camille [585]. Can [1090, 1017, 1241, 888, 43, 679, 1116, 931, 344, 979, 1009, 1277, 1107, 1046, 799, 509, 351, 161, 97, 300]. canonization [422]. can’t [718, 115]. Capable [1078]. CAPE [321]. caps [789]. Capturing [78]. car [307]. Carbon [1077, 313, 243, 51, 41]. Card [846]. care [748]. Career [290, 848, 1036, 452, 989, 404, 549, 1232, 715, 340, 950, 1096, 230, 654, 608, 502, 805, 1182, 903]. Cargo [1108]. Caribbean [254]. Carolina [180]. Cars [982, 680]. Case [1155, 651, 378, 377, 240, 252]. Cases [1004]. cast [302]. Catalyst [968]. catastrophic [618]. catch [466, 176]. Catching [407]. causal [60]. Center [1189]. Centered [964, 1110, 666]. centers [873]. Central [960, 254]. Centric [1089, 919, 34]. centricity [173]. century [730]. CEO [581, 630]. Cerf [22, 208, 123, 166, 267, 476, 526, 314, 583, 427, 365, 739, 631, 677, 827, 881, 1010, 1117, 1161, 1073, 926, 968, 1253, 1206, 784]. Certificates [1134]. Certification [1111, 1123]. certifications [541]. Ch [425]. Chain [981, 682]. chains [457]. Challenge [1000, 1058, 144, 956, 366, 826, 173, 664, 522, 768, 726]. Challenges [1018, 1064, 1145, 1043, 754, 599, 282, 266]. Challenging [1009, 778]. Chances [1048]. Change [1160]. Changing [806, 654, 384, 540]. channel [211]. characterizing [670]. Charles [458]. chase [803]. Chasm [901]. Chatbots [1162, 1041, 494]. ChatGPT [1166, 1009, 1277]. CHD [598]. cheap [774]. cheating [882]. check [152]. checking [241]. chemistry [334]. chess [268, 644]. Chief [737]. Chien [1000]. Chile [253]. chimes [104]. China [572, 563, 571, 560, 561, 566, 564, 565, 569, 570, 562, 567, 568]. Chinese [1277]. chip [334]. chiplet [448, 449]. chiplet-based [449]. chips [168, 528]. Choice [957, 548]. chronic [701]. Circuits [1226]. cities [509, 245]. citizen [30]. citizens [153]. citizenship [724]. city [384]. Civilizations [1168]. Clarke [345]. Clash [1168]. classes [411]. classification [843, 842, 447]. Clear [1037]. Clearing [1081]. Click [1095]. Click-Through [1095]. client [794]. Climate [1160, 869, 145]. clinician [352]. clinicians [701]. closer [235]. CLOSET [1059]. Closing [1101]. Cloud [1196, 1224, 1195, 919, 692, 34, 313, 577, 441, 419]. clouds [163, 415, 164, 493]. Cloudy [943]. Clue [1267]. CMD [148]. Co [1249, 1068, 442, 1072]. co-design [442]. Co-Designing [1068]. Co-Developing [1249]. Coalition [294]. Cobblers [914].

Cocktails [924]. CoCoPIE [442]. Code [1065, 1198, 1003, 1133, 510, 616, 668]. Codes [926, 410, 309, 735]. Coding [971, 42, 260, 411, 228]. coffee [172]. Cognitive [708, 91, 707]. cohort [23]. COI [1059]. coin [453]. cold [508]. collaborating [841]. collaboration [1, 192]. Collaborative [919]. Collusion [439, 120, 119]. Color [784, 64, 171]. combat [888]. Combination [1132]. Combining [1271, 980]. Come [67]. comes [317]. comfortable [499]. Coming [1102, 464]. Commercial [923]. Commercialization [88]. Commit [134]. Committee [949]. Common [584, 121, 233]. Commons [987, 531]. Communicate [135]. Communicating [581]. Communication [1137, 1204, 490, 476, 527, 540]. Communications [910, 1040, 100, 338, 737, 685]. Communicator [1146]. Communities [1243, 404]. Community [1281, 1074, 606, 391, 30, 738, 46, 760, 152, 517, 726, 724]. compact [250]. companies [481]. company [617, 834]. Comparing [1162]. compatibility [535]. competencies [669]. competent [894]. Competition [1126, 807]. competitions [550]. Competitive [535]. compilation [442]. compilers [646, 33]. Complaint [949]. complex [515]. Complexity [1003, 1007, 342]. Comprehension [1003]. comprehensive [604]. compression [442, 755]. compression-compilation [442]. Computation [757, 777, 813, 308, 306, 218, 814]. Computational [590, 930, 1087, 793, 328, 418, 897, 754, 143]. Computations [1232]. Compute [998, 437, 794]. Computer [983, 1020, 1211, 1146, 1075, 1018, 914, 999, 1288, 972, 993, 1215, 1287, 903, 273, 580, 500, 181, 848, 762, 321, 23, 9, 30, 687, 257, 325, 439, 218, 116, 765, 19, 180, 594, 841, 655, 541, 110, 727, 357, 488, 636, 470, 724, 562]. Computer-Assisted [1075]. Computer-Related [1018]. Computers [111, 24, 209, 161, 525, 883, 675, 29, 268, 138, 355]. Computing [204, 661, 860, 461, 869, 272, 920, 288, 826, 901, 910, 1221, 1212, 1036, 932, 348, 934, 1113, 1125, 46, 1058, 1055, 989, 1276, 1232, 950, 152, 1096, 1195, 203, 1079, 556, 145, 1140, 1181, 1230, 1078, 1062, 1110, 1182, 903, 703, 248, 15, 856, 371, 583, 121, 582, 606, 179, 290, 848, 840, 256, 391, 106, 605, 452, 27, 173, 893, 444, 404, 549, 210, 715, 340, 409, 17, 230, 441, 654, 419, 205, 608, 502, 805, 463, 740, 688, 533, 540, 40, 98, 548, 289, 847, 723, 254]. computing-research [548, 289]. Concepts [1154, 470]. Concurrent [921, 17]. Conference [1151, 1035]. Conferences [1216, 74, 51, 41, 232, 548]. conferencing [146]. Confidential [1197, 1195, 441]. configuration [733]. conflating [468]. Conflict [1200]. conflicts [217]. Confusion [1081]. congenital [598]. conjecture [756]. Conjoined [1020]. connected [533]. Connecting [394]. Connections [1012, 1007, 337]. Connectivity [1010, 1091]. conquer [534]. Consensus [1063]. Considerations [1078]. Considering [967]. Consistency [917, 202]. Constant [309]. constraint [240]. Constraint-Aware [1067]. consultant [805]. Consumers [1240, 153]. contact [207, 236, 621]. Content [297]. contest [828]. context [709, 886, 710, 727]. Contextual [138]. Contextualized [247]. Contracting [974]. Contracts [1095, 226, 225]. contribute [463]. Contributed [942, 962, 983, 964, 1026, 1027, 941, 982, 1049, 943, 963]. contributor [594]. Control [905, 559]. controller [32]. Controlling [1118]. Convenience [953]. Convergence [1192, 875]. conversation [667, 353, 470]. Conversational [1129]. conversions [504]. Converting [611].

Cookie [770]. cooking [602, 601]. cooling [745]. Cooperation [1084, 987, 1053]. Coordination [90]. cop [408]. Copilot [1025, 1246]. Copyright [237, 660, 1239, 554]. cordless [527]. core [311]. coronavirus [192]. corporate [494, 653]. correct [161]. Correctness [1249, 1252, 862]. Corrigendum [1231]. Cost [1179, 1048, 308, 588, 616, 432, 598]. cost-effective [598]. Costs [90]. could [678]. counterpoint [813, 814]. Counting [1115, 1114, 385]. countries [687]. Courses [932, 865]. COVID [871, 461, 123, 294, 474, 487, 621, 319, 622, 246, 269, 98]. COVID-19 [871, 461, 123, 294, 474, 487, 621, 319, 622, 246, 98]. Cracks [1]. Crash [917]. crazy [132]. Creating [1197, 839]. creation [595]. Creativity [1070, 931, 1009, 1096, 351, 178]. crime [500]. crime-solving [500]. crisis [181, 828, 662]. Critical [1281, 916, 1123, 1228, 171, 239, 761, 114]. criticality [724]. Crossing [787]. crossroads [260]. Crowd [1177]. crowdfarms [797]. Crowdsensing [570]. Crowdsourcing [54, 1107]. Crowdworkers [1285]. Crypto [906, 1097]. Crypto-NFT [1097]. cryptocurrencies [223]. Cryptographic [977, 1032]. Cryptography [1008, 929, 1157, 615, 332]. CS [167, 366, 411, 101, 866, 938, 102, 291, 405, 678, 189, 412, 266, 1259, 239, 62]. CSIRO [1057]. CSRB [892]. culprits [300]. Cult [1108]. Cultural [78]. cultures [556]. curation [595]. Currencies [960]. Currency [1083]. Current [1167, 282, 398]. Curricula [1215, 412]. curtail [129]. custodes [827]. custodiet [827]. Cutting [95]. CXL’s [1174]. Cyber [978, 358, 389, 1064, 92, 959, 281, 728]. Cyber-Physical [1064, 728]. Cyberattack [1183]. cyberinfrastructure [792]. Cybersecurity [73, 296, 983, 1121, 323, 1242, 874, 79, 1048, 80, 723, 454, 831, 186]. cycle [158].

D [1076, 363]. damage [6]. Dana [791]. Dangers [1275]. Dark [25, 198, 281, 1099]. DARPA [664]. Data [1059, 1275, 833, 1188, 1155, 946, 379, 1175, 141, 1254, 259, 640, 1089, 173, 1105, 928, 914, 1060, 1049, 70, 944, 69, 1217, 1284, 990, 591, 253, 360, 250, 575, 651, 175, 303, 696, 671, 492, 462, 126, 278, 292, 574, 478, 850, 452, 212, 612, 793, 877, 748, 691, 730, 889, 873, 554, 203, 689, 891, 158, 753, 472, 238, 493, 388, 669, 604, 799, 1066, 1192]. Data-Centric [1089]. Data-centricity [173]. Data-Driven [1059, 141, 575, 730]. Data-Enabled [70]. Data-Poisoning [946]. Data61 [1057]. Database [393, 798, 185, 466]. Databases [1248]. datacenter [163, 164]. datasets [807, 600, 670]. Datasheets [600]. date [716]. Daughter [950]. dawn [797]. day [416, 307]. DDT [121]. Dead [67]. Dear [1214]. debloating [733]. Debugging [220, 176, 265]. decade [224]. decades [393, 678]. Deceiving [107, 430]. Decentraland [1023]. Decentralization [1023, 320]. Decentralized [1192, 818]. deception [751]. decide [433]. Deciphering [308]. Decision [946, 396]. declaration [217]. Declarative [619]. decline [355]. Deconstructing [819]. Deep [467, 1244, 1055, 368, 1131, 1038, 1029, 360, 867, 821, 497, 6, 292, 616, 758, 448, 159, 842, 743, 449, 361, 1255]. Deep-Learning [1244, 1029, 449]. deeper [743]. Deepfake [1042]. deepfakes [6, 348]. default [326]. defenses [417]. definability [422]. defined [602]. defining [248]. Degrades [1243]. degree [863]. degrees [356]. Deliver [1015, 320]. Delivering [622]. delivery [494, 493]. Democracy [976, 153, 804]. Democratizing [920, 864]. demon [432]. Demonstrate [931]. Demystifying [156]. Departments [101, 100, 103, 938, 144, 97, 40, 143, 98, 2, 187, 402, 548, 339, 451, 289, 501, 653, 847, 763, 804, 607, 714, 988, 902, 1035, 1095, 1140, 948, 1181, 1273, 1230]. Dependability [15]. dependable [256]. Dependency [1067]. Dependency/Constraint [1067]. Dependency/Constraint-Aware [1067]. Deploy [69]. Deploying [818]. Depth [1226, 438]. Deriving [472, 471]. Describe [1225]. Design [1286, 952, 965, 1289, 1079, 1203, 674, 328, 442, 141, 9, 675, 545, 629, 533]. design-space [674]. Designing [958, 1064, 1129, 643, 1068]. destroy [607]. detachment [746]. Detecting [85, 124, 1266, 116]. Detecting/preventing [124]. Detection [1059, 224, 709, 539, 710, 566, 264]. Determines [1262]. Deterministic [986]. develop [31]. Developer [958, 464, 440, 666]. Developers [1199, 673]. Developing [1249, 83]. Development [1004, 1061, 1154, 708, 109, 537, 388]. DevEX [1153]. Device [1264, 1110, 499]. Device-Centered [1110]. Devices [1183, 1068, 551, 442, 363, 854, 844]. DevOps [1282, 221]. diagnosis [247]. diagnostics [795]. dialectal [394]. DIAMETRICS [899]. Diary [1214]. DIAS [1066]. didn’t [9]. differ [896]. Different [1162, 1273, 207, 396]. Differential [326, 462, 332]. difficult [673]. Digging [596]. Digital [960, 84, 1021, 739, 599, 236, 705, 1067, 304, 178, 1119, 169, 1101, 1083, 1164, 1191, 81, 252, 863, 864, 201, 208, 685, 512, 663, 316, 469, 622, 722, 502, 561]. Digitization [1275, 351]. digitized [174]. Dijkstra [341]. Dilemma [1117, 1104, 545]. Dilemmas [47]. dimension [142]. dimensional [472]. Dimensions [1147, 976, 591, 320, 536, 483, 151, 238, 769]. diplomacy [208]. direction [122]. Directions [35]. directives [618]. Dirty [915]. disability [722]. disassembler [485]. Disaster [1015, 689]. disciplined [158]. Disclose [949]. Disclosure [949, 223]. discontinuum [14]. discourse [641]. Discovery [1087, 970, 952]. discreetly [269]. discrimination [614]. disease [598]. diseases [566]. Disentangling [1006]. Disinformation [1242]. disorders [649]. disputes [315]. Disputing [341]. disruption [682]. disruptive [537]. Distance [1176, 146]. Distilling [1100]. Distinct [1115]. Distributed [916, 1031, 220, 202, 819, 60, 735]. distribution [601]. Diversity [1263, 890, 870]. Divide [534]. Divided [901, 288, 606, 740]. Divides [1101]. Division [94, 655]. Do [994, 1200, 1230, 672, 896, 631, 121, 221, 6, 111, 348, 618]. doctoral [841]. Does [523, 303, 940, 1032, 484, 365, 8]. dogged [491]. Doha [384]. Domain [920, 157, 159]. Domain-Specific [920, 157, 159]. domestic [187, 488]. done [509]. don’t [360]. Door [1051]. doppelgängers [25]. Down [992, 325, 429]. DP [332]. DP-cryptography [332]. Draw [1065]. Draws [1076]. drive [97]. Driven [1283, 1059, 1004, 575, 141, 58, 666, 730, 473]. Drives [953, 1153]. Driving [313, 828, 214, 383, 680, 809]. Drug [952]. dumb [325]. dupe [25]. duplicates [634]. durability [22]. during [461, 201, 91]. Dynamic [1291, 1111, 980, 60]. Dynamics [444, 764].

e-infrastructure [692]. E-Waste

[1235, 149]. Each [1230, 268]. ear [499]. ear-worn [499]. early [390]. earth [408, 1255, 1066]. East [78]. Easy [1225, 202]. eBP [499, 498]. Echoes [1152]. Economic [591, 320, 536, 1147, 483, 151, 238, 976, 769]. Economics [1230, 160, 143]. Economy [1088, 81, 101, 365, 889, 561]. Ecosystem [1189, 80, 321, 251, 389, 196, 382]. ecosystems [183]. Edge [919, 15, 191, 310, 311, 284]. Edge-Cloud [919]. Edit [1176]. Editing [57]. editor [477, 41, 145, 188, 122, 231, 584, 315, 849, 655, 740, 1, 121, 338, 288, 313, 582, 426, 826, 712, 737, 606, 901, 1072, 783, 760, 21, 403, 1231, 453, 503, 1037, 927]. Editor-in-Chief [737]. Editorial [762, 713, 738, 1190, 761]. Edmund [345]. Education [856, 484, 411, 789, 1281, 1221, 1155, 932, 321, 996, 1259, 239, 173, 46, 62, 557, 1079, 110, 636, 724, 273, 366, 864, 866, 762, 291, 609, 478, 266, 728, 741, 607]. Educational [356, 1060, 512]. Effective [980, 598]. effectively [189]. effectiveness [823, 645, 109]. effects [833, 668]. Efficiency [1159, 953, 978, 1284, 98]. Efficient [1055, 1276, 980, 802, 879, 801]. Effort [1054, 657]. eggs [815]. Eight [983]. Elderly [979]. Election [1274, 434, 99, 717]. electromagnetic [845]. electronic [621, 854]. elegant [471]. Elements [1115, 698]. Elephant [933]. Elevating [1195]. ElGamal [1033]. eliminating [149]. Embedded [1097, 985, 285]. Embracing [761]. emergency [546]. emergent [854]. emerging [332]. Emirates [380]. Emissions [1224]. emoji [706]. emphasizing [454]. empirical [181]. Empower [1240]. empowered [608]. Empowering [82]. ’empty [676]. emulation [660]. Enabled [70, 381, 452, 404]. Enablers [1191]. Enabling [340, 284, 442]. Enclaves [415]. encoding [446]. encouraging [550]. encrypted [843]. Encryption [1005]. End [1037, 913, 186, 40]. Energy [1055, 1224, 698, 851, 676, 879, 340]. Energy-Efficient [1055, 879]. Enforcement [83]. engage [487]. engagement [152]. engine [609]. engineer [443, 594]. Engineering [935, 1057, 703, 866, 555, 413, 453, 550]. Engines [1260, 899]. Enhance [1284, 627]. Enhancing [1191]. Enigmatic [94]. Enough [1022]. enrollment [789, 356]. enrollments [789]. entanglement [578]. entering [694]. Enterprise [1011, 876, 12]. Entity [182, 183]. Entrepreneurship [382]. entropy [651]. envelope [511]. Environment [1066, 729]. environments [274]. Envy [95]. Envy-Free [95]. Epic [613]. Epigenomics [1184]. equality [97]. equations [471, 472]. Equity [1263, 890, 321]. Era [1221, 538, 706, 842, 793, 51]. erasure [432]. Error [905, 1022]. errors [308]. esports [809]. Essential [941, 31]. Estimating [243]. estimation [651, 650]. Ethical [958, 934, 1078, 197, 556]. Ethicist [915]. Ethics [1283, 962, 1212, 630, 10, 934, 1084, 997, 1078, 661, 461, 272, 410, 348, 514, 152, 556, 486, 653, 307]. ethnicity [668]. Europe [692, 703, 686, 696, 1155, 690, 693, 697, 687, 694, 699, 700, 691, 701, 695, 698, 689, 688, 702]. European [692, 687, 699, 700, 691, 483, 558, 702]. Evaluate [957]. Evaluating [778]. evaluation [58]. Even [618]. event [685]. Ever [951, 384, 580]. ever-changing [384]. every [594]. Everyday [1137, 1138]. everyone [301]. Everything [1010, 395, 39]. Everywhere [1175]. Evidence [1060, 19, 347]. Evil [1191]. evolution [895, 1026]. Evolve [1254]. Evolves [969]. Examining [180]. Example [1263]. exascale [387]. Excellence [1245]. Excessive [300]. exchange [891]. Exclusivity [782]. execute [333]. execution [162]. Executives [1171]. Exercise [1009]. EXP [1008]. Expanded [962]. expander [309]. expanding [273]. Expectations [1167]. expected [639]. expensive [51]. experienced [662]. Experiences [1093]. experiment [523]. experiments [754, 524]. Expert [1161, 200]. experts [672, 669]. Explain [1260]. Explainable [684, 867]. Explicative [557]. explode [625]. Exploiting [1134, 162]. exploits [307]. Exploration [91, 674, 649, 730]. Explorations [728]. Exploring [762, 605, 707]. Expressing [966]. Expressive [781, 1204]. external [228]. Extracting [941]. extraction [898]. extreme [567]. extreme-scale [567]. Eye [1182]. Eyelid [628, 627].

Fabrication [1068]. Fabs [1278]. face [789, 383, 659]. Facebook [303, 632]. Faces [918]. facial [344, 665]. Fact [343, 536]. Fact-finding [343]. factor [279]. Factual [1206]. Faculty [938, 979, 893, 514]. fails [39]. failure [295, 176]. Fair [94, 904, 396, 459, 914]. fairly [799]. Fairness [1220, 1200, 800, 117, 396, 786]. Fake [85, 1136, 1135, 870, 119, 709, 710, 241]. Fakes [1131, 407]. false [151]. Family [74]. famous [664]. FANG [710]. fantasy [536]. Far [1173]. Fast [1268, 1142]. Faster [1013]. FATE [1074]. Fault [1276, 309]. Fault-Tolerant [1276]. FDO [1264]. Feature [1062, 100]. Features [1154]. February [1231]. Federated [704, 277, 691]. federation [692]. Feedback [20]. feel [167]. female [356]. Fenix [692]. fever [322]. few [664]. FHIR [891]. Field [1137, 331]. fields [625, 626]. Fifty [623]. Fight [1120, 635]. figures [636]. Filtering [553]. filters [400]. Finances [973, 1187, 103, 589]. financial [240]. Finding [900, 504, 1096, 1007, 343, 576]. Findings [949]. Finds [930]. Fine [1287]. finger [845]. fire [749]. First [1197, 1288, 1140, 338, 685, 207, 850, 1172, 1190]. first-contact [207]. Fit [1292]. Fits [1258]. Five [1171, 737, 258, 385]. Fix [92, 535, 113]. fix-it [113]. fixed [344]. Fixing [480]. Flatter [528]. Flexible [469, 1174]. Flight [1025]. Flood [206, 205]. Flood-risk [206]. Flow [1179, 1075, 1178]. Fluid [1075]. Focus [1105]. Footprint [1077]. Forecast [943]. forget [679]. forgetting [764]. form [446]. Formal [455, 868, 46]. Formalizing [786, 519]. Formats [1151]. Forward [1234, 884, 900]. fosters [564]. foundational [253]. FoundationDB [1030, 1031]. foundations [450]. Founded [1251]. Four [712, 932, 49, 887]. Fourier [1142]. foxes [20]. FPGA [802]. FPGAs [794, 219]. Fragility [1045]. Fragmented [1267]. Frama [491]. Frama-C [491]. Framework [1248, 1060, 1217, 1129, 321, 598]. frameworks [465]. Fran [213]. Frederick [908]. Free [95, 1127, 976, 491, 456]. Free-Speech [1127]. frequent [499]. friction [891]. friend [262]. frontiers [117]. fueled [290]. Fugaku [293]. fugitives [803]. Fully [1005]. Fun [1002, 549]. function [472]. Functional [966]. Functions [1007]. fungible [507]. Fusing [1070]. Fuss [1158]. Future [1216, 1063, 1190, 71, 381, 871, 296, 526, 288, 613, 282, 171, 507, 776, 457, 219, 198, 215, 285, 695, 517, 740, 665, 723, 769, 568]. Futures [663, 536, 858]. Fuzzer [1159]. Fuzzing [1158, 36, 1002].

G [88]. Gaining [1217, 559]. galaxy [803]. game [540]. game-changing [540]. gamer [580]. games [578, 512]. Gamification [1147]. gap [394, 116, 249]. gas [379, 226]. GDPR [329]. geek [18]. GenAI [1072]. Gender [357, 893, 116, 444, 668, 97]. gene [18]. General [1274, 99, 717, 355]. generalization [361]. generalized [418]. generate [882]. Generated [1266]. Generating [1134]. generation [751]. Generative [1243, 1122, 1189, 1017, 1221, 263, 1259, 1009, 1258, 1145, 1043, 1239]. Geometric [292]. Geschke [458]. gestures [628, 627]. get [766]. Getting [683, 911, 429]. giants [535]. Giga$ [1072]. GigaTons [1072]. GitHub [642, 1246]. Glimpse [1135]. Global [890, 27, 682, 533, 508, 606, 96, 389, 392]. globe [5]. Go [851, 1128, 673, 729]. goes [316, 112]. Gold [1205]. gone [187]. Good [582, 988, 8, 13]. Goods [1039]. Google [220, 1001]. Goose [803]. Governance [1082, 663]. Governments [1256]. GPA [789]. GPA-based [789]. GPS [878, 1186, 879]. GPUs [1197]. Grace [549]. grad [231]. grades [19]. graduate [187, 488]. graduates [894]. Grand [1000, 1058, 366, 826]. Graph [1268, 261, 710, 517, 421]. graphs [422, 359, 517, 567]. Green [1055, 280, 207]. grids [340]. Groundbreaking [1178]. Group [1208, 1093, 657]. growing [391, 616]. Growth [1187]. guaranteeing [519]. guesswork [650]. Guide [1245, 372]. Guiding [31, 701]. guilds [55].

hack [366, 36]. hackcraft [580]. hackers [373]. Half [314]. Half-baked [314]. Halfway [1169]. Handle [911]. Handles [911]. Hands [915, 1181, 820]. handwriting [832]. Happened [951]. happiness [594]. harbors [237]. Hard [1225, 484]. Hardness [1008]. Hardware [1196, 157, 597, 318, 794]. harm [468]. Hartmanis [830]. harvest [44]. having [882]. HCDA [418]. HCI [495, 654]. Heads [1110]. Heads-Up [1110]. health [498, 721]. Healthcare [84, 891, 252]. heap [501]. heart [101, 733, 598]. hearts [172]. heat [601, 432]. heats [745]. Help [1098, 989, 825, 404, 97]. Here [673, 948]. Heritage [78]. heroes [373]. heterogeneous [583]. Hidden [831, 90, 636]. Hiding [1250]. Hierarchy [921, 785]. High [966, 314, 432]. High-Performance [966]. high-resolution [314]. Higher [1155, 609]. highest [144]. Highlights [1291, 1159, 1227, 1179, 922, 1033, 946, 1158, 1290, 966, 1136, 1137, 1251, 1030, 1052, 1270, 1226, 1228, 1271, 985, 986, 1115, 1032, 965, 923, 1225, 1135, 1051, 1178, 1204, 945, 1203, 1138, 1250, 1114, 1031]. him [9]. Hiring [128, 659]. Historic [1098]. Historical [9, 298, 373, 510, 638, 1020, 1172, 1213]. history [325, 219]. hits [4]. holistic [215]. Holograms [552]. home [129]. Homomorphic [1005]. honest [414]. Hoping [969]. Hopper [549]. horizon [552]. horsemen [887]. Hot [384, 381, 692, 248, 250, 867, 868, 1059, 1065, 1061, 696, 862, 243, 690, 379, 378, 377, 383, 693, 1067, 1063, 563, 244, 1071, 697, 1066, 1069, 687, 694, 865, 1070, 245, 385, 1057, 1058, 247, 246, 1060, 691, 695, 249, 1064, 380, 870, 689, 688, 561, 566, 382, 1068, 564, 565, 562, 567, 1062]. Hours [1072]. house [749, 711]. How-To [1193]. HPC [894, 290, 943, 196]. Human [73, 1065, 964, 1229, 539, 1069, 1285, 1223, 666, 1080, 1053, 1284, 30, 519, 299, 659, 699, 730, 114]. Human-Centered [964, 666]. human-computer [30]. human-in-the-loop [699]. Human-Machine [1069]. Human-Robot [1223, 519]. Human-Scale [73]. Human-Subjects [1285]. Humanity [903]. humans [169, 29, 1262]. Humble [1112]. humbly [102]. Hummingbird [879]. hunger [635]. hXDP [802, 801]. Hype [1006]. Hyperscalers [1091]. Hypothesis [91].

ICT4D [1046]. idea [132]. ideal [732]. Ideas [1200, 1107]. identifying [868]. identity [718, 301]. if [618]. II [1145]. IIT [863]. illuminator [723]. Im [396]. image [57]. Images [1233]. Imaging [255]. Immersion [745]. immersive [885]. immortal [298]. Impact [947, 922, 967, 923, 1246, 868, 474, 853, 681, 21, 392, 109, 41]. Impactful [862, 152]. impacts [645]. impaired [860, 271, 569]. impairments [628]. implants [766]. Implementation [1289]. Implementing [417]. Implications [123, 197, 660]. Implicit [1029]. Importance [1275, 521]. impossible [623]. impostor [438]. Improve [1027]. Improving [201, 1244, 1290, 1079, 976, 575, 742]. in-copyright [554]. In-sensor [447]. inch [551]. incidents [220]. included [752]. Inclusion [1263, 1144, 890]. inclusive [649]. Incompleteness [1248]. Inconsistency [1248]. increase [109]. Increasing [7, 1263]. increasingly [843]. incrementally [305]. India [872, 871, 860, 867, 869, 868, 864, 866, 862, 859, 1263, 865, 874, 873, 870, 861, 863]. Indian [873]. Indistinguishability [118, 1251]. Individualized [1083]. Industry [1070, 1054, 379, 750, 232, 240, 694]. inequality [893]. infections [124]. inference [448, 449]. Inferring [575]. Inflection [1087]. influence [501]. Influencers [954]. Informatics [1155, 1056, 698]. Information [1159, 1206, 1207, 725, 204, 137, 161, 769]. Informing [347]. Infrapolitics [1150]. infrastructure [692, 691]. Infrastructures [1090, 469]. Ingredient [1203]. ingredients [855]. Inherent [1220]. initiative [685]. initiatives [392]. Innovating [81]. Innovation [1126, 1238, 1122, 955, 1070, 828]. Innovations [561, 390]. innovative [246]. innovators [195]. insects [362]. insider [417, 281]. Insights [299, 388, 403]. inspired [715]. Installer [1264]. Institute [1056, 253]. institutions [866, 873]. instruction [124]. instruments [304]. Insurance [959]. integer [228]. Integrating [196]. Integration [1290, 1066, 1069]. integrity [647, 439]. Intelligence [1020, 1077, 970, 71, 974, 1217, 1152, 919, 1104, 454, 47, 886, 610, 731, 617, 705, 721, 635, 775, 380, 558, 154]. Intelligent [858, 865]. intensive [175]. Intent [876]. Intent-based [876]. Intentional [1036]. inter [757]. Interaction [1137, 1223, 627, 30, 519]. interactive [566]. interdisciplinary [253, 478, 564]. Interest [89, 217]. Interface [983]. Interfaces [1129, 459]. internet [480, 473, 1247, 286, 520, 166, 677, 508, 1240, 492, 535, 282, 474, 796, 694, 5, 1190, 941, 346, 456, 1164, 550, 445]. Internets [49]. Internship [1093]. Interoperability [1253]. Interpretable [795, 16, 247]. intersectional [171]. interventions [356, 645, 863]. Interview [791]. introduction [286]. Intuition [1284]. Invention [1020, 304]. Inversion [1228]. Investigating [1088, 923]. invisible [137, 652, 216]. Involvement [746]. ion [675]. IoT [1014, 493]. ipsos [827]. irrevocably [606]. ISA [1286]. Island [79]. Isolation [1196]. Isomorphism [422, 1208, 261, 421]. Issarny [909]. Issues [316]. item [496, 779]. Iterative [997]. ITIL [221]. itinerary [577].

Jack [431]. Japan [77, 1056]. Java [896]. JavaScript [1289]. Jawbreaker [1265]. jewels [834]. JIZAI [1069]. journal [478, 857]. journalism [178]. Jr. [908]. Jump [1272]. Juris [830]. Just [1105, 1086, 295, 132, 488]. justice [724].

Kabin [322]. KAUST [377]. keel [330]. keep [43]. Keeping [330, 202, 466]. kernel [733, 136]. Key [1203, 1031, 855, 698]. keys [773]. kill [28]. kind [207, 888]. Kindle [115]. kingdom [773]. Knives [1124]. Know [1100, 1219, 594, 438]. Knowing [1219]. Knowledge [359, 1258, 324]. Knowledgeable [563]. Kode [172, 29, 275, 216, 435, 534, 485, 593, 887, 683, 748, 834, 637, 975, 933, 1080, 1169, 1019, 1124, 1214, 130, 322, 372, 790, 1261]. Kolmogorov [1007]. Korea [88, 1054].

lab [378]. labeling [811]. Labels

[1240, 855]. labor [744, 730]. laboratories [255, 334]. laboratories-on-a-chip [334]. Labs [327]. lack [434]. Lamboozling [751]. lamps [139]. Landscape [960, 962, 1104, 840]. Language [1154, 1073, 1103, 1241, 1202, 1094, 776, 1141, 1289, 1222, 1269, 1078, 729, 386, 563, 780, 571, 45, 234, 673]. Language-Based [1269]. Language-Capable [1078]. Languages [67, 1004, 1162, 965, 672, 450]. Large [1073, 1103, 1202, 1094, 1222, 285]. large-scale [285]. Last [287, 580, 947, 207, 364, 500, 652, 736, 880, 1229, 142, 39, 186, 266, 337, 605, 525, 450, 825, 900, 759, 676, 967, 1116, 1034, 1180, 1252, 1094, 1160, 803, 425, 1292, 1009, 20, 104, 229, 61, 165, 401, 547, 312, 475, 711, 846, 782, 629, 987, 924, 1039, 1139, 1272, 1205]. LATAM [249]. Latency [916, 986]. Latin [242, 248, 253, 250, 243, 255, 256, 251, 244, 257, 245, 247, 246, 249, 252, 254]. Launching [100]. Law [994, 508, 410, 1279, 810, 108, 7, 193, 614, 297, 1190, 1101, 722, 911, 83, 43, 412, 640, 788, 317]. Laws [1279, 611]. Lawyers [1185]. layer [821]. Ld [425]. lead [293]. leadership [702]. leak [161, 264]. leaks [265]. Lean [944, 221]. Learn [1199, 1277, 672, 856, 719, 679, 673]. Learning [1044, 1283, 962, 1199, 964, 1244, 1202, 24, 932, 759, 935, 1270, 1055, 1271, 68, 1060, 1224, 936, 1038, 1029, 945, 1148, 1182, 146, 867, 467, 140, 704, 795, 277, 117, 462, 821, 16, 497, 705, 292, 563, 616, 758, 448, 865, 635, 842, 368, 247, 619, 743, 449, 361]. Learning-Based [1270]. Lebanon [382]. lecture [646, 467, 895]. led [502, 38]. LEDs [986]. Legal [1145, 956, 1043, 554, 854, 558]. Legally [150, 237, 346, 459, 554, 768, 660, 854, 1145, 956, 1043, 1239]. lens [48, 143]. Lessons [460, 1141, 487, 194, 863]. Let [414, 433, 325, 815, 1128, 535]. Letter [901, 1231, 1, 121, 338, 288, 313, 582, 426, 826, 712, 737, 606, 1072, 783, 760, 21, 403, 453, 503]. Letters [477, 41, 145, 188, 122, 231, 584, 315, 849, 655, 740, 1037, 927]. level [50, 592]. Leveraging [709, 1212, 126, 1191, 710, 1136]. liability [847]. Libraries [1164]. library [864]. Life [67, 1183, 436, 278, 158, 527]. Life-Saving [1183]. Lifetime [1271]. Lifetime-Based [1271]. Light [1099, 506, 244, 186, 801]. Lighting [37]. Like [936, 365, 437, 119]. limit [4]. Limitations [1220]. Limits [1256, 462]. line [53]. Linear [1179, 838, 651, 185, 184]. Linearizability [921]. Linearizability-based [921]. Linguistics [930]. Link [1114]. Liquid [543, 544]. Little [207]. Live [1044]. Lives [979, 510, 636]. living [588]. LLM [1266]. LLM-Generated [1266]. load [735]. loads [734]. Locality [436]. localization [363]. Localizing [362, 1207]. Locating [1138]. location [203]. lockdown [474]. Locked [1282]. Locking [992]. logic [446]. logical [421]. Long [999, 615]. Long-Term [999]. loo [460]. look [365]. looking [810]. loop [699]. losing [27]. Lost [188, 1046, 857]. lottery [597]. Low [1226, 986, 878, 408, 1133]. Low-Code [1133]. Low-Depth [1226]. Low-Latency [986]. low-power [878]. Lower [228, 1226]. LTE [311]. Luce [803].

M [458, 345]. Machine

[1283, 964, 1069, 935, 1271, 68, 1224, 936, 945, 1148, 1078, 795, 117, 462, 16, 539, 9, 298, 563, 635, 819, 247, 619]. machine-manipulated [539]. Machines [1241, 66, 953, 54, 304, 730]. mad [683]. Made [1267]. MadMax [226, 225]. Madras [863]. Magellan [183, 182]. Make [1083, 509, 8, 195, 111]. maker [351]. makes [807]. makeup [767]. Making [907, 914, 904, 1012, 993, 396, 805]. mall [715]. malpractice [410]. malware [831]. Management [1059, 910, 1122, 1189, 1188, 1052, 995, 1271, 974, 1258, 8, 214, 295, 531, 613, 833, 32, 865, 370, 682, 196, 753, 702]. Manager [1245]. Managing [518, 90, 372]. manifesto [784]. Manifold [1091]. manipulated [539]. Manufacturing [1067, 1070, 101, 354]. Many [1116, 918, 848]. Map [955]. Mapping [960]. maps [575]. Margo [667]. Market [706, 92, 151]. Marketing [92]. Marketplace [1107]. Markets [1088]. markings [216]. marrying [332]. masses [107]. Master [72]. Mastering [17]. matching [183, 35, 182]. materiality [137]. Materials [970, 286]. Math [1013, 1140, 741]. mathematical [895]. Mathematicians [984]. matrix [735]. matter [772, 636]. maturity [390]. MAX [295]. Maximal [924]. Maximum [1179, 1178]. Maxwell [432]. may [236, 115]. Me [1009, 244, 452, 404, 715, 475, 608, 502]. Mean [940, 1032, 221]. Meaning [727]. measures [823, 455]. Measuring [824, 120, 282, 1246]. Mechanical [984, 538]. Mechanism [93]. mechanisms [396]. Media [1127, 85, 1136, 1273, 539, 107, 151]. mediated [641]. Medical [166, 558, 1183, 247, 352]. medicine [608]. Meet [68]. Meeting [77, 1167]. meetings [853]. Meets [1115, 703, 777, 424, 640, 30, 307]. megabytes [549]. Meltdown [136]. Members [925, 783, 763]. Memoriam [908, 1016, 1237, 830]. Memory [1270, 1271, 1173, 1174, 228, 134, 199, 136, 265]. Menlo [880]. mental [721]. Mentoring [999]. Mesh [986, 1284]. message [207]. metaverse [852]. meter [774]. Methods [752]. Metrics [957, 1003, 779, 778]. Mexico [252]. MEXT [1071]. micro [56, 564]. micro-program [564]. Microarchitecture [1194]. microfluidics [333]. microprocessor [861]. microservice [305]. microwave [601, 602]. microworlds [502]. Middleware [816]. Mike [667]. mile [837]. milestone [812]. mind [299]. Minding [249]. minicomputers [549]. minimal [651]. Minimum [1179]. Minimum-Cost [1179]. Mining [676, 554]. Minker [431]. MIP* [579]. Misinformation [1053]. Mismatch [1218]. Missed [1048]. missing [866, 855]. Mission [1187, 1228, 343].

Mission-Critical [1228]. Misunderstand [1171]. misunderstanding [234]. misuse [462]. Mitigate [1235]. mitigating [120]. Mix [1096]. Mixed [1093, 568]. ML [34, 1285, 1287, 1062]. ML-centric [34]. MobiGuide [701]. Mobile [855, 689, 442, 627, 498, 670]. Mobile-app [855]. Model [1264, 650, 1115, 963, 1114, 813, 58, 505, 471, 814]. model-driven [58]. Modeling [1004, 1076, 871]. Models [1073, 1103, 1202, 1133, 1094, 936, 1222, 705, 776]. moderation [297]. Modern [718, 1288, 876]. modulation [297]. modulo [420]. momentum [27]. Monetizing [612]. monitoring [499, 60]. Monopoly [1021]. monster [770]. Moore [317, 43]. moral [54]. morality [681]. Most [1141, 94, 580]. Motives [1180]. motor [628]. move [235]. movement [351]. moves [330, 884]. Moving [1110, 795, 124, 311]. MRI [368]. mugs [172]. Multi [577, 496]. multi-item [496]. Multi-itinerary [577]. Multilingual [1211, 375]. Multimedia [493]. multiparty [306]. multiple [877]. Multiplication [4, 735]. Multisensory [63]. music [248, 502]. Must [1284, 273, 856, 789]. mutation [896]. my [148]. mythos [795].

name [11]. Naming [923, 504]. Nand [1288]. Nano [1083]. Nano-Currency [1083]. nanosats [878]. nanosheet [43]. Nation [82]. National [864, 1056, 328]. Nations [79]. Nationwide [1054]. Natural [1202, 234, 566, 386, 563, 780]. nature [437, 806]. Navigates [950]. Navigating [274]. Near [1137, 800, 735, 675]. Near-Field [1137]. near-perfect [735]. near-term [675]. neat. [9]. Need [1105, 1219, 109, 217, 488, 438]. needs [487, 10, 659]. Neighborhood [927]. neighbors [800]. NeRF [626]. nets [360]. Network [1178, 833, 876, 228]. Networking [1034, 986, 392, 876]. Networks [1045, 952, 985, 1091, 868, 719, 120, 283, 263, 235, 647, 159, 311, 695, 119, 284]. Neumann [9]. Neural [625, 719, 1045, 952, 624, 868, 159, 626]. neuro [715]. neuro-symbolic [715]. neuroage [132]. neurological [566]. Neuromorphic [168, 1055]. Neurosymbolic [829]. Neurotechnology [962, 788]. News [85, 909, 24, 126, 211, 367, 479, 294, 552, 433, 317, 530, 807, 851, 657, 719, 610, 1040, 991, 905, 1098, 1163, 1255, 1209, 213, 458, 345, 431, 830, 908, 1016, 1237, 1121, 1185, 191, 43, 168, 235, 6, 106, 455, 292, 506, 551, 588, 369, 766, 679, 831, 745, 884, 971, 1013, 930, 1100, 1076, 1234, 1186, 45, 107, 234, 149, 26, 212, 529, 480, 587, 408, 319, 721, 852, 612, 808, 680, 1144, 953, 931, 1077, 1015, 1257, 4, 1208, 505, 432, 786, 633, 5, 271, 170]. News [507, 318, 481, 635, 744, 787, 809, 853, 681, 1120, 907, 1014, 954, 1143, 1235, 25, 192, 344, 368, 457, 659, 190, 147, 720, 1119, 1042, 1278, 1210, 44, 210, 169, 105, 233, 430, 528, 293, 342, 406, 586, 658, 765, 743, 829, 883, 970, 1075, 952, 929, 1142, 1184, 1276, 125, 127, 269, 148, 407, 343, 429, 742, 634, 992, 1012, 906, 1041, 1165, 1277, 128, 270, 456, 553, 832, 611, 788, 885, 767, 972, 1164, 1099, 993, 1256, 1236, 482, 709, 645, 710, 870, 241]. next [317, 419]. NFC [1138]. NFT [1097]. NICs [802]. NII [77]. Niklaus [1237]. Nim [846]. NLP [367]. No [1065, 1172, 574, 466]. No-Code [1065]. noise [868]. Non [507, 388, 372]. Non-fungible [507]. non-psychopath [372]. Non-traditional [388]. nonobvious [258]. Nonverbal [1204]. North [180]. novel [446]. novelty [645]. NP [623]. NPTEL [863]. NSA [307]. NSF [1000]. Nullify [1095]. Numbers [29, 135]. Nurturing [1061]. nutrition [855]. NVIDIA [1189].

OAuth [120]. Obfuscation [1251, 479]. Obituary [937]. Objects [1137, 1204, 1138]. obsessive [580]. obstacles [759]. Obtain [452]. OCCAM [980]. OCCAM-v2 [980]. occupations [541]. Oceania [84, 78]. odds [383]. off [442, 683]. off-the-shelf [442]. Offer [65]. Office [1239]. Offline [1052]. offload [801]. oil [379]. Old [1143, 298, 527]. Olson [667]. OMSCS [174]. On-skin [688]. Onboarding [1264]. One [1007, 227, 892, 815]. One-Way [1007]. ones [669]. Online [1243, 1260, 1290, 1052, 124, 588, 647, 189, 641, 237, 524]. Open [1082, 936, 992, 1035, 490, 1, 585, 372, 463, 726, 716]. Open-Conference [1035]. Opening [32, 1051]. OpenPGP [1033]. operating [733, 732]. Operational [1245]. Operationalizing [272, 1057]. operations [184]. Opinion [1171, 1127, 1283, 1170, 1126, 1238, 1243, 1188, 1106, 1282, 1166, 1262, 1260, 1151, 1216, 1211, 1146, 1259, 1263, 1125, 1105, 1167, 1242, 1150, 1215, 1128, 1217, 1192, 1284, 1148, 1104, 1193, 1107, 349]. opportunities [599, 870]. Opportunity [1048, 609, 173, 717]. Optical [1040]. Optimal [497, 104, 496]. optimization [577]. Optimizations [33, 966]. Optimizing [1048]. opus [892]. Orbit [629, 408]. organizations [540]. Organizing [74]. Oriented [1057]. original [178]. OSS [981]. Other [996, 1230, 405, 268]. Our [749, 738, 1105, 1101, 1200, 171, 815]. out-of-gas [226]. Out-of-this-world [354]. outdoors [362]. outer [427]. Outlook [906]. Outsmarting [1042]. oven [602]. Overcoming [944]. overfit [360]. overhead [309]. Overrated [1238]. oversensing [129]. Overwhelmed [961]. Own [1083, 500].

P [908, 623]. Pacific [85, 79]. packet [802, 801]. paddling [165]. Pan [692]. Pan-European [692]. pandemic [123, 236, 622, 473]. pandemic-driven [473]. Pandemics [1092]. panopticon [115]. panorama [254]. panoramic [386].

Paradigm [1110, 418]. paradox [548, 501]. parallelism [592]. Parameterized [1268]. Parchment [1019]. Park [880]. Part [1217, 1145, 1043]. participation [273, 371, 171, 180, 532]. Participatory [997]. Partnership [696]. Partnerships [97]. Pascal [350]. passionate [48]. Past [907, 776, 198, 885]. Patch [64]. Patent [593]. Path [916, 1098, 1019, 515, 683]. Paths [1036, 989, 1232, 950, 1096, 1182, 903, 290, 848, 452, 404, 549, 715, 340, 722, 230, 654, 608, 502, 805]. patient [352]. patient-clinician [352]. patients [701]. Pattern [1057]. Pattern-Oriented [1057]. patterns [198, 329]. payback [231]. paying [232]. Payments [1088]. Peer [951, 1059, 754]. Peer-Review [1059]. Peer-to-Peer [951]. People [1211, 648, 628, 177, 111, 209, 649, 339, 569]. Perception [1228]. perfect [735]. Performance [966, 1027, 923, 576, 742, 114]. Performs [1030]. Perils [1191]. period [96]. perish [2]. permissioned [320]. permissionless [320]. perpetuator [609]. Persistence [817, 438]. Persistent [912, 199]. Personal [1206, 1192, 1068, 469, 612]. Personalized [648, 1068, 705]. Perspective [1159, 1227, 85, 922, 1158, 1290, 1137, 1030, 1270, 1058, 1242, 985, 1032, 94, 965, 1225, 1007, 1135, 1051, 1178, 945, 1203, 1250, 1114, 601, 578, 672, 333, 360, 523, 227, 446, 757, 140, 898, 648, 674, 823, 709, 878, 496, 625, 821, 362, 262, 283, 308, 37, 627, 18, 310, 576, 498, 161, 780, 399, 59, 335, 448, 257, 225, 732, 842, 650, 521, 755, 184, 285, 707, 163, 473, 545, 471, 421, 203, 205, 603, 734, 423, 558, 182, 778, 119, 543, 264, 799, 844, 801]. perspectives [890]. Ph.D [452]. Phantom [982]. phase [694, 419]. phases [230]. Philosopher [950]. philosophy [10]. phone [689, 543]. photography [897]. Photonic [506]. Physical [821, 1064, 728]. Picking [656]. Picture [64, 1193]. Pieces [1267]. pillars [705]. PiM [1024]. pipeline [877]. Pipelines [1228]. Pitting [268]. Pivot [60, 59]. PL [495, 1141]. PL/I [1141]. place [155]. placebo [836]. Placement [1291]. PlanAlyzer [524]. planet [164]. planet-scale [164]. planning [748]. Platform [1122, 491, 490, 690, 139, 540]. Platformizing [8]. Platforms [1127, 1062, 34, 833, 108, 346]. Playing [525]. Plays [972, 500, 130]. plea [412]. Please [998, 717]. pocket [301]. Point [813, 814, 959]. Point/counterpoint [813, 814]. Pointers [1173]. points [122]. PoIs [648, 649]. Poisoning [946]. Polanyi [324]. Polarization [1053]. policing [7]. political [10]. politics [276]. polling [229]. Polymorphic [822]. Polynomial [1268]. Polynomial-Time [1268]. Polynomials [1225]. Pondering [1233]. Portal [1155]. possibility [623, 396]. possible [3]. Post [929, 856, 615]. Post-Quantum [929, 615]. post-secondary [856]. Potential [1005, 391, 392]. poverty [529]. Power [177, 1014, 1038, 878, 342, 755]. Powered [1182]. powerful [283]. practical [765]. Practicality [1006, 733]. Practice [917, 1109, 1002, 1060, 571, 875, 486, 109, 569]. Practices [964, 1280, 1062, 824, 413]. praise [485]. precise [845]. Predict [1254]. prediction [205]. predictions [771]. preeminent [378]. Prefixes [1081]. Premature [937]. prepares [387]. Preprocessing [1268]. present [776, 198, 568]. presents [446]. Preserving [1044, 677, 647, 885, 976, 277, 695, 229, 818]. President [925, 96]. pressure [499]. prestige [893]. Prevent [1027]. preventing [124]. price [227, 849]. Principle [1035]. Principles [1089, 1288, 1062, 272, 482]. priorities [581]. Prioritize [983]. Priority [1228]. Privacy [1044, 960, 236, 347, 460, 770, 855, 957, 1240, 1149, 174, 1074, 555, 695, 1195, 229, 109, 662, 704, 277, 462, 37, 193, 545, 326, 849, 818, 501, 332, 38].

Privacy-Preserving [1044, 695, 229, 277, 818]. Private [1090, 1091]. privilege [609]. probes [782]. Problem [73, 1098, 1208, 915, 1178, 140, 886, 261, 335, 750, 421]. Problems [1268, 175, 218, 608, 423]. Process [997, 690, 52, 702]. processes [244]. Processing [567, 802, 386, 563, 571, 517, 801]. processors [506]. producers [599]. Producing [894]. Product [1036]. Production [694]. Productivity [1153, 1246, 440, 428]. Products [923, 1207]. Profession [1017, 1168, 955, 1103, 1241, 195, 47, 274, 131, 511, 590, 436, 349, 812, 886, 746, 40]. Professional [1212, 71, 436, 518, 35, 541]. professionals [410, 590]. Profiling [916]. Program [1003, 980, 451, 862, 374, 630, 839, 863, 564]. programmer [850]. Programming [333, 996, 1133, 1141, 913, 937, 672, 484, 817, 729, 450, 397, 865, 334, 557, 673, 132, 502]. Programs [1249, 1250, 491, 218, 660, 611, 727]. Progress [205, 214, 716]. Project [1071, 372, 887]. projects [463, 527]. promise [605]. promises [214]. Proof [1134, 1249, 105]. Proofs [1075, 578]. Proposal [151, 122]. protect [236]. Protecting [508, 209, 982, 1183, 976, 603]. Protection [1174, 604, 38]. protections [326]. Protein [991]. protocols [333]. prototype [448]. Provable [1032]. provenance [596]. providers [237, 300]. providing [388]. Proving [946]. provocateur [736]. Provocation [1238]. proximity [818]. Prutor [865]. Psychology [1121]. psychopath [372]. Public [988, 153]. publication [656, 289]. Publications [973, 1187, 103, 589]. Publicly [923]. Publish [2]. purpose [355]. purposes [303]. pursuit [491, 380, 326]. push [346]. pushback [668]. put [815].

Putting [996, 138]. Puzzles [1255].

Python [1165].

Q&A [1194]. Qatar [381]. QCRI [378]. QR [926]. Quantum [910, 905, 1006, 883, 929, 1142, 1276, 516, 703, 674, 840, 24, 309, 308, 796, 605, 615, 515, 147, 210, 342, 765, 675]. quantum-computing [840]. Quantum-safe [516]. Qubit [551]. Qubits [1163]. query [899]. querying [781]. quest [580, 496]. Question [94]. Questions [1239, 762]. Quis [827].

R [1026]. R&D [255]. Race

[446, 291, 668, 730, 516, 447]. radiance [625, 626]. radios [821]. Raising [720]. Rampant [942]. random [244]. Randomness [1022, 244]. Randomower [547]. range [336, 400]. rank [422]. Ransomware [1088, 369]. rapid [622]. Rare [1135]. ratchets [831]. rate [651]. Rateless [735]. Re [942, 579]. reaching [144]. reaction [503]. read [618]. Reading [115, 136]. Real [1227, 86, 1233, 954, 222, 1062, 490, 274, 766, 442, 545, 562]. Real-Time [1227, 1062, 490, 274, 442]. Real-World [86, 545]. Realistically [1006]. Reality [1210, 1128, 235, 551, 587, 318, 540, 451, 568]. Realization [1071]. really [484]. Reason [241]. Reason-checking [241]. Reasoning [1134]. Reasons [983, 540]. Reboot [289]. Rebuttal [1193]. Rebuttals [1106]. Rebutting [1106]. receiver [879]. receivers [878, 822]. recipe [603]. recognition [394, 344, 665]. recommendation [648, 779]. Recommendations [956, 282]. Recommender [700, 649]. reconnaissance [358]. Reconsidering [965]. reconstruction [336, 335]. record [176]. record-and-replay [176]. recruit [488]. red [664]. reduce [308]. Reducing [149, 891]. redux [267, 526]. referencing [314]. Reflecting [1280]. Reflections [371, 638, 1020, 1172, 1213, 9, 298, 373, 510]. Refugee [1291, 1290]. regarding [664]. Region [872, 871, 860, 867, 869, 868, 864, 572, 866, 862, 563, 571, 560, 859, 865, 874, 87, 873, 870, 561, 566, 861, 863, 564, 565, 569, 570, 562, 567, 568, 381, 379, 378, 377, 699, 389, 692, 703, 686, 696, 690, 693, 697, 687, 694, 699, 700, 691, 701, 695, 698, 689, 688, 702]. Regional [242, 248, 253, 250, 243, 255, 256, 251, 244, 257, 245, 247, 246, 249, 252, 254, 392, 338]. regrets [527]. Regulate [1095]. regulating [108, 483]. regulation [699]. regulations [347, 700]. Regulatory [1104]. Reimagining [644]. Reimplementing [459]. Reinventing [66, 1001]. Related [1018]. relational [185]. relationships [352]. release [462]. Relevant [1262]. Reliability [877, 12]. reliable [621, 311]. rely [631]. Remaining [533]. Remediation [1018]. Remembering [909, 764]. Remote [1092, 243, 258, 370, 413]. Removing [275]. renaissance [193]. renewable [340]. Repair [1120, 854]. Repairability [267]. replay [176]. replication [181]. Report [998, 798]. Repository [1130]. representation [710]. representations [138]. representing [626]. require [396]. requirements [443]. requires [361]. Rescue [1107]. Research [1201, 1291, 1247, 917, 1111, 942, 1134, 1159, 1199, 1227, 1220, 1154, 1249, 1179, 922, 1033, 1221, 1155, 946, 1202, 1175, 1158, 1290, 938, 966, 1136, 1132, 86, 1071, 1268, 1133, 1137, 1251, 1109, 1030, 1223, 1248, 875, 1112, 1113, 1052, 1270, 1177, 1267, 1226, 1228, 1271, 1176, 985, 1002, 1060, 986, 977, 1224, 1115, 87, 1280, 1032, 1156, 965, 1200, 923, 1289, 1157, 1225, 1288, 1135, 1131, 1222, 1079, 1051, 1269, 1287, 1266, 1178, 1204, 945, 1203, 1138, 1250, 1114, 1193, 1110, 1031, 1246, 798, 393, 872, 253, 296, 197, 347, 862, 256, 391, 378, 377, 687, 806]. research [439, 389, 196, 392, 109, 398, 548, 289, 723]. Researchers [999, 438]. Resettlement [1291]. Resilience [918, 978, 1284, 869, 821, 682, 98]. resist [789]. Resistance [1150, 515]. Resolution [756, 314]. Resolving [1285]. Resource [1290, 1271]. resourceful [803]. respecting [238]. Response [91, 622, 403, 689]. responses [319]. Responsible [461, 223, 1212, 1057, 486, 753, 948]. restoration [505]. result [406]. Results [942, 923, 232]. Rethinking [1151, 857, 361]. Return [1279]. Revamping [1165]. Revealing [114]. revenge [324]. Review [984, 1059, 1028, 1050, 944, 1029, 331, 668, 754]. Reviews [1136, 1135]. Revisited [923, 374]. Revisiting [1035]. revolution [174, 102]. revolutionary [839]. revolutionize [106]. Rewrite [966]. Rewrites [971]. rides [498]. Right [994, 975, 1096, 509, 854]. Rights [1028, 591]. rings [439]. RISC [106]. RISC-V [106]. Rise [1041]. rising [398]. Risk [995, 206]. Risks [1123, 1018, 129, 831, 28, 434, 215, 747]. Rites [1028]. Rituals [1254]. road [587, 808, 615]. roadmap [871, 520]. Robot [1116, 1223, 519]. Robotic [690]. robotics [696, 884]. Robots [1234, 1028, 1078, 825, 270]. robovoices [720]. Robust [423]. Robustness [424, 946]. role [443, 108, 152, 145, 741, 541, 114]. Roll [1234]. romance [324]. Room [933]. Roots [374]. Roulette [401]. Route [1036]. Router [1027]. routing [384, 445]. rule [412]. rules [500, 346]. Rust [397].

Safe [397, 945, 334, 237, 516]. Safety

[926, 1123, 122, 194]. Safety-Critical [1123]. Salary [315]. salsa [248]. same [122]. sampled [779, 778]. Sampling [800, 1039]. sand [501]. sand-heap [501]. Sanity [216]. SAT [1026]. satellites [879]. satisfiability [777]. satisfying [406]. SATURN [286]. Save [979, 851]. Saving [1164, 1183]. Say [1230, 484]. saying [484]. Scaffolding [968]. Scalable [336, 185]. Scale [73, 1117, 1057, 599, 899, 328, 335, 285, 522, 870, 164, 863, 567]. Scaling [494, 1170, 449]. scans [368]. scenario [207]. scene [625]. scenes [626]. Schedulable [965]. schedule [3]. Scheduling [1109]. schema [522]. Schemes [1032]. school [231]. Science [349, 991, 487, 1020, 1254, 1211, 1146, 928, 914, 1060, 1049, 999, 993, 1215, 1217, 1266, 273, 330, 835, 175, 181, 848, 671, 379, 762, 321, 251, 36, 292, 23, 478, 850, 640, 30, 687, 257, 325, 439, 116, 793, 792, 19, 180, 873, 196, 655, 158, 110, 40, 357, 488, 636, 724]. sciences [179, 255]. Scientific [1087, 409]. Scientist [944, 903, 594]. scientists [841]. Scott [791]. Scrambled [1154]. screen [169]. SDN [820]. seabed [454]. Search [1065, 1260, 800, 11, 780, 218, 428, 3]. Seattle [798]. second [279, 353]. secondary [856]. secret [718]. Secrets [1250]. Section [872, 242, 248, 253, 250, 871, 860, 867, 869, 868, 864, 866, 862, 243, 255, 256, 251, 244, 859, 865, 257, 245, 247, 246, 874, 249, 873, 870, 252, 976, 861, 863, 254, 384, 390, 381, 393, 692, 394, 703, 376, 572, 686, 696, 531, 386, 690, 379, 391, 378, 377, 383, 693, 563, 697, 571, 560, 687, 694, 387, 385, 699, 700, 691, 701, 695, 389, 698, 380, 392, 689, 688, 561, 566, 382, 388, 564, 565, 569, 570, 562, 567, 568, 702]. Sections [338]. Secure [1014, 306, 992, 109, 945]. Securing [454, 52, 546, 834, 445, 990]. Security [296, 509, 1240, 279, 981, 811, 615, 194, 995, 912, 409, 87, 1280, 1032, 1195, 1191, 1265, 1218, 723, 1102, 823, 824, 479, 174, 281, 618, 389, 545, 109, 1033]. Security-Enhancing [1191]. Seeing [186, 897, 127]. Seeking [585, 233]. Seesaw [1205]. segmentation [706]. SeL4 [86]. selection [140]. Self [214, 57, 680]. Self-driving [214, 680]. self-image [57]. Seltzer [667]. Semantic [331]. Semantics [1004, 264]. semantics-aware [264]. Semantics-Driven [1004]. Semiconductor [1054]. sense [233]. sensing [243, 285, 1137]. sensitive [303]. sensor [472, 447]. Sensors [1014, 1148, 543]. Sequencing [1050]. sequential [17]. series [466]. Seriously [1046]. Server [1270]. Serverless [1113, 419]. Service [1234, 494, 520, 577, 237]. Services [74, 81, 692, 240]. Set [733]. Setting [1256]. shake [820]. SHAKTI [861]. Shape [905, 168]. shapes [848]. Shaping [450, 556, 171]. Shared [1105]. Sharpening [1085]. shedding [244]. shelf [442]. shield [718]. Shining [1099]. ship [803]. Shoes [914, 837]. Shonan [77]. shoppers [26]. shopping [715]. shortage [464]. shortages [744]. Shortcut [1202]. Shortest [1098, 3]. Shortest-Path [1098]. Should [1199, 1240, 1198, 1211, 1028, 1200, 841, 1083, 949, 582, 344, 790, 594, 240, 419, 463, 669]. Show [1132]. Shrinking [610]. SI [1081]. side [211]. side-channel [211]. sides [453]. siege [237]. signal [336, 335]. significant [580]. Silent [1026]. Silicon [276, 857]. Simba [449]. simple [160, 9]. Simplicity [941]. simulations [719]. Singapore [81, 82, 80, 83]. situ [591]. Six [540]. sized [363]. Sizing [23]. sketching [780]. Skew [1260]. skills [31, 564]. skin [897, 688]. Skins [1124]. SkyCore [310, 311]. Sliders [975]. Small [1267, 599, 879, 21, 844]. small-scale [599]. Smallness [1103]. Smart [82, 509, 226, 245, 225, 639, 38]. smarter [340]. smartphone [631, 621, 212, 513, 544]. smartphone-based [621]. Smartphones [1224]. Smarts [953]. smartwatches [758]. Smell [1236, 523]. SMS [279]. snapshot [117]. Snatch [1272]. Social [1127, 85, 1136, 1273, 1107, 201, 179, 224, 709, 647, 107, 710, 35, 151]. Societal [1192]. Society [1127, 1071, 967, 1084, 1022, 487]. Softening [270]. Software [1126, 1199, 410, 935, 602, 912, 1245, 936, 992, 1269, 537, 330, 703, 491, 464, 802, 708, 895, 821, 455, 750, 28, 887, 624, 459, 768, 413, 463, 453, 550, 527, 665, 114]. Software-defined [602]. SolarWinds [366]. Solid [1192]. Solution [1178, 601, 870]. solutions [860, 183, 754]. Solvable [1268]. Solving [335, 633, 777, 500]. Some [160, 487]. sorting [228]. soul [298]. sound [758]. SoundWatch [758]. Source [1082, 936, 992, 616, 586, 372, 463, 726]. sources [388]. South [1054]. Space [1097, 14, 1094, 427, 674, 136, 803, 440]. space-traveling [803]. spaces [838, 691]. SPADE [596]. Spake [1166]. Spans [5, 393]. Spatial [470, 205]. spatiotemporal [204]. Speaking [1145, 956, 1043, 1239, 150, 237, 346, 459, 554, 768, 660, 854]. Special [872, 242, 248, 253, 250, 871, 860, 867, 869, 868, 864, 866, 338, 862, 243, 255, 256, 251, 244, 859, 865, 257, 245, 247, 246, 874, 249, 873, 870, 252, 861, 863, 254, 384, 390, 381, 393, 692, 394, 703, 376, 572, 686, 696, 386, 690, 379, 391, 378, 377, 383, 693, 563, 697, 571, 560, 687, 694, 387, 385, 699, 700, 691, 701, 695, 389, 698, 380, 392, 689, 688, 561, 566, 382, 388, 564, 565, 569, 570, 562, 567, 568, 702]. Specialization [980]. specializing [163, 164]. Specific [920, 157, 159]. Specifications [921]. Specifying [1201]. Spectre [162]. spectrum [649, 128]. speculation [603]. Speculative [604, 162]. speculatively [604]. Speech [1127, 976, 394, 571]. speed [719, 282, 4]. speeds [368]. Speedup [1142]. spent [222]. sphere [153]. spirit [125]. Split [961]. spot [576]. Spotify [55]. Spreads [991]. spying [115]. SQL [184]. SSD [1052, 1051]. stability [337]. stablecoins [156]. Stack [963]. stages [877]. standard [154]. Standards [1014]. Start [21, 565]. start-ups [565]. state [651, 369, 318, 819, 650]. states [273]. Static [642, 980, 620, 365, 666, 643]. static-analysis [643]. static-analysis-driven [666]. statistics [423]. Status [1285, 219]. Stay [312]. STEM [65]. step [421]. Stewart [857]. Stifles [1126]. Still [1034, 680, 750, 361]. stock [243]. stole [327]. Stone [1124]. stop [236, 841]. Stopping [61]. Store [1031]. stores [613]. Storied [1178]. stories [250]. Story [1009]. Storytelling [835]. Strategic [165, 380]. Strategies [1282, 966, 1193]. Strategy [1170, 910, 1122, 1189, 974, 1258, 8, 214, 295, 531, 613, 833, 370, 682, 805]. Streaming [1114]. Street [139, 575]. strength [399]. stricter [346]. Strides [993]. String [475]. strings [29]. Striving [1187]. structure [650]. structures [250, 799]. STT [604]. Student [69, 23]. students [167, 856, 31, 463, 187, 488]. Study [1155, 1287]. Studying [132]. subcentimeter [363]. subcentimeter-sized [363]. Subfield [893]. Subgraph [1065]. Subjects [996, 1285]. Subsetting [1001]. subway [155]. Success [957, 250, 855]. Succinct [400]. Sufficient [1260]. Summer [1093]. supercomputer [159]. supercomputers [567]. supercomputing [692, 251]. superheroes [555]. Superpolynomial [1226]. Supply [981, 457, 682]. Supply-Chain [981]. Support [74, 1071, 178]. Supporting [649, 184]. sure [28, 115]. SuRF [399]. Surface [1255]. surgical [598]. Surging [89, 789]. Surveillance [774, 70, 767]. survey [386]. Sustainability [1000, 1187, 826, 58, 404]. Sustainable [1156, 365, 582, 245, 388]. Swahili [1277]. Swarm [884]. sweet [576]. switch [530]. switched [651]. Symbol [986]. Symbol-Synchronous [986]. Symbolic [945, 777, 715]. Synchronous [986]. syndrome [438]. synergy [739]. synthesis [626, 472]. synthetic [731]. System [1261, 1288, 448, 865, 733, 546, 732, 185, 649, 747, 698, 1066]. Systematic [1049, 675]. Systems [1111, 951, 1227, 1244, 86, 935, 1123, 1064, 393, 494, 651, 220, 396, 397, 60, 25, 504, 619, 700, 35, 517, 728, 489, 289, 562].

tacit [324]. Tackle [1053, 608, 423]. Tackling [266]. Tactics [1193]. taint [604]. Take [981, 1075, 1046, 168, 678, 717]. takes [294, 310, 650, 293, 726]. Taking [1025, 432]. Tale [1088, 670]. Tales [1163, 477]. Talking [291, 1222, 197]. Taming [1228]. Tandem [1289]. Tapping [1114]. targeting [119]. targets [656]. task [858]. Taxonomy [75]. Teach [1116, 351, 405]. Teachers [65, 405]. teaches [832, 98]. Teaching [1209, 102, 65, 405, 189, 1074, 562, 218, 532]. Teams [961]. Teaspoon [996]. Tech [958, 963, 902, 1095, 535, 718, 483, 300, 339, 714]. Tech-Stack [963]. Technical [601, 578, 672, 898, 1227, 674, 922, 262, 1158, 1290, 308, 37, 18, 310, 576, 780, 1137, 1030, 1270, 842, 985, 1032, 94, 965, 1225, 421, 1007, 1135, 603, 1051, 423, 1178, 778, 945, 1203, 1250, 1114, 264, 844, 333, 523, 227, 446, 757, 140, 648, 823, 709, 878, 496, 625, 821, 362, 283, 627, 498, 161, 399, 59, 335, 448, 225, 732, 785, 650, 521, 755, 184, 285, 707, 163, 473, 545, 471, 203, 205, 734, 182, 119, 543, 799, 801, 360]. Techniques [16, 258, 358, 109, 566]. TECHNOLOchicas [171]. Technological [319, 1192, 246]. Technologies [271, 559, 885]. Technology [994, 1171, 8, 214, 295, 531, 613, 833, 910, 1122, 1189, 131, 1279, 967, 1144, 681, 1190, 989, 1101, 722, 1096, 911, 370, 682, 974, 1258, 1207, 804, 860, 137, 508, 410, 810, 108, 10, 7, 193, 614, 297, 170, 481, 112, 13, 355, 300]. telementoring [598]. Telephones [990]. Tell [1132, 1009, 9]. Temptation [70]. Terahertz [235]. TeraWatt [1072]. TeraWatt-Hours [1072]. Term [999, 675]. terrains [206]. Test [1283]. Test-Driven [1283]. Testimony [1161]. Testing [1244, 512, 1269, 543, 544]. Tetris [1288]. Text [554, 1266, 882, 526, 1132]. Textiles [1137, 1138]. Thanks [716]. thee [631].

Their [1249, 1132, 646, 360, 837, 618, 841, 109]. them [488]. Theoretic [1159]. Theoretical [1050, 1176, 257]. theoretician [48]. theories [420]. Theory [872, 218]. Therapy [1210]. There [574, 1172, 1261, 18, 466, 488]. thereof [434]. thermodynamically [437]. These [810]. Thing [994, 907]. Things [1199, 759, 844, 1247, 520, 166, 1240, 456]. Think [62, 323]. thinking [590, 418, 793, 17]. third [498, 142]. Thirsty [1278]. Thorny [175]. Thought [1069, 197, 9]. Thoughts [1253]. threat [147]. threaten [439]. Threats [1127, 181, 995, 912, 281, 524]. Three [250, 1217, 670, 287]. Three-Part [1217]. throughout [321, 533]. Thus [1166]. Thwarting [211]. Tickets [72]. Tidy [1139]. Time [1227, 1179, 426, 1268, 1128, 1062, 490, 338, 274, 530, 442, 14, 316, 39, 239, 327, 466, 222]. time-series [466]. times [537]. TinyML [1156]. today [360]. together [495]. token [120]. tokens [507]. tolerance [309]. Tolerant [1276]. tomorrow [698]. Too [1281, 1211, 1273, 774]. tooling [862]. Tools [1085, 896, 666, 109, 643]. top [873]. topic [645]. Topics [248, 250, 867, 868, 1059, 1065, 1061, 862, 243, 1067, 1063, 244, 1071, 1066, 1069, 865, 1070, 245, 1057, 1058, 247, 246, 1060, 249, 1064, 870, 1068, 1062, 384, 381, 692, 696, 690, 379, 378, 377, 383, 693, 563, 697, 687, 694, 385, 691, 695, 380, 689, 688, 561, 566, 382, 564, 565, 562, 567]. Topological [1163, 651]. topromote [404]. total [747]. total-system [747]. Touch [1229, 1080]. tour [256]. Towers [1139]. Toxicity [1260]. Tracing [916, 236, 59, 621, 60, 818]. track [844]. Tracking [26, 473, 269, 1135, 1182, 845, 604]. tradition [393]. traditional [388]. Traffic [384, 843, 907, 842, 474, 408, 130, 473]. tragedy [531]. Trained [1162]. training [360, 492, 159]. Transform [756, 1142]. transformation [201, 352]. Transformed [1209]. Transformers [398]. Transforming [792]. transistors [43]. Transition [977, 305, 615]. Transitioning [146]. translation [188]. Transparent [1082, 790]. trapped [675]. traveling [803]. treadmill [113]. treat [799]. Treatment [1049]. Trees [946, 447]. Trends [872, 253, 871, 860, 869, 864, 866, 255, 256, 251, 687, 1054, 1056, 257, 1055, 874, 873, 1053, 252, 861, 863, 254, 390, 393, 394, 703, 572, 386, 391, 571, 387, 699, 700, 701, 389, 698, 392, 398, 561, 357, 388, 569, 570, 568, 702]. triad [281]. Trials [88]. Trick [92]. tricky [283]. trilemma [402]. trip [715]. trolley [307]. Trouble [586, 632]. troubling [665]. True [1009, 313]. Trust [1198, 1084, 940, 699, 1022, 24, 516]. Trusted [1017]. Trustworthiness [1201, 693, 747]. Trustworthy [86, 409, 542]. truth [881, 151]. tunnel [186]. Turing [646, 467, 426, 895, 9, 503]. Turings [477]. Turning [959]. turnover [518]. Twins [1067, 1020, 1119, 1191, 705]. Two [227, 1177, 1088, 722, 453, 426, 694, 477]. tyranny [61].

U.K. [56]. U.S. [273, 23, 893, 670, 1239]. UAV [311]. UAV-based [311]. Ugly [1233]. UiPath [690]. UIs [643]. Ulterior [1180]. unappreciated [167]. uncanny [787]. Uncertain [943]. uncertainties [96]. Uncertainty [1248]. Uncloneable [1157]. unconventional [383]. Uncover [1090]. Underbelly [1233]. Undergrads [65, 189]. Undergraduate [1215, 180]. undergraduates [562]. underserved [404]. Understanding [248, 1202, 412, 361, 743]. unexpected [492]. unfolds [230]. Unifying [1248]. union [483]. unionization [481]. United [380]. universal [9, 638]. universe [883]. Universities [69, 789, 1, 888]. university [377]. unlabeled [126]. Unleashing [390, 1038]. unlikely [236]. unlimited [428]. Unlock [1098]. Unlocking [1005, 391]. unplugged [411, 637]. unrestricted [849]. Unsafe [50]. untethered [311]. Untrusted [1264]. Unveiling [492]. Updates [995, 912]. Upholding [482]. ups [565]. Uruguay [246]. Usability [957]. usage [689]. Use [1004, 904, 303, 459, 300, 942]. Used [1105]. Useless [912]. User [965, 1203, 136]. User-Schedulable [965]. users [618, 559, 238]. Using [1264, 1121, 1254, 1137, 1045, 1015, 1055, 218, 989, 767, 238, 1138, 1068, 243, 602, 135, 710, 472, 543, 38]. UX [1036, 654].

V [106]. v2 [980]. Valérie [909]. Validating [1206]. validity [524]. valley [787, 857]. Value [1031, 204, 396, 203]. values [630]. Valuing [1079]. ’Vardi [403]. Varied [1093]. vector [735]. vehicle [214, 194, 639]. vehicles [516]. Vending [953]. Vendor [1282]. Vendor-Locked [1282]. Venues [1059]. Veracity [1058, 645]. Verification [1279, 455, 451]. verified [775]. Verifying [1252]. versions [435]. versus [613, 259]. via [228, 442, 244]. Vicious [29, 216, 435, 534, 485, 593, 887, 683, 748, 834, 637, 975, 933, 1080, 1169, 1019, 1124, 1214, 172, 275, 130, 322, 372, 790, 1261]. Video [1042, 492]. Vietnam [1061]. View [1091, 626, 215, 517]. Viewpoint [1000, 835, 890, 1021, 836, 175, 464, 197, 50, 277, 685, 11, 374, 414, 535, 462, 888, 617, 512, 1023, 351, 663, 487, 772, 938, 958, 616, 111, 412, 592, 10, 640, 30, 437, 935, 1082, 1045, 324, 749, 325, 1084, 31, 513, 978, 439, 218, 664, 815, 299, 857, 618, 323, 793, 771, 133, 1022, 112, 514, 352, 914, 49, 276, 977, 792, 979, 889, 594, 51, 641, 936, 559, 48, 240, 999, 915, 196, 998, 1083, 486, 132, 217, 413, 463, 153, 997, 558, 684, 154, 1081, 639, 300, 241, 375, 665, 913, 488, 858]. Viewpoint [350, 959, 438, 937, 1046, 725]. Virtual [1216, 66, 954, 1093, 192, 1210, 634, 1128, 146, 318, 853, 540]. virtualization [768]. viruses [209]. Vision [1287, 437, 110, 451]. visual [37, 781, 38]. Visualization [1132, 780]. visually [860, 271, 569]. VM [1196]. Vogels [353]. Voice [930, 1078]. voices [761]. vote [717]. voting [42]. VPN [395]. vs [623, 216, 1224, 153, 98, 339, 725]. vulnerability [223].

waiting [680]. wake [154]. Walden [287]. Walk [837]. walls [127]. War [714, 902]. warming [145]. Warranties [92]. wars [508]. Was [1172, 1160, 167, 9]. Waste [1235, 149]. Watch [927]. watching [102]. wave [498]. waves [474]. Way [966, 1007, 506, 37, 900, 466]. Wayfinding [1186]. Ways [1171, 932, 1116, 62]. web [843, 490, 25, 265, 264, 331, 1099, 1192]. Web3 [889]. WebRTC [490]. Weighing [231]. Welcome [242, 376, 686, 560, 859, 76, 1047]. Well [1251]. Well-Founded [1251]. Went [1213, 573]. were [373]. Werner [353]. Wheeler [756]. Where [977, 240, 187, 757, 466]. Whether [1233]. which [669]. while [273]. Whiskey [68]. Who [1262, 212, 948]. Whole [980]. Whole-Program [980]. whom [484]. Whose [513]. Wide [901]. width [422]. Wikipedia [375]. wild [707]. Will [6, 106, 1254, 981, 607, 174, 28, 109]. William [1016]. willing [294]. Win [902]. wind [44]. winners [807]. Winograd [522]. WINOGRANDE [521, 522]. Winter [1172]. Wireless [986, 337, 546, 822]. Wirth [1237]. wise [509]. wish [148]. within [179, 687, 699]. Without [1186, 574, 914]. wizard [880]. Women [510, 993, 866]. Won [1160, 167]. word [138]. Words [772, 138]. Work [91, 1092, 71, 1258, 189, 469, 258, 715, 370, 769]. workforce [387]. working [370]. Workloads [1270]. works [554]. World [384, 390, 381, 393, 394, 376, 580, 386, 379, 391, 378, 377, 383, 86, 387, 954, 385, 1232, 1169, 389, 1165, 380, 392, 382, 388, 1046, 56, 338, 354, 226, 635, 641, 545, 654]. worn [499]. worse [323]. worsening [369]. Worst [651]. Worst-case [651]. wrong [888, 573]. Wulf [1016]. WWII [167].

X [564]. XAI [1167]. XNOR [283, 284]. XNOR-networks [284, 283].

Year [1097, 474]. years [737, 623, 385, 350]. young [841]. yourself [144].

Zealand [1058]. Zero [939, 940, 313, 307]. zero-day [307].

References

Chien:2020:ELC

[1] Andrew A. Chien. Editor’s letter: Cracks in open collaboration in universities. Communications of the ACM, 63(1):5, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                              URL https://dl.acm.org/doi/abs/
                                              10.1145/3374766.

Vardi:2020:DPP

[2] Moshe Y. Vardi. Departments: Publish and perish. Communications of the ACM, 63(1):7, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                              URL https://dl.acm.org/doi/abs/
                                              10.1145/3373386.

Meyer:2020:BSS

[3] Bertrand Meyer. BLOG@CACM: In search of the shortest possible schedule. Communications of the ACM, 63(1):8– 9, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3371385.

Klarreich:2020:NMH

[4] Erica Klarreich. News: Multiplication hits the speed limit. Communications of the ACM, 63(1):11–13, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3371387.

Kugler:2020:NHI

[5] Logan Kugler. News: How the Internet spans the globe. Communications of the ACM, 63(1):14–16, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3371411.

Greengard:2020:NWD

[6] Samuel Greengard. News: Will deepfakes do deep damage? Communications of the ACM, 63(1):17–19, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3371409.

Joh:2020:LTI

[7] Elizabeth E. Joh. Law and technology: Increasing automation in policing. Communications of the ACM, 63(1):20–22, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3372912.

Cusumano:2020:TSMa

[8] Michael A. Cusumano. Technology strategy and management: ‘platformizing’ a bad business does not make it a good business. Communications of the ACM, 63(1):23–25, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3372918.

Haigh:2020:HRN

[9] Thomas Haigh and Mark Priestley. Historical reflections: von Neumann thought Turing’s universal machine was ‘simple and neat.’: but that didn’t tell him how to design a computer. Communications of the ACM, 63(1):26–32, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3372920.

Himmelreich:2020:VET

[10] Johannes Himmelreich. Viewpoint: Ethics of technology needs more political philosophy. Communications of the ACM, 63(1):33–35, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3339905.

Davis:2020:VSW

[11] James W. Davis and Jeff Hachtel. Viewpoint: A* search: what’s in a name? Communications of the ACM, 63(1):36–37, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3326600.

Sha:2020:REA

[12] Sanjay Sha. The reliability of enterprise applications. Communications of the ACM, 63(1):38–45, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3369756.

Ruoti:2020:BTW

[13] Scott Ruoti, Ben Kaiser, Arkady Yerukhimovich, Jeremy Clark, and Robert Cunningham. Blockchain technology: what is it good for? Communications of the ACM, 63(1):46–53, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3369752.

Helland:2020:STD

[14] Pat Helland. Space time discontinuum. Communications of the ACM, 63(1):54–56, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3369744.

Bagchi:2020:DEC

[15] Saurabh Bagchi, Muhammad-Bilal Siddiqui, Paul Wood, and Heng Zhang. Dependability in edge computing. Communications of the ACM, 63(1):58–66, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3362068.

Du:2020:TIM

[16] Mengnan Du, Ninghao Liu, and Xia Hu. Techniques for interpretable machine learning. Communications of the ACM, 63(1):68–77, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3359786.

Rajsbaum:2020:MCC

[17] Sergio Rajsbaum and Michel Raynal. Mastering concurrent computing through sequential thinking. Communications of the ACM, 63(1):78–87, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/abs/10.1145/3363823.

Guzdial:2020:TPT

[18] Mark Guzdial. Technical perspective: Is there a geek gene? Communications of the ACM, 63(1):90, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3372157.

Patitsas:2020:ECS

[19] Elizabeth Patitsas, Jesse Berlin, Michelle Craig, and Steve Easterbrook. Evidence that computer science grades are not bimodal. Communications of the ACM, 63(1):91–98, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3372161.

Shasha:2020:LBF

[20] Dennis Shasha. Last byte: Feedback for foxes. Communications of the ACM, 63(1):104–ff, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3372389.

Pancake:2020:ELS

[21] Cherri M. Pancake. Editor’s letter: Start small, then achieve big impact. Communications of the ACM, 63(2): 5, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3377932.

Cerf:2020:CD

[22] Vinton G. Cerf. Cerf’s up: On durability. Communications of the ACM, 63(2):7, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3377424.

Guzdial:2020:SUS

[23] Mark Guzdial. Sizing the U.S. student cohort for computer science. Communications of the ACM, 63(2):10–11, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3374764.

Edwards:2020:NLT

[24] Chris Edwards. News: Learning to trust quantum computers. Communications of the ACM, 63(2):13–15, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3374874.

Marks:2020:NDW

[25] Paul Marks. News: Dark web’s doppelgängers aim to dupe antifraud systems. Communications of the ACM, 63(2):16–18, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3374878.

Kirkpatrick:2020:NTS

[26] Keith Kirkpatrick. News: Tracking shoppers. Communications of the ACM, 63(2):19–21, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3374876.

Iglesias:2020:GCW

[27] Carlos Iglesias, Dhanaraj Thakur, and Michael L. Best. Global computing: Are we losing momentum? Communications of the ACM, 63(2):22–24, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/abs/10.1145/3376125.

Leveson:2020:IRY

[28] Nancy Leveson. Inside risks: Are you sure your software will not kill anyone? Communications of the ACM, 63(2):25–28, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376127.

Neville-Neil:2020:KVN

[29] George V. Neville-Neil. Kode vicious: Numbers are for computers, strings are for humans. Communications of the ACM, 63(2):29–30, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376630.

Hsu:2020:VWH

[30] Yen-Chia Hsu and Illah Nourbakhsh. Viewpoint: When human-computer interaction meets community citizen science. Communications of the ACM, 63(2):31–34, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376892.

Kussmaul:2020:VGS

[31] Clif Kussmaul. Viewpoint: Guiding students to develop essential skills. Communications of the ACM, 63(2):35–37, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376893.

Frazelle:2020:OBM

[32] Jessie Frazelle. Opening up the baseboard management controller. Communications of the ACM, 63(2):38–40, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3369758.

Godbolt:2020:OCC

[33] Matt Godbolt. Optimizations in C++ compilers. Communications of the ACM, 63(2):41–49, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3369754.

Bianchini:2020:TMC

[34] Ricardo Bianchini, Marcus Fontoura, Eli Cortez, Anand Bonde, Alexandre Muzio, Ana-Maria Constantin, Thomas Moscibroda, Gabriel Magalhaes, Girish Bablani, and Mark Russinovich. Toward ML-centric cloud platforms. Communications of the ACM, 63(2):50–59, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3364684.

Olsson:2020:DPS

[35] Thomas Olsson, Jukka Huhtamäki, and Hannu Kärkkäinen. Directions for professional social matching systems. Communications of the ACM, 63(2):60–69, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3363825.

Godefroid:2020:FHA

[36] Patrice Godefroid. Fuzzing: hack, art, and science. Communications of the ACM, 63(2):70–76, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3363824.

Gruteser:2020:TPL

[37] Marco Gruteser. Technical perspective: Lighting the way to visual privacy. Communications of the ACM, 63(2): 80, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3375569.

Zhu:2020:AVP

[38] Shilin Zhu, Chi Zhang, and Xinyu Zhang. Automating visual privacy protection using a smart LED. Communications of the ACM, 63(2):81–89, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3375571.

Hoffmann:2020:LBE

[39] Leah Hoffmann. Last byte: ‘everything fails all the time’. Communications of the ACM, 63(2):96–ff, February 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3374762.

Vardi:2020:DAC

[40] Moshe Y. Vardi. Departments: Advancing computing as a science and profession: but to what end? Communications of the ACM, 63(3):5, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3381047.

Staff:2020:LECa

[41] CACM Staff. Letters to the editor: Conferences and carbon impact. Communications of the ACM, 63(3):6–7, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3380448.

Hill:2020:CV

[42] Robin K. Hill. Coding for voting. Communications of the ACM, 63(3):8–9, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3379491.

Greengard:2020:NCN

[43] Samuel Greengard. News: Can nanosheet transistors keep Moore’s law alive? Communications of the ACM, 63(3):10–12, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379493.

Monroe:2020:NAH

[44] Don Monroe. News: Algorithms to harvest the wind. Communications of the ACM, 63(3):13–14, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379497.

Kirkpatrick:2020:NAL

[45] Keith Kirkpatrick. News: Across the language barrier. Communications of the ACM, 63(3):15–17, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379495.

Lachney:2020:ECC

[46] Michael Lachney and Aman Yadav. Education: Computing and community in formal education. Communications of the ACM, 63(3):18–21, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379918.

Denning:2020:PID

[47] Peter J. Denning and Dorothy E. Denning. The profession of IT: Dilemmas of artificial intelligence. Communications of the ACM, 63(3):22–24, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379920.

Reingold:2020:VTL

[48] Omer Reingold. Viewpoint: Through the lens of a passionate theoretician. Communications of the ACM, 63(3):25–27, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3378543.

OHara:2020:VFI

[49] Kieron O’Hara and Wendy Hall. Viewpoint: Four Internets. Communications of the ACM, 63(3):28–30, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3341722.

Canellas:2020:VUL

[50] Marc Canellas and Rachel Haga. Viewpoint: Unsafe at any level. Communications of the ACM, 63(3):31–34, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3342102.

Pierce:2020:VCE

[51] Benjamin C. Pierce, Michael Hicks, Crista Lopes, and Jens Palsberg. Viewpoint: Conferences in an era of expensive carbon. Communications of the ACM, 63(3):35–37, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380445.

Frazelle:2020:SBP

[52] Jessie Frazelle. Securing the boot process. Communications of the ACM, 63(3):38–42, March 2020.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379512.

Cook:2020:ALB

[53] Richard I. Cook. Above the line, below the line. Communications of the ACM, 63(3):43–46, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3379510.

Awad:2020:CMM

[54] Edmond Awad, Sohan Dsouza, Jean-François Bonnefon, Azim Shariff, and Iyad Rahwan. Crowdsourcing moral machines. Communications of the ACM, 63(3):48–55, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3339904.

Smite:2020:SG

[55] Darja Smite, Nils Brede Moe, Marcin Floryan, Georgiana Levinta, and Panagiota Chatzipetrou. Spotify guilds. Communications of the ACM, 63(3):56–61, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3343146.

Austin:2020:BMB

[56] Jonny Austin, Howard Baker, Thomas Ball, James Devine, Joe Finney, Peli De Halleux, Steve Hodges, Michał Moskal, and Gareth Stockdale. The BBC micro:bit: from the U.K. to the world. Communications of the ACM, 63(3):62–69, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3368856.

Fried:2020:ESI

[57] Ohad Fried, Jennifer Jacobs, Adam Finkelstein, and Maneesh Agrawala. Editing self-image. Communications of the ACM, 63(3):70–79, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3326601.

Kienzle:2020:TMD

[58] Jörg Kienzle, Gunter Mussbacher, Benoit Combemale, Lucy Bastin, Nelly Bencomo, Jean-Michel Bruel, Christoph Becker, Stefanie Betz, Ruzanna Chitchyan, Betty H. C. Cheng, Sonja Klingert, Richard F. Paige, Birgit Penzenstadler, Norbert Seyff, Eugene Syriani, and Colin C. Venters. Toward model-driven sustainability evaluation. Communications of the ACM, 63(3):80–91, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3371906.

Isaacs:2020:PPT

[59] Rebecca Isaacs. A perspective on pivot tracing: technical perspective. Communications of the ACM, 63(3):93, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3378931.

Mace:2020:PTD

[60] Jonathan Mace, Ryan Roelke, and Rodrigo Fonseca. Pivot tracing: dynamic causal monitoring for distributed systems. Communications of the ACM, 63(3):94–102, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3378933.

Shasha:2020:LBSa

[61] Dennis Shasha. Last byte: Stopping tyranny. Communications of the ACM, 63(3):104–ff, March 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3380444.

Pancake:2020:NWT

[62] Cherri M. Pancake. New ways to think about CS education. Communications of the ACM, 63(4):5, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cerf:2020:MA

[63] Vinton G. Cerf. Multisensory adventures. Communications of the ACM, 63(4):7, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2020:AMC

[64] CACM Staff. Adding more color to patch picture. Communications of the ACM, 63(4):9, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guzdial:2020:TTO

[65] Mark Guzdial. Teaching teachers to offer STEM to undergrads. Communications of the ACM, 63(4):10–11, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2020:RVM

[66] Leah Hoffmann. Reinventing virtual machines. Communications of the ACM, 63(4):128ff, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anthes:2020:DLC

[67] Gary Anthes. Dead languages come to life. Communications of the ACM, 63(4):13–15, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mone:2020:MLM

[68] Gregory Mone. Machine learning, meet whiskey. Communications of the ACM, 63(4):16–17, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shein:2020:HUD

[69] Esther Shein. How universities deploy student data. Communications of the ACM, 63(4):18–20, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubel:2020:TDE

[70] Alan Rubel and Kyle M. L. Jones. The temptation of data-enabled surveillance. Communications of the ACM, 63(4):22–24, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sako:2020:AIF

[71] Mari Sako. Artificial intelligence and the future of professional work. Communications of the ACM, 63(4):25–27, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2020:MT

[72] George V. Neville-Neil. Master of tickets. Communications of the ACM, 63(4):28–29, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Banga:2020:WCH

[73] Gaurav Banga. Why is cybersecurity not a human-scale problem anymore? Communications of the ACM, 63(4):30–34, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Girouard:2020:OFS

[74] Audrey Girouard, Jon E. Froehlich, Regan L. Mandryk, and Mark Hancock. Organizing family support services at ACM conferences. Communications of the ACM, 63(4):35–38, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grochow:2020:TAA

[75] Jerrold M. Grochow. A taxonomy of automated assistants. Communications of the ACM, 63(4):39–41, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moon:2020:W

[76] Sue Moon, Ann Nicholson, and Abhik Roychoudhury. Welcome. Communications of the ACM, 63(4):44–46, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kawarabayashi:2020:NSM

[77] Ken-Ichi Kawarabayashi. The NII Shonan meeting in Japan. Communications of the ACM, 63(4):48–49, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cunningham:2020:CCH

[78] Sally Jo Cunningham, Fariz Darari, Adila Krisnadhi, and Annika Hinze. Capturing cultural heritage in East Asia and Oceania. Communications of the ACM, 63(4):50–52, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rudolph:2020:CPI

[79] Carsten Rudolph, Sadie Creese, and Sameer Sharma. Cybersecurity in Pacific Island nations. Communications of the ACM, 63(4):53–54, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teh:2020:SCE

[80] Karen Teh, Vivy Suhendra, Soon Chia Lim, and Abhik Roychoudhury. Singapore’s cybersecurity ecosystem. Communications of the ACM, 63(4):55–57, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tat:2020:ISD

[81] Thomas Ho Chee Tat and George Loh Chee Ping. Innovating services and digital economy in Singapore. Communications of the ACM, 63(4):58–59, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teddy-Ang:2020:ASE

[82] Sintia Teddy-Ang and Abigail Toh. AI Singapore: Empowering a smart nation. Communications of the ACM, 63(4):60–63, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yan:2020:DAL

[83] Lam Kwok Yan and Campbell Wilson. Developing AI for law enforcement in Singapore and Australia. Communications of the ACM, 63(4):62, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bain:2020:DHA

[84] Chris Bain and Abraham Oshni Alvandi. Digital healthcare across Oceania. Communications of the ACM, 63(4):64–67, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cha:2020:DFN

[85] Meeyoung Cha, Wei Gao, and Cheng-Te Li. Detecting fake news in social media: An Asia–Pacific perspective. Communications of the ACM, 63(4):68–71, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heiser:2020:SAR

[86] Gernot Heiser, Gerwin Klein, and June Andronick. SeL4 in Australia: From research to real-world trustworthy systems. Communications of the ACM, 63(4):72–75, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phan:2020:ASR

[87] Raphaël C. W. Phan, Masayuki Abe, Lynn Batten, Jung Hee Cheon, Ed Dawson, Steven Galbraith, Jian Guo, Lucas Hui, Kwangjo Kim, Xuejia Lai, Dong Hoon Lee, Mitsuru Matsui, Tsutomu Matsumoto, Shiho Moriai, Phong Nguyen, Dingyi Pei, Duong Hieu Phan, Josef Pieprzyk, Huaxiong Wang, Hank Wolfe, Duncan Wong, Tzong-Chen Wu, Bo-Yin Yang, Siu-Ming Yiu, Yu Yu, and Jianying Zhou. Advances in security research in the Asiacrypt region. Communications of the ACM, 63(4):76–81, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2020:CTK

[88] Dong Ku Kim, Hyeonwoo Lee, Seong-Choon Lee, and Sunwoo Lee. 5G commercialization and trials in Korea. Communications of the ACM, 63(4):82–85, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cha:2020:ASI

[89] Sang Kil Cha and Zhenkai Liang. Asia’s surging interest in binary analysis. Communications of the ACM, 63(4):86–88, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maguire:2020:MHC

[90] Laura M. D. Maguire. Managing the hidden costs of coordination. Communications of the ACM, 63(4):90–96, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grayson:2020:CWH

[91] Marisa R. Grayson. Cognitive work of hypothesis exploration during anomaly response. Communications of the ACM, 63(4):97–103, April 2020. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woods:2020:CWM

[92] Daniel W. Woods and Tyler Moore. Cyber warranties: Market fix or marketing trick? Communications of the ACM, 63(4):104–107, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruderer:2020:AM

[93] Herbert Bruderer. The Antikythera mechanism. Communications of the ACM, 63(4):108–115, April 2020. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Procaccia:2020:TPA

[94] Ariel D. Procaccia. Technical perspective: An answer to fair division’s most enigmatic question. Communications of the ACM, 63(4):118, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aziz:2020:BEF

[95] Haris Aziz and Simon Mackenzie. A bounded and envy-free cake cutting algorithm. Communications of the ACM, 63(4):119–126, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pancake:2020:PHA

[96] Cherri M. Pancake. From the President: How ACM is adapting in this period of global uncertainties. Communications of the ACM, 63(5):5, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3392777.

Tims:2020:DPC

[97] Jodi Tims. Departments: Partnerships can help drive gender equality. Communications of the ACM, 63(5):7, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3388983.

Vardi:2020:DEV

[98] Moshe Y. Vardi. Departments: Efficiency vs. resilience: what COVID-19 teaches computing. Communications of the ACM, 63(5):9, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3388890.

Segal:2020:AEA

[99] Gerald Segal. ACM’s election: ACM’s 2020 general election. Communications of the ACM, 63(5):11–21, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386158.

Chien:2020:DLN

[100] Andrew A. Chien. Departments: Launching a new feature in Communications. Communications of the ACM, 63(5):22, May 2020.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3391913.

Canedo:2020:DAA

[101] Arquimedes Canedo. Departments: Automating automation: CS at the heart of the manufacturing economy. Communications of the ACM, 63(5):23, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3391911.

Guzdial:2020:TCH

[102] Mark Guzdial and Jiajie Zhang. Teaching CS humbly, and watching the AI revolution. Communications of the ACM, 63(5):24–25, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386312.

Davidson:2020:DAP

[103] Jack W. Davidson, Joseph A. Konstan, and Scott E. Delman. Departments: ACM publications finances. Communications of the ACM, 63(5):53–57, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3389687.

Shasha:2020:LBO

[104] Dennis Shasha. Last byte: Optimal chimes. Communications of the ACM, 63(5):112, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386912.

Monroe:2020:NPB

[105] Don Monroe. News: A proof from ‘the book’. Communications of the ACM, 63(5):27–29, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386366.

Greengard:2020:NWR

[106] Samuel Greengard. News: Will RISC-V revolutionize computing? Communications of the ACM, 63(5):30–32, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386377.

Kirkpatrick:2020:NDM

[107] Keith Kirkpatrick. News: Deceiving the masses on social media. Communications of the ACM, 63(5):33–35, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386375.

Hemphill:2020:LTW

[108] C. Scott Hemphill. Law and technology: What role for antitrust in regulating platforms? Communications of the ACM, 63(5):36–38, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386906.

Shostack:2020:PSS

[109] Adam Shostack and Mary Ellen Zurko. Privacy and security: Secure development tools and techniques need more research that will increase their impact and effectiveness in practice. Communications of the ACM, 63(5):39–41, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3386908.

Tissenbaum:2020:EVK

[110] Mike Tissenbaum and Anne Ottenbreit-Leftwich. Education: A vision of K–12 computer science education for 2030. Communications of the ACM, 63(5): 42–44, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3386910.

Hertzmann:2020:VCD

[111] Aaron Hertzmann. Viewpoint: Computers do not make art, people do. Communications of the ACM, 63(5): 45–48, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3347092.

Newport:2020:VWT

[112] Cal Newport. Viewpoint: When technology goes awry. Communications of the ACM, 63(5):49–52, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3391975.

Reed:2020:BFI

[113] J. Paul Reed. Beyond the ‘fixit’ treadmill. Communications of the ACM, 63(5):58–63, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380322.

Woods:2020:RCR

[114] David D. Woods and John Allspaw. Revealing the critical role of human performance in software. Communications of the ACM, 63(5):64–67, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380468.

Wicker:2020:RPY

[115] Stephen B. Wicker and Dipayan Ghosh. Reading in the panopticon — your Kindle may be spying on you, but you can’t be sure. Communications of the ACM, 63(5):68–73, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376899.

Mattauch:2020:BAD

[116] Sandra Mattauch, Katja Lohmann, Frank Hannig, Daniel Lohmann, and Jürgen Teich. A bibliometric approach for detecting the gender gap in computer science. Communications of the ACM, 63(5):74–80, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376901.

Chouldechova:2020:SFF

[117] Alexandra Chouldechova and Aaron Roth. A snapshot of the frontiers of fairness in machine learning. Communications of the ACM, 63(5):82–89, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376898.

Attiya:2020:I

[118] Hagit Attiya and Sergio Rajsbaum. Indistinguishability. Communications of the ACM, 63(5):90–99, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376902.

Voelker:2020:FLT

[119] Geoffrey M. Voelker. Fake ‘likes’ and targeting collusion networks: technical perspective. Communications of the ACM, 63(5):102, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3387722.

Farooqi:2020:MMO

[120] Shehroze Farooqi, Fareed Zaffar, Nektarios Leontiadis, and Zubair Shafiq. Measuring and mitigating OAuth access token abuse by collusion networks. Communications of the ACM, 63(5):103–111, May 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3387720.

Chien:2020:ELW

[121] Andrew A. Chien. Editor’s letter: What do DDT and computing have in common? Communications of the ACM, 63(6):5–6, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3397339.

Staff:2020:LES

[122] CACM Staff. Letters to the editor: Safety proposal points in same direction. Communications of the ACM, 63 (6):6, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3397169.

Cerf:2020:CIC

[123] Vinton G. Cerf. Cerf’s up: Implications of the COVID-19 pandemic. Communications of the ACM, 63(6):7, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3397262.

DeFranco:2020:DPI

[124] Terrence DeFranco and Jeremy Roschelle. Detecting/preventing infections, and moving instruction online. Communications of the ACM, 63(6):8–9, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3392370.

Savage:2020:NAS

[125] Neil Savage. News: An animating spirit. Communications of the ACM, 63(6):10–12, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3392520.

Edwards:2020:NLU

[126] Chris Edwards. News: Leveraging unlabeled data. Communications of the ACM, 63(6):13–14, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3392496.

Savage:2020:NST

[127] Neil Savage. News: Seeing through walls. Communications of the ACM, 63(6):15–16, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.1145/3392516.

Shein:2020:NHA

[128] Esther Shein. News: Hiring from the autism spectrum. Communications of the ACM, 63(6):17–19, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3392509.

Bolton:2020:IRH

[129] Connor Bolton, Kevin Fu, Josiah Hester, and Jun Han. Inside risks: How to curtail oversensing in the home. Communications of the ACM, 63(6):20–24, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3396261.

Neville-Neil:2020:KVK

[130] George V. Neville-Neil. Kode Vicious: Kode Vicious plays in traffic. Communications of the ACM, 63(6):25–26, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3396263.

Denning:2020:PIT

[131] Peter J. Denning and Ted G. Lewis. The profession of IT: Technology adoption. Communications of the ACM, 63(6):27–29, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3396265.

Siegmund:2020:VSP

[132] Janet Siegmund, Norman Peitek, André Brechmann, Chris Parnin, and Sven Apel. Viewpoint: Studying programming in the neuroage: just a crazy idea? Communications of the ACM, 63(6):30–34, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3347093.

Morris:2020:VAA

[133] Meredith Ringel Morris. Viewpoint: AI and accessibility. Communications of the ACM, 63(6):35–37, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3356727.

Frazelle:2020:CM

[134] Jessie Frazelle. Commit to memory. Communications of the ACM, 63(6): 38–41, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3386522.

Limoncelli:2020:CUN

[135] Thomas A. Limoncelli. Communicate using the numbers 1, 2, 3, and more. Communications of the ACM, 63(6): 42–44, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3386267.

Lipp:2020:MRK

[136] Moritz Lipp, Michael Schwarz, Daniel Gruss, Thomas Prescher, Werner Haas, Jann Horn, Stefan Mangard, Paul Kocher, Daniel Genkin, Yuval Yarom, Mike Hamburg, and Raoul Strackx. Meltdown: reading kernel memory from user space. Communications of the ACM, 63(6):46–56, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3357033.

Borning:2020:IMI

[137] Alan Borning, Batya Friedman, and Nick Logler. The ‘invisible’ materiality of information technology. Communications of the ACM, 63(6):57–64, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3360647.

Smith:2020:CWR

[138] Noah A. Smith. Contextual word representations: putting words into computers. Communications of the ACM, 63(6):66–74, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3347145.

Muhlhauser:2020:SLP

[139] Max Mühlhäuser, Christian Meurisch, Michael Stein, Jörg Daubert, Julius Von Willich, Jan Riemann, and Lin Wang. Street lamps as a platform. Communications of the ACM, 63(6):75–83, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376900.

Blum:2020:ASL

[140] Avrim Blum. Algorithm selection as a learning problem: technical perspective. Communications of the ACM, 63 (6):86, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3394623.

Gupta:2020:DDA

[141] Rishi Gupta and Tim Roughgarden. Data-driven algorithm design. Communications of the ACM, 63(6):87–94, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394625.

Hoffmann:2020:LBA

[142] Leah Hoffmann. Last byte: Attaining the third dimension. Communications of the ACM, 63(6):96–ff, June 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3392697.

Vardi:2020:DCL

[143] Moshe Y. Vardi. Departments: A computational lens on economics. Communications of the ACM, 63(7):5, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3402561.

Ozawa:2020:DCY

[144] Yosuke Ozawa. Departments: Challenge yourself by reaching for the highest bar. Communications of the ACM, 63(7):7, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3399746.

Staff:2020:LECb

[145] CACM Staff. Letters to the editor: Computing’s role in climate warming. Communications of the ACM, 63(7):9, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3402933.

Arquilla:2020:TDL

[146] John Arquilla and Mark Guzdial. Transitioning to distance learning and virtual conferencing. Communications of the ACM, 63(7):10–11, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3398386.

Mone:2020:NQT

[147] Gregory Mone. News: The quantum threat. Communications of the ACM, 63(7):12–14, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3398388.

Savage:2020:NYW

[148] Neil Savage. News: Your wish is my CMD. Communications of the ACM, 63(7):15–16, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3398392.

Kirkpatrick:2020:NRE

[149] Keith Kirkpatrick. News: Reducing and eliminating e-waste. Communications of the ACM, 63(7):17–19, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3398390.

Samuelson:2020:LSA

[150] Pamela Samuelson. Legally speaking: AI authorship? Communications of the ACM, 63(7):20–22, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3401718.

VanAlstyne:2020:EBDa

[151] Marshall W. Van Alstyne. Economic and business dimensions: Proposal: a market for truth to address false ads on social media. Communications of the ACM, 63(7):23–25, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3401724.

Pine:2020:CEI

[152] Kathleen H. Pine, Margaret M. Hinrichs, Jieshu Wang, Dana Lewis, and Erik Johnston. Computing ethics: For impactful community engagement: check your role. Communications of the ACM, 63(7):26–28, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3401720.

Stanger:2020:VCV

[153] Allison Stanger. Viewpoint: Consumers vs. citizens in democracy’s public sphere. Communications of the ACM, 63(7):29–31, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3359553.

Subirana:2020:VCW

[154] Brian Subirana. Viewpoint: Call for a wake standard for artificial intelligence. Communications of the ACM, 63(7): 32–35, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/abs/10.1145/3402193.

Helland:2020:BPB

[155] Pat Helland. The best place to build a subway. Communications of the ACM, 63(7):36–39, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3386269.

Clark:2020:DS

[156] Jeremy Clark, Didem Demirag, and Seyedehmahsa Moosavi. Demystifying stablecoins. Communications of the ACM, 63(7):40–46, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3386275.

Dally:2020:DSH

[157] William J. Dally, Yatish Turakhia, and Song Han. Domain-specific hardware accelerators. Communications of the ACM, 63(7):48–57, July 2020.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3361682.

Stodden:2020:DSL

[158] Victoria Stodden. The data science life cycle: a disciplined approach to advancing data science as a science. Communications of the ACM, 63(7):58–66, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3360646.

Jouppi:2020:DSS

[159] Norman P. Jouppi, Doe Hyun Yoon, George Kurian, Sheng Li, Nishant Patil, James Laudon, Cliff Young, and David Patterson. A domainspecific supercomputer for training deep neural networks. Communications of the ACM, 63(7):67–78, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3360307.

Catalini:2020:SSE

[160] Christian Catalini and Joshua S. Gans. Some simple economics of the blockchain. Communications of the ACM, 63(7):80–90, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3359552.

Hill:2020:WCC

[161] Mark D. Hill. Why ‘correct’ computers can leak your information: technical perspective. Communications of the ACM, 63(7):92, July 2020.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3399740.

Kocher:2020:SAE

[162] Paul Kocher, Jann Horn, Anders Fogh, Daniel Genkin, Daniel Gruss, Werner Haas, Mike Hamburg, Moritz Lipp, Stefan Mangard, Thomas Prescher, Michael Schwarz, and Yuval Yarom. Spectre attacks: exploiting speculative execution. Communications of the ACM, 63(7):93–101, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3399742.

Ranganathan:2020:ACS

[163] Parthasarathy Ranganathan. ASIC clouds: specializing the datacenter: technical perspective. Communications of the ACM, 63(7):102, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3399738.

Taylor:2020:ACS

[164] Michael Bedford Taylor, Luis Vega, Moein Khazraee, Ikuo Magaki, Scott Davidson, and Dustin Richmond. ASIC clouds: specializing the datacenter for planet-scale applications. Communications of the ACM, 63(7):103–109, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3399734.

Shasha:2020:LBSb

[165] Dennis Shasha. Last byte: Strategic paddling. Communications of the ACM, 63(7):112–ff, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3401749.

Cerf:2020:CIM

[166] Vinton G. Cerf. Cerf’s up: On the Internet of Medical Things. Communications of the ACM, 63(8):5, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3406779.

Arquilla:2020:HWW

[167] John Arquilla and Jeremy Roschelle. How WWII was won, and why CS students feel unappreciated. Communications of the ACM, 63(8):6–7, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3403958.

Greengard:2020:NNC

[168] Samuel Greengard. News: Neuromorphic chips take shape. Communications of the ACM, 63(8):9–11, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3403960.

Monroe:2020:NDH

[169] Don Monroe. News: Digital humans on the big screen. Communications of the ACM, 63(8):12–14, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3403972.

Kugler:2020:NWA

[170] Logan Kugler. News: Are we addicted to technology? Communications of the ACM, 63(8):15–16, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3403966.

Fernandez:2020:BPT

[171] Jannie Fernandez and JeffriAnne Wilder. Broadening participation: TECHNOLOchicas: a critical intersectional approach shaping the color of our future. Communications of the ACM, 63(8):18–21, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3408052.

Neville-Neil:2020:KVB

[172] George V. Neville-Neil. Kode vicious: Broken hearts and coffee mugs. Communications of the ACM, 63(8):22–23, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3408054.

Krishnamurthi:2020:EDC

[173] Shriram Krishnamurthi and Kathi Fisler. Education: Data-centricity: a challenge and opportunity for computing education. Communications of the ACM, 63(8):24–26, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3408056.

Galil:2020:PSO

[174] Zvi Galil. Privacy and security: OM-SCS: the revolution will be digitized. Communications of the ACM, 63(8): 27–29, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3366191.

Borgman:2020:VTP

[175] Christine L. Borgman, Michael J. Scroggins, Irene V. Pasquetto, R. Stuart Geiger, Bernadette M. Boscoe, Peter T. Darch, Charlotte Cabasse-Mazel, Cheryl Thompson, and Milena S. Golshan. Viewpoint: Thorny problems in data (-intensive) science. Communications of the ACM, 63(8):30–32, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3408047.

Staff:2020:CFR

[176] CACM Staff. To catch a failure: the record-and-replay approach to debugging. Communications of the ACM, 63(8):34–40, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386277.

Frazelle:2020:PP

[177] Jessie Frazelle. Power to the people. Communications of the ACM, 63(8): 41–45, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3397888.

Maiden:2020:DCS

[178] Neil Maiden, Konstantinos Zachos, Amanda Brown, Dimitris Apostolou, Balder Holm, Lars Nyre, Aleksander Tonheim, and Arend van den Beld. Digital creativity support for original journalism. Communications of the ACM, 63(8):46–53, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3386526.

Connolly:2020:WCB

[179] Randy Connolly. Why computing belongs within the social sciences. Communications of the ACM, 63(8):54–59, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3383444.

Payton:2020:EUC

[180] Fay Cobb Payton and Alexa Busch. Examining undergraduate computer science participation in North Carolina. Communications of the ACM, 63(8): 60–68, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3372122.

Cockburn:2020:TRC

[181] Andy Cockburn, Pierre Dragicevic, Lonni Besançon, and Carl Gutwin. Threats of a replication crisis in empirical computer science. Communications of the ACM, 63(8):70–79, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3360311.

Tan:2020:EMM

[182] Wang-Chiew Tan. Entity matching with Magellan: technical perspective. Communications of the ACM, 63(8):82, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3405464.

Doan:2020:MTB

[183] AnHai Doan, Pradap Konda, Paul Suganthan G. C., Yash Govind, Derek Paulsen, Kaushik Chandrasekhar, Philip Martinkus, and Matthew Christie. Magellan: toward building ecosystems of entity matching solutions. Communications of the ACM, 63(8):83–91, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3405476.

Papakonstantinou:2020:SLA

[184] Yannis Papakonstantinou. Supporting linear algebra operations in SQL: technical perspective. Communications of the ACM, 63(8):92, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3405468.

Luo:2020:SLA

[185] Shangyu Luo, Zekai J. Gao, Michael Gubanov, Luis L. Perez, Dimitrije Jankov, and Christopher Jermaine. Scalable linear algebra on a relational database system. Communications of the ACM, 63(8):93–101, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3405470.

Hoffmann:2020:LBS

[186] Leah Hoffmann. Last byte: Seeing light at the end of the cybersecurity tunnel. Communications of the ACM, 63(8): 104–ff, July 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/abs/10.1145/3403976.

Vardi:2020:DWA

[187] Moshe Y. Vardi. Departments: Where have all the domestic graduate students gone? Communications of the ACM, 63(9):5, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3410470.

Staff:2020:LEL

[188] CACM Staff. Letters to the editor: Lost in translation. Communications of the ACM, 63(9):9–10, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3411279.

Hazzan:2020:TCU

[189] Orit Hazzan. Teaching CS undergrads online to work with others effectively. Communications of the ACM, 63(9):12–13, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3409780.

Mone:2020:NA

[190] Gregory Mone. News: It’s alive! Communications of the ACM, 63(9):15–17, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3409800.

Greengard:2020:NAE

[191] Samuel Greengard. News: AI on edge. Communications of the ACM, 63(9):18–20, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3409977.

Marks:2020:NVC

[192] Paul Marks. News: Virtual collaboration in the age of the coronavirus. Communications of the ACM, 63(9):21–23, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3409803.

Kaminski:2020:LTR

[193] Margot Kaminski. Law and technology: A recent renaissance in privacy law. Communications of the ACM, 63(9):24–27, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3411049.

Lala:2020:SAV

[194] Jaynarayan H. Lala, Carl E. Landwehr, and John F. Meyer. Security: Autonomous vehicle safety: lessons from aviation. Communications of the ACM, 63(9):28–31, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3411053.

Denning:2020:PIA

[195] Peter J. Denning. The profession of IT: Avalanches make us all innovators. Communications of the ACM, 63(9):32–34, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3411055.

Schryen:2020:VIM

[196] Guido Schryen. Viewpoint: Integrating management science into the HPC research ecosystem. Communications of the ACM, 63(9):35–37, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3377477.

Bruckman:2020:VYT

[197] Amy Bruckman. Viewpoint: ’have you thought about… ’: talking about ethical implications of research. Communications of the ACM, 63(9):38–40, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3377405.

Narayanan:2020:DPP

[198] Arvind Narayanan, Arunesh Mathur, Marshini Chetty, and Mihir Kshirsagar. Dark patterns: past, present, and future. Communications of the ACM, 63(9):42–47, August 2020. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3397884.

Kelly:2020:PMP

[199] Terence Kelly. Is persistent memory persistent? Communications of the ACM, 63(9):48–54, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3397882.

Burke:2020:BAE

[200] Quinn Burke and Cinamon Sunrise Bailey. Becoming an ‘adaptive’ expert. Communications of the ACM, 63(9):56–64, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3366171.

Burton-Jones:2020:ISA

[201] Andrew Burton-Jones, Alicia Gilchrist, Peter Green, and Michael Draheim. Improving social alignment during digital transformation. Communications of the ACM, 63(9):65–71, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3410429.

Hellerstein:2020:KCW

[202] Joseph M. Hellerstein and Peter Alvaro. Keeping CALM: when distributed consistency is easy. Communications of the ACM, 63(9):72–81, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3369736.

Shahabi:2020:CVL

[203] Cyrus Shahabi. Computing the value of location data: technical perspective. Communications of the ACM, 63(9):84, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3410385.

Aly:2020:CVS

[204] Heba Aly, John Krumm, Gireeja Ranade, and Eric Horvitz. Computing value of spatiotemporal information. Communications of the ACM, 63(9):85–92, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3410387.

Shekhar:2020:PSC

[205] Shashi Shekhar. Progress in spatial computing for flood prediction: technical perspective. Communications of the ACM, 63(9):93, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3410410.

Lowe:2020:FRA

[206] Aaron Lowe, Pankaj K. Agarwal, and Mathias Rav. Flood-risk analysis on terrains. Communications of the ACM, 63(9):94–102, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3410413.

Clegg:2020:LBL

[207] Brian Clegg. Last byte: Little green message: a different kind of first-contact scenario. Communications of the ACM, 63(9):104–ff, August 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3409981.

Cerf:2020:CDD

[208] Vinton G. Cerf. Cerf’s up: On digital diplomacy. Communications of the ACM, 63(10):5, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3418557.

Hill:2020:PCP

[209] Robin K. Hill. Protecting computers and people from viruses. Communications of the ACM, 63(10):8, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3415748.

Monroe:2020:NBB

[210] Don Monroe. News: Bouncing balls and quantum computing. Communications of the ACM, 63(10):10–12, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3416076.

Edwards:2020:NTS

[211] Chris Edwards. News: Thwarting side-channel attacks. Communications of the ACM, 63(10):13–14, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416080.

Kirkpatrick:2020:NWA

[212] Keith Kirkpatrick. News: Who has access to your smartphone data? Communications of the ACM, 63(10):15–17, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3416078.

Garfinkel:2020:NFA

[213] Simson Garfinkel and Eugene H. Spafford. News: Fran Allen: 1932– 2020. Communications of the ACM, 63 (10):18–19, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3418560.

Cusumano:2020:TSMb

[214] Michael A. Cusumano. Technology strategy and management: Selfdriving vehicle technology: progress and promises. Communications of the ACM, 63(10):20–22, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3417074.

Neumann:2020:IRH

[215] Peter G. Neumann. Inside risks: A holistic view of future risks. Communications of the ACM, 63(10):23–27, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3417095.

Neville-Neil:2020:KVS

[216] George V. Neville-Neil. Kode vicious: Sanity vs. invisible markings. Communications of the ACM, 63(10): 28–29, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3417099.

Snodgrass:2020:VWN

[217] Richard T. Snodgrass and Marianne Winslett. Viewpoint: We need to automate the declaration of conflicts of interest. Communications of the ACM, 63(10):30–32, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3414556.

MacCormick:2020:VUC

[218] John MacCormick. Viewpoint: Using computer programs and search problems for teaching theory of computation. Communications of the ACM, 63 (10):33–35, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3382036.

Mencer:2020:HSF

[219] Oskar Mencer, Dennis Allison, Elad Blatt, Mark Cummings, Michael J. Flynn, Jerry Harris, Carl Hewitt, Quinn Jacobson, Maysam Lavasani, Mohsen Moazami, Hal Murray, Masoud Nikravesh, Andreas Nowatzyk, Mark Shand, and Shahram Shirazi. The history, status, and future of FP-GAs. Communications of the ACM, 63 (10):36–39, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3410669.

Chan:2020:DIG

[220] Charisma Chan and Beth Cooper. Debugging incidents in Google’s distributed systems. Communications of the ACM, 63(10):40–46, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3397880.

Galup:2020:WDA

[221] Stuart Galup, Ronald Dattero, and Jing Quan. What do agile, lean, and ITIL mean to DevOps? Communications of the ACM, 63(10):48–53, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3372114.

Wilhelm:2020:RTS

[222] Reinhard Wilhelm. Real time spent on real time. Communications of the ACM, 63(10):54–60, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3375545.

Bohme:2020:RVD

[223] Rainer Böhme, Lisa Eckey, Tyler Moore, Neha Narula, Tim Ruffing, and Aviv Zohar. Responsible vulnerability disclosure in cryptocurrencies. Communications of the ACM, 63 (10):62–71, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3372115.

Cresci:2020:DSB

[224] Stefano Cresci. A decade of social bot detection. Communications of the ACM, 63(10):72–83, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3409116.

Livshits:2020:ASC

[225] Benjamin Livshits. Analyzing smart contracts with MadMax: technical perspective. Communications of the ACM, 63(10):86, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3416259.

Grech:2020:MAG

[226] Neville Grech, Michael Kong, Anton Jurisevic, Lexi Brent, Bernhard Scholz, and Yannis Smaragdakis. Mad-Max: analyzing the out-of-gas world of smart contracts. Communications of the ACM, 63(10):87–95, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416262.

Beame:2020:TPO

[227] Paul Beame. Two for the price of one: technical perspective. Communications of the ACM, 63(10):96, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416264.

Farhadi:2020:LBE

[228] Alireza Farhadi, Mohammad Taghi Hajiaghayi, Kasper Green Larsen, and Elaine Shi. Lower bounds for external memory integer sorting via network coding. Communications of the ACM, 63(10):97–105, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416268.

Shasha:2020:LBP

[229] Dennis Shasha. Last byte: Privacypreserving polling. Communications of the ACM, 63(10):108–ff, September 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416266.

Rohlfing:2020:CPC

[230] Celeste M. Rohlfing. Career paths in computing: A career unfolds in phases. Communications of the ACM, 63(11):7, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3422082.

Staff:2020:LEW

[231] CACM Staff. Letters to the editor: Weighing grad school payback. Communications of the ACM, 63(11):8–9, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3425745.

Patterson:2020:BBI

[232] David Patterson and Yegor Bugayenko. BLOG@CACM: Bringing industry back to conferences, and paying for results. Communications of the ACM, 63(11):12–13, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3422628.

Monroe:2020:NSA

[233] Don Monroe. News: Seeking artificial common sense. Communications of the ACM, 63(11):14–16, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422588.

Kirkpatrick:2020:NNL

[234] Keith Kirkpatrick. News: Natural language misunderstanding. Communications of the ACM, 63(11):17–18, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422600.

Greengard:2020:NTN

[235] Samuel Greengard. News: Terahertz networks move closer to reality. Communications of the ACM, 63 (11):19–21, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3422630.

Cranor:2020:PDC

[236] Lorrie Faith Cranor. Privacy: Digital contact tracing may protect privacy, but it is unlikely to stop the pandemic. Communications of the ACM, 63(11):22–24, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3423993.

Samuelson:2020:LSC

[237] Pamela Samuelson. Legally speaking: Copyright’s online service providers safe harbors under siege. Communications of the ACM, 63(11):25–27, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3423995.

VanAlstyne:2020:EBDb

[238] Marshall W. Van Alstyne and Alisa Lenart. Economic and business dimensions: Using data and respecting users. Communications of the ACM, 63(11):28–30, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3423998.

Ko:2020:EIT

[239] Amy J. Ko, Alannah Oleson, Neil Ryan, Yim Register, Benjamin Xie, Mina Tari, Matthew J. Davidson, Stefania Druga, and Dastyni Loksa. Education: It is time for more critical CS education. Communications of the ACM, 63(11):31–33, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3424000.

Ronen:2020:VWS

[240] Boaz Ronen and Alex Coman. Viewpoint: Where should your IT constraint be?: The case of the financial services industry. Communications of the ACM, 63(11):34–37, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3382081.

Visser:2020:VRC

[241] Jacky Visser, John Lawrence, and Chris Reed. Viewpoint: Reasonchecking fake news. Communications of the ACM, 63(11):38–40, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3397189.

Almeida:2020:LAR

[242] Virgilio Almeida, Gonzalo Navarro, and Sergio Rajsbaum. Latin America regional special section: Welcome. Communications of the ACM, 63(11):42–44, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3416955.

DeFreitas:2020:LAR

[243] Rosiane De Freitas, João M. B. Cavalcanti, Sergio Cleger, Niro Higuchi, Carlos Henrique Celes, and Adriano Lima. Latin America regional special section: Hot topics: Estimating Amazon carbon stock using AI-based remote sensing. Communications of the ACM, 63 (11):46–48, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3416957.

Hevia:2020:LAR

[244] Alejandro Hevia and Camilo Gómez. Latin America regional special section: Hot topics: Why me?: shedding light on random processes via randomness beacons. Communications of the ACM, 63(11):49–50, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3416959.

Kon:2020:LAR

[245] Fabio Kon, Kelly Braghetto, Eduardo Z. Santana, Roberto Speicys, and Jorge Guerra Guerra. Latin America regional special section: Hot topics: Toward smart and sustainable cities. Communications of the ACM, 63(11):51–52, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3416961.

Milano:2020:LAR

[246] Gastón Milano, Diego Vallespir, and Alfredo Viola. Latin America regional special section: Hot topics: A technological and innovative approach to COVID-19 in Uruguay. Communications of the ACM, 63(11):53–55, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422826.

Meira:2020:LAR

[247] Wagner Meira, Antonio L. P. Ribeiro, Derick M. Oliveira, and Antonio H. Ribeiro. Latin America regional special section: Hot topics: Contextualized interpretable machine learning for medical diagnosis. Communications of the ACM, 63(11):56–58, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416965.

Arce-Lopera:2020:LAR

[248] Carlos Arce-Lopera and Gerardo M. Sarria M. Latin America regional special section: Hot topics: Understanding salsa: how computing is defining Latin music. Communications of the ACM, 63(11):59–60, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3416967.

Poblete:2020:LAR

[249] Barbara Poblete and Jorge Pérez. Latin America regional special section: Hot topics: Minding the AI gap in LATAM. Communications of the ACM, 63(11):61–63, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3416969.

Arroyuelo:2020:LAR

[250] Diego Arroyuelo, José Fuentes-Sepülveda, and Diego Seco. Latin America regional special section: Hot topics: Three success stories about compact data structures. Communications of the ACM, 63(11):64–65, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3416971.

Gitler:2020:LAR

[251] Isidoro Gitler, Antônio Tadeu A. Gomes, and Sergio Nesmachnow. Latin America regional special section: Big trends: The Latin American supercomputing ecosystem for science. Communications of the ACM, 63(11):66– 71, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3419977.

Tentori:2020:LAR

[252] Monica Tentori, Artur Ziviani, Débora C. Muchaluat-Saade, and Jesus Favela. Latin America regional special section: Big trends: Digital healthcare in Latin America: the case of Brazil and Mexico. Communications of the ACM, 63 (11):72–77, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3423923.

Arenas:2020:LAR

[253] Marcelo Arenas and Pablo Barceló. Latin America regional special section: Big trends: Chile’s new interdisciplinary institute for foundational research on data. Communications of the ACM, 63(11):78–83, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3416975.

Zelaya:2020:LAR

[254] Gerardo Torres Zelaya. Latin America regional special section: Big trends: A panorama of computing in central America and the Caribbean. Communications of the ACM, 63(11):84– 89, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3419979.

Delrieux:2020:LAR

[255] Claudio Delrieux, Virginia Ballarín, Cristian García Bauza, and Mario A. López. Latin America regional special section: Big trends: Imaging sciences R&D laboratories in Argentina. Communications of the ACM, 63(11):90– 95, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3416977.

Duarte:2020:LAR

[256] Elias P. Duarte, Raimundo J. A. Macêdo, Eliane Martins, and Sergio Rajsbaum. Latin America regional special section: Big trends: A tour of dependable computing research in Latin America. Communications of the ACM, 63(11):96–101, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3416979.

Kiwi:2020:LAR

[257] Marcos Kiwi, Yoshiharu Kohayakawa, Sergio Rajsbaum, Francisco Jose Rodríguez-Henríquez, Jayme Luiz Szwarcfiter, and Alfredo Viola. Latin America regional special section: Big trends: A perspective on theoretical computer science in Latin America. Communications of the ACM, 63 (11):102–107, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3419975.

Limoncelli:2020:FNR

[258] Thomas A. Limoncelli. Five nonobvious remote work techniques. Communications of the ACM, 63(11):108– 110, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3410627.

Helland:2020:DOV

[259] Pat Helland. Data on the outside versus data on the inside. Communications of the ACM, 63(11):111–118, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3410623.

Resnick:2020:CC

[260] Mitchel Resnick and Natalie Rusk. Coding at a crossroads. Communications of the ACM, 63(11):120–127, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3375546.

Grohe:2020:GIP

[261] Martin Grohe and Pascal Schweitzer. The graph isomorphism problem. Communications of the ACM, 63(11):128– 134, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3372123.

Efros:2020:TPW

[262] Alexei A. Efros and Aaron Hertzmann. Technical perspective: When the adversary is your friend. Communications of the ACM, 63(11):138, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422602.

Goodfellow:2020:GAN

[263] Ian Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair Aaron C. Courville, and Yoshua Bengio. Generative adversarial networks. Communications of the ACM, 63(11):139–144, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422622.

Xu:2020:TPB

[264] Harry Xu. Technical perspective: BLeak: semantics-aware leak detection in the web. Communications of the ACM, 63(11):145, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422590.

Vilk:2020:BAD

[265] John Vilk and Emery D. Berger. BLeak: automatically debugging memory leaks in web applications. Communications of the ACM, 63(11):146– 153, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3422598.

Hoffmann:2020:LBT

[266] Leah Hoffmann. Last byte: Tackling the challenges of CS education. Communications of the ACM, 63(11):160– ff, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3422672.

Cerf:2020:CRR

[267] Vinton G. Cerf. Cerf’s up: Repairability redux. Communications of the ACM, 63(12):5, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3429267.

Newborn:2020:PCA

[268] Monroe “Monty” Newborn. Pitting computers against each other … in chess. Communications of the ACM, 63(12):6–7, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3427786.

Savage:2020:NTC

[269] Neil Savage. News: Tracking COVID, discreetly. Communications of the ACM, 63(12):9–11, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3427933.

Shein:2020:NSR

[270] Esther Shein. News: Softening up robots. Communications of the ACM, 63(12):12–14, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3427942.

Kugler:2020:NTV

[271] Logan Kugler. News: Technologies for the visually impaired. Communications of the ACM, 63(12):15–17, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3427936.

Canca:2020:CEO

[272] Cansu Canca. Computing ethics: Operationalizing AI ethics principles. Communications of the ACM, 63(12): 18–21, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3430368.

Adrion:2020:BPU

[273] W. Richards Adrion, Sarah T. Dunton, Barbara Ericson, Renee Fall, Carol Fletcher, and Mark Guzdial. Broadening participation: U.S. states must broaden participation while expanding access to computer science education. Communications of the ACM, 63 (12):22–25, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3430375.

Denning:2020:PIN

[274] Peter J. Denning. The profession of IT: Navigating in real-time environments. Communications of the ACM, 63(12):26–28, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3430370.

Neville-Neil:2020:KVR

[275] George V. Neville-Neil. Kode vicious: Removing kode. Communications of the ACM, 63(12):29, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3430378.

OMara:2020:VSP

[276] Margaret O’Mara. Viewpoint: Silicon politics. Communications of the ACM, 63(12):30–32, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3431480.

Cheng:2020:VFL

[277] Yong Cheng, Yang Liu, Tianjian Chen, and Qiang Yang. Viewpoint: Federated learning for privacy-preserving AI. Communications of the ACM, 63 (12):33–36, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3387107.

Frazelle:2020:LDB

[278] Jessie Frazelle. The life of a data byte. Communications of the ACM, 63 (12):38–45, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3410625.

Jover:2020:SAS

[279] Roger Piqueras Jover. Security analysis of SMS as a second factor of authentication. Communications of the ACM, 63(12):46–52, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3424260.

Schwartz:2020:GA

[280] Roy Schwartz, Jesse Dodge, Noah A. Smith, and Oren Etzioni. Green AI. Communications of the ACM, 63 (12):54–63, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3381831.

Maasberg:2020:DTI

[281] Michele Maasberg, Craig Van Slyke, Selwyn Ellis, and Nicole Beebe. The dark triad and insider threats in cyber security. Communications of the ACM, 63(12):64–80, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3408864.

Feamster:2020:MIS

[282] Nick Feamster and Jason Livingood. Measuring Internet speed: current challenges and future recommendations. Communications of the ACM, 63 (12):72–80, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3372135.

Forsyth:2020:XNP

[283] David Alexander Forsyth. XNORnetworks — powerful but tricky: technical perspective. Communications of the ACM, 63(12):82, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3429950.

Rastegari:2020:EAE

[284] Mohammad Rastegari, Vicente Ordonez, Joseph Redmon, and Ali Farhadi. Enabling AI at the edge with XNOR-networks. Communications of the ACM, 63(12):83–90, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3429945.

Paradiso:2020:FLS

[285] Joseph A. Paradiso. The future of large-scale embedded sensing: technical perspective. Communications of the ACM, 63(12):91, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3429952.

Arora:2020:SII

[286] Nivedita Arora, Thad Starner, and Gregory D. Abowd. SATURN: an introduction to the Internet of Materials. Communications of the ACM, 63 (12):92–99, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3429948.

Bainbridge:2020:LBW

[287] William Sims Bainbridge. Last byte: Walden Three. Communications of the ACM, 63(12):104–ff, December 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3427948.

Chien:2021:ELC

[288] Andrew A. Chien. Editor’s letter: 2021: Computing’s divided future. Communications of the ACM, 64(1): 5, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3439708.

Vardi:2021:DRC

[289] Moshe Y. Vardi. Departments: Reboot the computing-research publication systems. Communications of the ACM, 64(1):7, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3437991.

Crawford:2021:CPC

[290] Dona Crawford. Career paths in computing: A career fueled by HPC. Communications of the ACM, 64(1):9, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3434141.

Guzdial:2021:TAR

[291] Mark Guzdial. Talking about race in CS education. Communications of the ACM, 64(1):10–11, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3433921.

Greengard:2021:NGD

[292] Samuel Greengard. News: Geometric deep learning advances data science. Communications of the ACM, 64(1):13–15, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3433951.

Monroe:2021:NFT

[293] Don Monroe. News: Fugaku takes the lead. Communications of the ACM, 64(1):16–18, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3433954.

Edwards:2021:NCW

[294] Chris Edwards. News: Coalition of the willing takes aim at COVID-19. Communications of the ACM, 64(1):19–21, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3433952.

Cusumano:2021:TSMa

[295] Michael A. Cusumano. Technology strategy and management: Boeing’s 737 MAX: a failure of management, not just technology. Communications of the ACM, 64(1):22–25, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3436231.

Benzel:2021:SCR

[296] Terry Benzel. Security: Cybersecurity research for the future. Communications of the ACM, 64(1):26–28, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3436241.

Klonick:2021:LTC

[297] Kate Klonick. Law and technology: Content moderation modulation. Communications of the ACM, 64(1):29–31, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3436247.

Haigh:2021:HRI

[298] Thomas Haigh. Historical reflections: The immortal soul of an old machine. Communications of the ACM, 64(1):32–37, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3436249.

Marcus:2021:VIA

[299] Gary Marcus and Ernest Davis. Viewpoint: Insights for AI from the human mind. Communications of the ACM, 64(1):38–41, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3392663.

Turel:2021:VEU

[300] Ofir Turel and Christopher Ferguson. Viewpoint: Excessive use of technology: can tech providers be the culprits? Communications of the ACM, 64(1):42–44, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3392664.

Vachon:2021:IEP

[301] Phil Vachon. The identity in everyone’s pocket. Communications of the ACM, 64(1):46–55, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3424262.

Levine:2021:C

[302] Edlyn V. Levine. The die is cast. Communications of the ACM, 64(1):56–60, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3424264.

Cabanas:2021:DFU

[303] José González Cabañas, Ángel Cuevas, Aritz Arrate, and Rubén Cuevas. Does Facebook use sensitive data for advertising purposes? Communications of the ACM, 64(1):62–69, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3426361.

Koutroumpis:2021:DII

[304] Pantelis Koutroumpis, Aija Leiponen, and Llewellyn D. W. Thomas. Digital instruments as invention machines. Communications of the ACM, 64(1):70–78, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3377476.

Bozan:2021:HTI

[305] Karoly Bozan, Kalle Lyytinen, and Gregory M. Rose. How to transition incrementally to microservice architecture. Communications of the ACM, 64(1):79–85, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3378064.

Lindell:2021:SMC

[306] Yehuda Lindell. Secure multiparty computation. Communications of the ACM, 64(1):86–96, January 2021. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3387108.

Wicker:2021:EZD

[307] Stephen B. Wicker. The ethics of zeroday exploits: the NSA meets the trolley car. Communications of the ACM, 64(1):97–103, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3393670.

Gottesman:2021:TPD

[308] Daniel Gottesman. Technical perspective: Deciphering errors to reduce the cost of quantum computation. Communications of the ACM, 64(1):105, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3434157.

Fawzi:2021:COQ

[309] Omar Fawzi, Antoine Grospellier, and Anthony Leverrier. Constant overhead quantum fault tolerance with quantum expander codes. Communications of the ACM, 64(1):106–114, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3434163.

Han:2021:TPS

[310] Richard Han. Technical perspective: SkyCore’s architecture takes it to the ‘edge’. Communications of the ACM, 64(1):115, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3434151.

Moradi:2021:SMC

[311] Mehrdad Moradi, Karthikeyan Sundaresan, Eugene Chai, Sampath Rangarajan, and Z. Morley Mao. SkyCore: moving core to the edge for untethered and reliable UAV-based LTE networks. Communications of the ACM, 64(1):116–124, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3434161.

Shasha:2021:LBSa

[312] Dennis Shasha. Last byte: Stay in balance. Communications of the ACM, 64(1):128, January 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3434645.

Chien:2021:ELD

[313] Andrew A. Chien. Editor’s letter: Driving the cloud to true zero carbon. Communications of the ACM, 64(2):5, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3445037.

Cerf:2021:CHB

[314] Vinton G. Cerf. Cerf’s up: Half-baked high-resolution referencing. Communications of the ACM, 64(2):7, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442186.

Staff:2021:LES

[315] CACM Staff. Letters to the editor: Salary disputes. Communications of the ACM, 64(2):8–9, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442383.

Hill:2021:IAW

[316] Robin K. Hill. Issues arise when time goes digital. Communications of the ACM, 64(2):10–11, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3440990.

Edwards:2021:NML

[317] Chris Edwards. News: Moore’s Law: what comes next? Communications of the ACM, 64(2):12–14, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3440992.

Kugler:2021:NSV

[318] Logan Kugler. News: The state of virtual reality hardware. Communications of the ACM, 64(2):15–16, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3441290.

Kirkpatrick:2021:NTR

[319] Keith Kirkpatrick. News: Technological responses to COVID-19. Communications of the ACM, 64(2):17–19, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3441288.

Bakos:2021:EBD

[320] Yannis Bakos, Hanna Halaburda, and Christoph Mueller-Bloch. Economic and business dimensions: When permissioned blockchains deliver more decentralization than permissionless. Communications of the ACM, 64(2): 20–22, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3442371.

Fletcher:2021:ECF

[321] Carol L. Fletcher and Jayce R. Warner. Education: CAPE: a framework for assessing equity throughout the computer science education ecosystem. Communications of the ACM, 64(2):23–25, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3442373.

Neville-Neil:2021:KVK

[322] George V. Neville-Neil. Kode Vicious: Kabin fever. Communications of the ACM, 64(2):26–27, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442375.

Maurer:2021:VCI

[323] Chris Maurer, Kevin Kim, Dan Kim, and Leon A. Kappelman. Viewpoint: Cybersecurity: is it worse than we think? Communications of the ACM, 64(2):28–30, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3399667.

Kambhampati:2021:VPR

[324] Subbarao Kambhampati. Viewpoint: Polanyi’s revenge and AI’s new romance with tacit knowledge. Communications of the ACM, 64(2):31–32, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3446369.

Knuth:2021:VLD

[325] Donald E. Knuth and Len Shustek. Viewpoint: Let’s not dumb down the history of computer science. Communications of the ACM, 64(2):33–35, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3442377.

Staff:2021:DPP

[326] CACM Staff. Differential privacy: the pursuit of protections by default. Communications of the ACM, 64(2):36–43, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3434228.

Limoncelli:2021:TSB

[327] Thomas A. Limoncelli. The time I stole $10,000 from Bell Labs. Communications of the ACM, 64(2):44–46, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3434224.

Feit:2021:AAC

[328] Anna Maria Feit, Mathieu Nancel, Maximilian John, Andreas Karrenbauer, Daryl Weir, and Antti Oulasvirta. AZERTY amélioré: computational design on a national scale. Communications of the ACM, 64(2): 48–58, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3382035.

Shastri:2021:GAP

[329] Supreeth Shastri, Melissa Wasserman, and Vijay Chidambaram. GDPR anti-patterns. Communications of the ACM, 64(2):59–65, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3378061.

Ahn:2021:KSK

[330] Dong H. Ahn, Allison H. Baker, Michael Bentley, Ian Briggs, Ganesh Gopalakrishnan, Dorit M. Hammerling, Ignacio Laguna, Gregory L. Lee, Daniel J. Milroy, and Mariana Vertenstein. Keeping science on keel when software moves. Communications of the ACM, 64(2):66–74, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3382037.

Hitzler:2021:RSW

[331] Pascal Hitzler. A review of the Semantic Web field. Communications of the ACM, 64(2):76–83, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3397512.

Wagh:2021:DCM

[332] Sameer Wagh, Xi He, Ashwin Machanavajjhala, and Prateek Mittal. DP-cryptography: marrying differential privacy and cryptography in emerging applications. Communications of the ACM, 64(2):84–93, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3418290.

Amin:2021:PME

[333] Nada Amin. Programming microfluidics to execute biological protocols: technical perspective. Communications of the ACM, 64(2):96, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3441684.

Ott:2021:BPS

[334] Jason Ott, Tyson Loveless, Chris Curtis, Mohsen Lesani, and Philip Brisk. BioScript: programming safe chemistry on laboratories-on-a-chip. Communications of the ACM, 64(2):97–104, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3441686.

Ives:2021:SSR

[335] Zachary G. Ives. Solving the signal reconstruction problem at scale: technical perspective. Communications of the ACM, 64(2):105, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3441688.

Asudeh:2021:SSR

[336] Abolfazl Asudeh, Jees Augustine, Saravanan Thirumuruganathan, Azade Nazi, Nan Zhang, Gautam Das, and Divesh Srivastava. Scalable signal reconstruction for a broad range of applications. Communications of the ACM, 64 (2):106–115, February 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3441689.

Hoffmann:2021:LBB

[337] Leah Hoffmann. Last byte: Bringing stability to wireless connections. Communications of the ACM, 64(2):120–ff, February 2021. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3441293.

Chien:2021:ELA

[338] Andrew A. Chien. Editor’s letter: Around the world: (the first time) with Communications’ Regional Special Sections. Communications of the ACM, 64(3):5–7, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3448648.

Vardi:2021:DPVa

[339] Moshe Y. Vardi. Departments: The people vs. tech. Communications of the ACM, 64(3):9, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3446769.

Oakes:2021:CPC

[340] Graham Oakes. Career paths in computing: Enabling renewable energy through smarter grids. Communications of the ACM, 64(3):11, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3441558.

Guzdial:2021:DDB

[341] Mark Guzdial and Joel C. Adams. Disputing Dijkstra, and birthdays in base 2. Communications of the ACM, 64(3):12–13, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3446806.

Monroe:2021:NPQ

[342] Don Monroe. News: The power of quantum complexity. Communications of the ACM, 64(3):15–17, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3446875.

Savage:2021:NFF

[343] Neil Savage. News: Fact-finding mission. Communications of the ACM, 64(3):18–19, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3446879.

Marks:2021:NCB

[344] Paul Marks. News: Can the biases in facial recognition be fixed; also, should they? Communications of the ACM, 64(3):20–22, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3446877.

Garfinkel:2021:NEM

[345] Simson Garfinkel and Eugene H. Spafford. News: Edmund M. Clarke (1945– 2020). Communications of the ACM, 64(3):23–24, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3447810.

Samuelson:2021:LSP

[346] Pamela Samuelson. Legally speaking: The push for stricter rules for Internet platforms. Communications of the ACM, 64(3):26–28, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447251.

Cranor:2021:PIC

[347] Lorrie Faith Cranor. Privacy: Informing California privacy regulations with evidence from research. Communications of the ACM, 64(3):29–32, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447253.

Johnson:2021:CEW

[348] Deborah G. Johnson and Nicholas Diakopoulos. Computing ethics: What to do about deepfakes. Communications of the ACM, 64(3):33–35, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447255.

Denning:2021:PIS

[349] Peter J. Denning and Jeffrey Johnson. The profession of IT: Science is not another opinion. Communications of the ACM, 64(3):36–38, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447257.

Wirth:2021:VYP

[350] Niklaus Wirth. Viewpoint: 50 years of Pascal. Communications of the ACM, 64(3):39–41, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447525.

Friesike:2021:VWC

[351] Sascha Friesike, Frédéric Thiesse, and George Kuk. Viewpoint: What can the maker movement teach us about the digitization of creativity? Communications of the ACM, 64(3):42–45, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447524.

Nov:2021:VTP

[352] Oded Nov, Yindalon Aphinyanaphongs, Yvonne W. Lui, Devin Mann, Maurizio Porfiri, Mark Riedl, John-Ross Rizzo, and Batia Wiesenfeld. Viewpoint: The transformation of patient-clinician relationships with AI-based medical advice. Communications of the ACM, 64(3):46–48, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3417518.

Staff:2021:SCW

[353] CACM Staff. A second conversation with Werner Vogels. Communications of the ACM, 64(3):50–57, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3434232.

Frazelle:2021:WAM

[354] Jessie Frazelle. Out-of-this-world additive manufacturing. Communications of the ACM, 64(3):58–62, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3424254.

Thompson:2021:DCG

[355] Neil C. Thompson and Svenja Spanuth. The decline of computers as a general purpose technology. Communications of the ACM, 64(3):64–72, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3430936.

Molnar:2021:EIF

[356] Andreea Molnar, Therese Keane, and Rosemary Stockdale. Educational interventions and female enrollment in IT degrees. Communications of the ACM, 64(3):73–77, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3387106.

Wang:2021:GTC

[357] Lucy Lu Wang, Gabriel Stanovsky, Luca Weihs, and Oren Etzioni. Gender trends in computer science authorship. Communications of the ACM, 64(3):78–84, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3430803.

Mazurczyk:2021:CRT

[358] Wojciech Mazurczyk and Luca Caviglione. Cyber reconnaissance techniques. Communications of the ACM, 64(3):86–95, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3418293.

Gutierrez:2021:KG

[359] Claudio Gutierrez and Juan F. Sequeda. Knowledge graphs. Communications of the ACM, 64(3):96–104, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3418294.

Arora:2021:WDT

[360] Sanjeev Arora. Why don’t today’s deep nets overfit to their training data?: Technical perspective. Communications of the ACM, 64(3):106, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3446773.

Zhang:2021:UDL

[361] Chiyuan Zhang, Samy Bengio, Moritz Hardt, Benjamin Recht, and Oriol Vinyals. Understanding deep learning (still) requires rethinking generalization. Communications of the ACM, 64(3):107–115, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3446776.

Dutta:2021:LIO

[362] Prabal Dutta. Localizing insects outdoors: technical perspective. Communications of the ACM, 64(3):116, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3446778.

Nandakumar:2021:LSS

[363] Rajalakshmi Nandakumar, Vikram Iyer, and Shyamnath Gollakota. 3D localization for subcentimeter-sized devices. Communications of the ACM, 64(3):117–125, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3446782.

Clegg:2021:LBAa

[364] Brian Clegg. Last byte: Awakening. Communications of the ACM, 64(3):128–ff, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3446803.

Cerf:2021:CWD

[365] Vinton G. Cerf. Cerf’s up: What does a static, sustainable economy look like? Communications of the ACM, 64(4):5, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3450610.

Arquilla:2021:BSH

[366] John Arquilla and Mark Guzdial. BLOG@CACM: The SolarWinds hack, and a grand challenge for CS education. Communications of the ACM, 64(4): 6–7, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3449047.

Edwards:2021:NBN

[367] Chris Edwards. News: The best of NLP. Communications of the ACM, 64(4):9–11, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3449049.

Marks:2021:NDL

[368] Paul Marks. News: Deep learning speeds MRI scans. Communications of the ACM, 64(4):12–14, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3449060.

Greengard:2021:NWS

[369] Samuel Greengard. News: The worsening state of ransomware. Communications of the ACM, 64(4):15–17, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3449054.

Sako:2021:TSM

[370] Mari Sako. Technology strategy and management: From remote work to working from anywhere. Communications of the ACM, 64(4):20–22, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3451223.

Brown:2021:BPR

[371] Q. Brown, T. Grandison, J. D. Burge, O. C. Jenkins, and T. Dillahunt. Broadening participation: Reflections on black in computing. Communications of the ACM, 64(4):23–24, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3451225.

Neville-Neil:2021:KVN

[372] George V. Neville-Neil. Kode Vicious: The non-psychopath’s guide to managing an open source project. Communications of the ACM, 64(4):25–27, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3451229.

Haigh:2021:HRWa

[373] Thomas Haigh. Historical reflections: When hackers were heroes. Communications of the ACM, 64(4):28–34, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3451227.

DeMol:2021:VRP

[374] Liesbeth De Mol and Maarten Bullynck. Viewpoint: Roots of ‘program’ revisited. Communications of the ACM, 64(4):35–37, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3419406.

Vrandecic:2021:VBM

[375] Denny Vrandecić. Viewpoint: Building a multilingual Wikipedia. Communications of the ACM, 64(4):38–41, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3425778.

Aly:2021:AWS

[376] Sherif G. Aly, Mohamed Mokbel, and Moustafa Youssef. Arab world special section: Welcome. Communications of the ACM, 64(4):42–44, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447754.

Elnozahy:2021:AWS

[377] E. N. (Mootaz) Elnozahy. Arab world special section: Hot topics: Building a research university in the Arab region: the case of KAUST. Communications of the ACM, 64(4):46–49, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3447717.

Elmagarmid:2021:AWS

[378] Ahmed Elmagarmid and Abdellatif Saoudi. Arab world special section: Hot topics: Building a preeminent research lab in the Arab region: the case of QCRI. Communications of the ACM, 64(4):50–53, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3447719.

ElSaban:2021:AWS

[379] Motaz El Saban. Arab world special section: Hot topics: Data science for the oil and gas industry in the Arab region. Communications of the ACM, 64(4):54–56, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3447721.

Shamout:2021:AWS

[380] Farah E. Shamout and Dana Abu Ali. Arab world special section: Hot topics: The strategic pursuit of artificial intelligence in the United Arab Emirates. Communications of the ACM, 64(4):57–58, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3447723.

Aboulnaga:2021:AWSa

[381] Ashraf Aboulnaga, Sanjay Chawla, Ahmed Elmagarmid, Mohammed Al-Mannai, and Hassan Al-Sayed. Arab world special section: Hot topics: An AI-enabled future for Qatar and the region. Communications of the ACM, 64(4):59–61, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447725.

Touma:2021:AWS

[382] Walid R. Touma and Saad El Zein. Arab world special section: Hot topics: Entrepreneurship ecosystem in Lebanon. Communications of the ACM, 64(4):62–63, April 2021. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3447727.

Eraqi:2021:AWS

[383] Hesham M. Eraqi and Ibrahim Sobh. Arab world special section: Hot topics: Autonomous driving in the face of unconventional odds. Communications of the ACM, 64(4):64–66, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447729.

Abbar:2021:AWS

[384] Sofiane Abbar, Rade Stanojevic, Shadab Mustafa, and Mohamed Mokbel. Arab world special section: Hot topics: Traffic routing in the everchanging city of Doha. Communications of the ACM, 64(4):67–68, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3447731.

Lazem:2021:AWS

[385] Shaimaa Lazem, Mennatallah Saleh, and Ebtisam Alabdulqader. Arab world special section: Hot topics:

ArabHCI: five years and counting. Communications of the ACM, 64(4): 69–71, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3447733.

Darwish:2021:AWS

[386] Kareem Darwish, Nizar Habash, Mourad Abbas, Hend Al-Khalifa, Huseein T. Al-Natsheh, Houda Bouamor, Karim Bouzoubaa, Violetta Cavalli-Sforza, Samhaa R. El-Beltagy, Wassim El-Hajj, Mustafa Jarrar, and Hamdy Mubarak. Arab world special section: Big trends: A panoramic survey of natural language processing in the Arab world. Communications of the ACM, 64(4):72–81, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3447735.

Keyes:2021:AWS

[387] David Keyes. Arab world special section: Big trends: The Arab world prepares the exascale workforce. Communications of the ACM, 64(4):82–87, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3447737.

Weber:2021:AWS

[388] Ingmar Weber, Muhammad Imran, Ferda Ofli, Fouad Mrad, Jennifer Colville, Mehdi Fathallah, Alissar Chaker, and Wigdan Seed Ahmed. Arab world special section: Big trends: Non-traditional data sources: providing insights into sustainable development. Communications of the ACM, 64(4):88–95, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447739.

Popper:2021:AWS

[389] Christina Pöpper, Michail Maniatakos, and Roberto Di Pietro. Arab world special section: Big trends: Cyber security research in the Arab region: a blooming ecosystem with global ambitions. Communications of the ACM, 64(4):96–101, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447741.

Abdennadher:2021:AWS

[390] Slim Abdennadher, Sherif G. Aly, Joe Tekli, and Karima Echihabi. Arab world special section: Big trends: Unleashing early maturity academic innovations. Communications of the ACM, 64(4):102–107, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447743.

Eldawlatly:2021:AWS

[391] Seif Eldawlatly, Mohamed Abouelhoda, Omar S. Al-Kadi, Takashi Gojobori, Boris Jankovic, Mohamad Khalil, Ahsan H. Khandoker, and Ahmed Morsy. Arab world special section: Big trends: Biomedical computing in the Arab world: unlocking the potential of a growing research community. Communications of the ACM, 64(4):108–113, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447746.

Shihada:2021:AWS

[392] Basem Shihada, Tamer Elbatt, Ahmed Eltawil, Mohammad Mansour, Essaid Sabir, Slim Rekhis, and Sanaa Sharafeddine. Arab world special section: Big trends: Networking research for the Arab world: from regional initiatives to potential global impact. Communications of the ACM, 64(4):114–119, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3447748.

Aboulnaga:2021:AWSb

[393] Ashraf Aboulnaga, Azza Abouzied, Karima Echihabi, and Mourad Ouzzani. Arab world special section: Big trends: Database systems research in the Arab world: a tradition that spans decades. Communications of the ACM, 64(4):120–123, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3447750.

Ali:2021:AWS

[394] Ahmed Ali, Shammur Chowdhury, Mohamed Afify, Wassim El-Hajj, Hazem Hajj, Mourad Abbas, Amir Hussein, Nada Ghneim, Mohammad Abushariah, and Assal Alqudah. Arab world special section: Big trends: Connecting Arabs: bridging the gap in dialectal speech recognition. Communications of the ACM, 64(4):124–129, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3451150.

Crawshaw:2021:EVN

[395] David Crawshaw. Everything VPN is new again. Communications of the ACM, 64(4):130–134, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3434230.

Friedler:2021:IPF

[396] Sorelle A. Friedler, Carlos Scheidegger, and Suresh Venkatasubramanian. The (im)possibility of fairness: different value systems require different mechanisms for fair decision making. Communications of the ACM, 64(4):136–143, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3433949.

Jung:2021:SSP

[397] Ralf Jung, Jacques-Henri Jourdan, Robbert Krebbers, and Derek Dreyer. Safe systems programming in Rust. Communications of the ACM, 64 (4):144–152, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3418295.

SouzaDosReis:2021:TAC

[398] Eduardo Souza Dos Reis, Cristiano André Da Costa, Diórgenes Eugênio Da Silveira, Rodrigo Simon Bavaresco, Rodrigo Da Rosa Righi, Jorge Luis Victória Barbosa, Rodolfo Stoffel Antunes, Márcio Miguel Gomes, and Gustavo Federizzi. Transformers aftermath: current research and rising trends. Communications of the ACM, 64(4):154–163, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3430937.

Idreos:2021:SST

[399] Stratos Idreos. The strength of SuRF: technical perspective. Communications of the ACM, 64(4):165, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3450261.

Zhang:2021:SRF

[400] Huanchen Zhang, Hyeontaek Lim, Viktor Leis, David G. Andersen, Michael Kaminsky, Kimberly Keeton, and Andrew Pavlo. Succinct range filters. Communications of the ACM, 64(4):166–173, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3450262.

Shasha:2021:LBRa

[401] Dennis Shasha. Last byte: Roulette angel. Communications of the ACM, 64(4):176–ff, April 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3450331.

Vardi:2021:DAT

[402] Moshe Y. Vardi. Departments: The agency trilemma and ACM. Communications of the ACM, 64(5):4, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3459084.

Pancake:2021:ELR

[403] Cherri M. Pancake and Andrew A. Chien. Editor’s letter: In response to ’Vardi’s insights’. Communications of the ACM, 64(5):5, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3459982.

Mitchell:2021:CPC

[404] Caven Cade Mitchell. Career paths in computing: Computing enabled me to… promote sustainability and help underserved communities. Communications of the ACM, 64(5):7, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453631.

Guzdial:2021:TOT

[405] Mark Guzdial. Teaching other teachers how to teach CS better. Communications of the ACM, 64(5):8–9, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453647.

Monroe:2021:NSR

[406] Don Monroe. News: A satisfying result. Communications of the ACM, 64(5):10–12, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453650.

Savage:2021:NCF

[407] Neil Savage. News: Catching the fakes. Communications of the ACM, 64(5):13–14, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453696.

Kirkpatrick:2021:NTC

[408] Keith Kirkpatrick. News: A traffic cop for low earth orbit. Communications of the ACM, 64(5):15–17, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453703.

Peisert:2021:STS

[409] Sean Peisert. Security: Trustworthy scientific computing. Communications of the ACM, 64(5):18–21, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3457191.

Choi:2021:LTS

[410] Bryan H. Choi. Law and technology: Software professionals, malpractice law, and codes of ethics. Communications of the ACM, 64(5):22–24, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3457193.

Bell:2021:ECU

[411] Tim Bell. Education: CS unplugged or coding classes? Communications of the ACM, 64(5):25–27, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3457195.

Hildebrandt:2021:VUL

[412] Mireille Hildebrandt. Viewpoint: Understanding law and the rule of law: a plea to augment CS curricula. Communications of the ACM, 64(5):28–31, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3425779.

Sochat:2021:VBP

[413] Vanessa Sochat. Viewpoint: The 10 best practices for remote software engineering. Communications of the ACM, 64(5):32–36, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3459613.

Dershowitz:2021:VLH

[414] Nachum Dershowitz. Viewpoint: Let’s be honest. Communications of the ACM, 64(5):37–41, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3431281.

Singh:2021:EC

[415] Jatinder Singh, Jennifer Cobbe, Do Le Quoc, and Zahra Tarkhani. Enclaves in the clouds. Communications of the ACM, 64(5):42–51, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3447543.

Frazelle:2021:BD

[416] Jessie Frazelle. Battery day. Communications of the ACM, 64(5):52–59, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3434222.

Grosse:2021:IID

[417] Eric Grosse, Fred B. Schneider, and Lynette L. Millett. Implementing insider defenses. Communications of the ACM, 64(5):60–65, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3418296.

Liu:2021:HCT

[418] Yuhang Liu, Xian-He Sun, Yang Wang, and Yungang Bao. HCDA: from computational thinking to a generalized thinking paradigm. Communications of the ACM, 64(5):66–75, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3418291.

Schleier-Smith:2021:WSC

[419] Johann Schleier-Smith, Vikram Sreekanti, Anurag Khandelwal, Joao Carreira, Neeraja J. Yadwadkar, Raluca Ada Popa, Joseph E. Gonzalez, Ion Stoica, and David A. Patterson. What serverless computing is and should become: the next phase of cloud computing. Communications of the ACM, 64(5): 76–84, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3406011.

DAntoni:2021:AMT

[420] Loris D’Antoni and Margus Veanes. Automata modulo theories. Communications of the ACM, 64(5):86–95, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3419404.

Schweitzer:2021:TPL

[421] Pascal Schweitzer. Technical perspective: a logical step toward the graph isomorphism problem. Communications of the ACM, 64(5):97, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453945.

Grohe:2021:ICD

[422] Martin Grohe and Daniel Neuen. Isomorphism, canonization, and definability for graphs of bounded rank width. Communications of the ACM, 64(5):98–105, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453943.

Steinhardt:2021:TPR

[423] Jacob S. Steinhardt. Technical perspective: Robust statistics tackle new problems. Communications of the ACM, 64(5):106, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453937.

Diakonikolas:2021:RMA

[424] Ilias Diakonikolas, Gautam Kamath, Daniel M. Kane, Jerry Li, Ankur Moitra, and Alistair Stewart. Robustness meets algorithms. Communications of the ACM, 64(5):107–115, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453935.

P-Ray:2021:LBB

[425] P-Ray. Last byte: Behold the Ch!Ld. Communications of the ACM, 64(5): 120–ff, May 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3453712.

Chien:2021:ELT

[426] Andrew A. Chien. Editor’s letter: Time for two annual Turing awards. Communications of the ACM, 64(6):5, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3462459. See letters

[477].

Cerf:2021:CIC

[427] Vinton G. Cerf. Cerf’s up: It came from outer space! Communications of the ACM, 64(6):7, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3462461.

Meil:2021:SUP

[428] Doug Meil. The search for unlimited productivity. Communications of the ACM, 64(6):10–11, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3460212.

Savage:2021:NGB

[429] Neil Savage. News: Getting down to basics. Communications of the ACM, 64(6):12–14, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3460734.

Monroe:2021:NDA

[430] Don Monroe. News: Deceiving AI. Communications of the ACM, 64(6): 15–16, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3460218.

Garfinkel:2021:NJM

[431] Simson Garfinkel and Eugene H. Spafford. News: Jack Minker (1927–2021). Communications of the ACM, 64(6):17, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3462465.

Krakovsky:2021:NTH

[432] Marina Krakovsky. News: Taking the heat: Maxwell’s demon and the high cost of erasure. Communications of the ACM, 64(6):18–20, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3460214.

Edwards:2021:NLA

[433] Chris Edwards. News: Let the algorithm decide? Communications of the ACM, 64(6):21–22, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3460216.

Mercuri:2021:IRR

[434] Rebecca T. Mercuri and Peter G. Neumann. Inside risks: The risks of election believability (or lack thereof). Communications of the ACM, 64(6): 24–30, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3461464.

Neville-Neil:2021:KVA

[435] George V. Neville-Neil. Kode vicious: Aversion to versions. Communications of the ACM, 64(6):31, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3461468.

Denning:2021:PIL

[436] Peter J. Denning. The profession of IT: Locality and professional life. Communications of the ACM, 64(6):32–34, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3461466.

Hylton:2021:VVC

[437] Todd Hylton, Thomas M. Conte, and Mark D. Hill. Viewpoint: A vision to compute like nature: thermodynamically. Communications of the ACM, 64(6):35–38, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3431282.

Yao:2021:VDP

[438] Danfeng (Daphne) Yao. Viewpoint: Depth and persistence: what researchers need to know about impostor syndrome. Communications of the ACM, 64(6):39–42, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3437255.

Littman:2021:VCR

[439] Michael L. Littman. Viewpoint: Collusion rings threaten the integrity of computer science research. Communications of the ACM, 64(6):43–44, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3429776.

Forsgren:2021:SDP

[440] Nicole Forsgren, Margaret-Anne Storey, Chandra Maddila, Thomas Zimmermann, Brian Houck, and Jenna Butler. The SPACE of developer productivity. Communications of the ACM, 64(6):46–53, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3453928.

Russinovich:2021:TCC

[441] Mark Russinovich, Manuel Costa, Cédric Fournet, David Chisnall, Antoine Delignat-Lavaud, Sylvan Clebsch, Kapil Vaswani, and Vikas Bhatia. Toward confidential cloud computing. Communications of the ACM, 64(6):54–61, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3453930.

Guan:2021:CER

[442] Hui Guan, Shaoshan Liu, Xiaolong Ma, Wei Niu, Bin Ren, Xipeng Shen, Yanzhi Wang, and Pu Zhao. CoCoPIE: enabling real-time AI on off-the-shelf mobile devices via compressioncompilation co-design. Communications of the ACM, 64(6):62–68, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3418297.

Franch:2021:RER

[443] Xavier Franch, Cristina Palomares, and Tony Gorschek. On the requirements engineer role. Communications of the ACM, 64(6):69–75, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3418292.

Misa:2021:DGB

[444] Thomas J. Misa. Dynamics of gender bias in computing. Communications of the ACM, 64(6):76–83, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3417517.

Sun:2021:SIA

[445] Yixin Sun, Maria Apostolaki, Henry Birge-Lee, Laurent Vanbever, Jennifer Rexford, Mung Chiang, and Prateek Mittal. Securing Internet applications from routing attacks. Communications of the ACM, 64(6):86–96, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3429775.

Bhattacharjee:2021:RLP

[446] Abhishek Bhattacharjee. Race logic presents a novel form of encoding: technical perspective. Communications of the ACM, 64(6):98, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3460221.

Tzimpragos:2021:SCB

[447] Georgios Tzimpragos, Advait Madhavan, Dilip Vasudevan, Dmitri Strukov, and Timothy Sherwood. In-sensor classification with boosted race trees. Communications of the ACM, 64(6):99–105, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3460223.

Jerger:2021:CPS

[448] Natalie Enright Jerger. A chiplet prototype system for deep learning inference: technical perspective. Communications of the ACM, 64(6):106, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3460225.

Shao:2021:SSD

[449] Yakun Sophia Shao, Jason Cemons, Rangharajan Venkatesan, Brian Zimmer, Matthew Fojtik, Nan Jiang, Ben Keller, Alicia Klinefelter, Nathaniel Pinckney, Priyanka Raina, Stephen G. Tell, Yanqing Zhang, William J. Dally, Joel Emer, C. Thomas Gray, Brucek Khailany, and Stephen W. Keckler. Simba: scaling deep-learning inference with chiplet-based architecture. Communications of the ACM, 64(6):107– 116, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3460227.

Hoffmann:2021:LBS

[450] Leah Hoffmann. Last byte: Shaping the foundations of programming languages. Communications of the ACM, 64(6):120–ff, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3460442.

Vardi:2021:DPVb

[451] Moshe Y. Vardi. Departments: Program verification: vision and reality. Communications of the ACM, 64(7):5, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3469113.

Holt:2021:CPC

[452] Victoria Holt. Career paths in computing: Computing enabled me to …: Obtain a Ph.D. and a career in data. Communications of the ACM, 64(7):7, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3464919.

Staff:2021:ELTa

[453] CACM Staff. Editor’s letter: Two sides of the software engineering coin. Communications of the ACM, 64(7):8– 9, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3466562.

Arquilla:2021:SSC

[454] John Arquilla, Judi Fusco, Pati Ruiz, and Jeremy Roschelle. Securing seabed cybersecurity, emphasizing intelligence augmentation. Communications of the ACM, 64(7):10–12, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3464931.

Greengard:2021:NFS

[455] Samuel Greengard. News: Formal software verification measures up. Communications of the ACM, 64(7):13–15, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3464933.

Shein:2021:NBF

[456] Esther Shein. News: a batteryfree Internet of Things. Communications of the ACM, 64(7):16–18, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3464937.

Marks:2021:NFS

[457] Paul Marks. News: The future of supply chains. Communications of the ACM, 64(7):19–21, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3464935.

Garfinkel:2021:NCM

[458] Simson Garfinkel and Eugene H. Spafford. News: Charles M. Geschke (1939–2021). Communications of the ACM, 64(7):22, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3467481.

Samuelson:2021:LSR

[459] Pamela Samuelson. Legally speaking: Reimplementing software interfaces is fair use. Communications of the ACM, 64(7):24–26, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466607.

Cranor:2021:PLL

[460] Lorrie Faith Cranor. Privacy: Lessons from the loo. Communications of the ACM, 64(7):27–29, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466609.

Barocas:2021:CER

[461] Solon Barocas, Asia J. Biega, Margarita Boyarskaya, Kate Crawford, Hal Daumé III, Miroslav Dudík, Benjamin Fish, Mary L. Gray, Brent Hecht, Alexandra Olteanu, Forough Poursabzi-Sangdeh, Luke Stark, Jennifer Wortman Vaughan, Hanna Wallach, and Marion Zepf. Computing ethics: Responsible computing during COVID-19 and beyond. Communications of the ACM, 64(7):30–32, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466612.

Domingo-Ferrer:2021:VLD

[462] Josep Domingo-Ferrer, David Sánchez, and Alberto Blanco-Justicia. Viewpoint: The limits of differential privacy (and its misuse in data release and machine learning). Communications of the ACM, 64(7):33–35, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3433638.

Spinellis:2021:VWC

[463] Diomidis Spinellis. Viewpoint: Why computing students should contribute to open source software projects. Communications of the ACM, 64(7):36–38, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3437254.

Breaux:2021:VSD

[464] Travis Breaux and Jennifer Moritz. Viewpoint: The 2021 software developer shortage is coming. Communications of the ACM, 64(7):39–41, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3440753.

Nokleberg:2021:AF

[465] Chris Nokleberg and Brad Hawkes. Application frameworks. Communications of the ACM, 64(7):42–49, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3446796.

Schlossnagle:2021:ATS

[466] Theo Schlossnagle, Justin Sheehy, and Chris McCubbin. Always-on timeseries database: keeping up where there’s no way to catch up. Communications of the ACM, 64(7):50–56, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442518.

Bengio:2021:TLD

[467] Yoshua Bengio, Yann LeCun, and Geoffrey Hinton. Turing lecture: Deep learning for AI. Communications of the ACM, 64(7):58–65, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3448250.

Knowles:2021:HCA

[468] Bran Knowles, Vicki L. Hanson, Yvonne Rogers, Anne Marie Piper, Jenny Waycott, Nigel Davies, Aloha Hufana Ambe, Robin N. Brewer, Debaleena Chattopadhyay, Marianne Dee, David Frohlich, Marisela Gutierrez-Lopez, Ben Jelen, Amanda Lazar, Radoslaw Nielek, Belén Barros Pena, Abi Roper, Mark Schlager, Britta Schulte, and Irene Ye Yuan. The harm in conflating aging with accessibility. Communications of the ACM, 64(7):66–71, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3431280.

Jarrahi:2021:FWP

[469] Mohammad Hossein Jarrahi, Gemma Newlands, Brian Butler, Saiph Savage, Christoph Lutz, Michael Dunn, and Steve Sawyer. Flexible work and personal digital infrastructures. Communications of the ACM, 64(7):72–79, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3419405.

Winter:2021:SCC

[470] Stephan Winter, Timothy Baldwin, Martin Tomko, Jochen Renz, Werner Kuhn, and Maria Vasardani. Spatial concepts in the conversation with a computer. Communications of the ACM, 64(7):82–88, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3440752.

Sankaranarayanan:2021:EMD

[471] Sriram Sankaranarayanan. An elegant model for deriving equations: technical perspective. Communications of the ACM, 64(7):90, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3465214.

Tsoutsouras:2021:DES

[472] Vasileios Tsoutsouras, Sam Willis, and Phillip Stanley-Marbell. Deriving equations from sensor data using dimensional function synthesis. Communications of the ACM, 64(7):91–99, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3465216.

Rexford:2021:TPD

[473] Jennifer Rexford. Tracking pandemicdriven internet traffic: technical perspective. Communications of the ACM, 64(7):100, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3465173.

Feldmann:2021:YLH

[474] Anja Feldmann, Oliver Gasser, Franziska Lichtblau, Enric Pujol, Ingmar Poese, Christoph Dietzel, Daniel Wagner, Matthias Wichtlhuber, Juan Tapiador, Narseo Vallina-Rodriguez, Oliver Hohlfeld, and Georgios Smaragdakis. A year in lockdown: how the waves of COVID-19 impact Internet traffic. Communications of the ACM, 64(7):101–108, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3465212.

Shasha:2021:LBSb

[475] Dennis Shasha. Last byte: String me along. Communications of the ACM, 64(7):112, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466564.

Cerf:2021:CC

[476] Vinton G. Cerf. Cerf’s up: On communication. Communications of the ACM, 64(8):5, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3472146.

Rector:2021:LET

[477] Gary Rector, Alexander Simonelis, Maurice van Swaaij, Alex Thomasian, Rebecca Mercuri, Andrew A. Chien, and Joseph P. Skudlarek. Letters to the editor: Tales of two Turings. Communications of the ACM, 64(8):7–9, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3469876. Comments on

proposal for two annual Turing awards [426].

Hazzan:2021:JID

[478] Orit Hazzan and Koby Mike. A journal for interdisciplinary data science education. Communications of the ACM, 64(8):10–11, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3469281.

Edwards:2021:NBS

[479] Chris Edwards. News: Better security through obfuscation. Communications of the ACM, 64(8):13–15, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3469283.

Kirkpatrick:2021:NFI

[480] Keith Kirkpatrick. News: Fixing the internet. Communications of the ACM, 64(8):16–17, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3469287.

Kugler:2021:NUT

[481] Logan Kugler. News: The unionization of technology companies. Communications of the ACM, 64(8):18–20, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3469285.

Wolf:2021:NUA

[482] Marty J. Wolf, Don Gotterbarn, and Michael Kirkpatrick. News: Upholding ACM’s principles. Communications of the ACM, 64(8):21, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3473051.

Petropoulos:2021:EBD

[483] Georgios Petropoulos. Economic and business dimensions: A European union approach to regulating big tech. Communications of the ACM, 64(8):24–26, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3469104.

Becker:2021:EWD

[484] Brett A. Becker. Education: What does saying that ‘programming is hard’ really say, and about whom? Communications of the ACM, 64(8):27–29, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3469115.

Neville-Neil:2021:KVPa

[485] George V. Neville-Neil. Kode vicious: In praise of the disassembler. Communications of the ACM, 64(8):30–31, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3469118.

Shneiderman:2021:VRA

[486] Ben Shneiderman. Viewpoint: Responsible AI: bridging from ethics to practice. Communications of the ACM, 64(8):32–35, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3445973.

Ghezzi:2021:VSN

[487] Carlo Ghezzi. Viewpoint: Science needs to engage with society: some lessons from COVID-19. Communications of the ACM, 64(8):36–38, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3473614.

Whalley:2021:VDC

[488] David Whalley, Xin Yuan, and Xiuwen Liu. Viewpoint: The domestic computer science graduate students are there, we just need to recruit them. Communications of the ACM, 64(8):39–43, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3448249.

Srinivasan:2021:BAS

[489] Ramya Srinivasan and Ajay Chander. Biases in AI systems. Communications of the ACM, 64(8):44–49, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3464903.

Blum:2021:WRT

[490] Niklas Blum, Serge Lachapelle, and Harald Alvestrand. WebRTC: real-time communication for the open web platform. Communications of the ACM, 64(8):50–54, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453182.

Baudin:2021:DPB

[491] Patrick Baudin, François Bobot, David Bühler, Loïc Correnson, Florent Kirchner, Nikolai Kosmatov, André Maroneze, Valentin Perrelle, Virgile Prevosto, Julien Signoles, and Nicky Williams. The dogged pursuit of bugfree C programs: the Frama-C software analysis platform. Communications of the ACM, 64(8):56–68, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3470569.

Dekel:2021:UUT

[492] Tali Dekel and Noah Snavely. Unveiling unexpected training data in Internet video. Communications of the ACM, 64(8):69–79, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3431283.

Wang:2021:MDD

[493] Xiaonan Wang and Xingwei Wang. Multimedia data delivery based on IoT clouds. Communications of the ACM, 64(8):80–86, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3474089.

Barros:2021:SCC

[494] Alistair Barros, Renuka Sindhgatta, and Alireza Nili. Scaling up chatbots for corporate service delivery systems. Communications of the ACM, 64(8):88–97, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3446912.

Chasins:2021:PHB

[495] Sarah E. Chasins, Elena L. Glassman, and Joshua Sunshine. PL and HCI: better together. Communications of the ACM, 64(8):98–106, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3469279.

Daskalakis:2021:QOM

[496] Constantinos Daskalakis. The quest for optimal multi-item auctions: technical perspective. Communications of the ACM, 64(8):108, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3470440.

Dutting:2021:OAT

[497] Paul Dütting, Zhe Feng, Harikrishna Narasimhan, David C. Parkes, and Sai S. Ravindranath. Optimal auctions through deep learning. Communications of the ACM, 64(8):109–116, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3470442.

Hester:2021:ERT

[498] Josiah D. Hester. eBP rides the third wave of mobile health: technical perspective. Communications of the ACM, 64(8):117, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3470444.

Bui:2021:EEW

[499] Nam Bui, Nhat Pham, Jessica Jacqueline Barnitz, Zhanan Zou, Phuc Nguyen, Hoang Truong, Taeho Kim, Nicholas Farrow, Anh Nguyen, Jianliang Xiao, Robin Deterding, Thang Dinh, and Tam Vu. eBP: an ear-worn device for frequent and comfortable blood pressure monitoring. Communications of the ACM, 64(8):118–125, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3470446.

Clegg:2021:LBAb

[500] Brian Clegg. Last byte: Agent algorithm: crime-solving computer plays by its own rules. Communications of the ACM, 64(8):128–ff, August 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3470629.

Vardi:2021:DSH

[501] Moshe Y. Vardi. Departments: The sand-heap paradox of privacy and influence. Communications of the ACM, 64(9):5, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3477583.

Sorensen:2021:CPC

[502] Andrew Sorensen. Career paths in computing: How music and programming led me to build digital microworlds. Communications of the ACM, 64(9):7, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3476462.

Staff:2021:ELTb

[503] CACM Staff. Editor’s letter: Turing reaction. Communications of the ACM, 64(9):9, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3477598.

Meil:2021:FAS

[504] Doug Meil and Mario Antoine Aoun. Finding the art in systems conversions, naming. Communications of the ACM, 64(9):10–11, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3474351.

Krakovsky:2021:NMR

[505] Marina Krakovsky. News: A model restoration. Communications of the ACM, 64(9):13–15, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3474353.

Greengard:2021:NPP

[506] Samuel Greengard. News: Photonic processors light the way. Communications of the ACM, 64(9):16–18, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3474357.

Kugler:2021:NNF

[507] Logan Kugler. News: Non-fungible tokens and the future of art. Communications of the ACM, 64(9):19–20, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3474355.

Chander:2021:LTP

[508] Anupam Chander. Law and technology: Protecting the global Internet from technology cold wars. Communications of the ACM, 64(9):22–24, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3473606.

Blumenthal:2021:SSD

[509] Marjory S. Blumenthal. Security: Security done right can make smart cities wise. Communications of the ACM, 64 (9):25–27, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3473608.

Haigh:2021:HRWb

[510] Thomas Haigh. Historical reflections: Women’s lives in code. Communications of the ACM, 64(9):28–34, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3473610.

Denning:2021:PIB

[511] Peter J. Denning. The profession of IT: Back of the envelope. Communications of the ACM, 64(9):35–37, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3473612.

Flowers:2021:VTE

[512] Lamont A. Flowers. Viewpoint: Testing educational digital games. Communications of the ACM, 64(9):38–40, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3450758.

Larus:2021:VWS

[513] James R. Larus. Viewpoint: Whose smartphone is it? Communications of the ACM, 64(9):41–42, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3454007.

Nourbakhsh:2021:VAE

[514] Illah Reza Nourbakhsh. Viewpoint: AI ethics: a call to faculty. Communications of the ACM, 64(9):43–45, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3478516.

Mashatan:2021:CPQ

[515] Atefeh Mashatan and Douglas Heintzman. The complex path to quantum resistance. Communications of the ACM, 64(9):46–53, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3464905.

Staff:2021:QST

[516] CACM Staff. Quantum-safe trust for vehicles: the race is already on. Communications of the ACM, 64(9): 54–61, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466174.

Sakr:2021:FBG

[517] Sherif Sakr, Angela Bonifati, Hannes Voigt, Alexandru Iosup, Khaled Ammar, Renzo Angles, Walid Aref, Marcelo Arenas, Maciej Besta, Peter A. Boncz, Khuzaima Daudjee, Emanuele Della Valle, Stefania Dumbrava, Olaf Hartig, Bernhard Haslhofer, Tim Hegeman, Jan Hidders, Katja Hose, Adriana Iamnitchi, Vasiliki Kalavri, Hugo Kapp, Wim Martens, M. Tamer Özsu, Eric Peukert, Stefan Plantikow, Mohamed Ragab, Matei R. Ripeanu, Semih Salihoglu, Christian Schulz, Petra Selmer, Juan F. Sequeda, Joshua Shinavier, Gábor Szárnyas, Riccardo Tommasini, Antonino Tumeo, Alexandru Uta, Ana Lucia Varbanescu, Hsiang-Yun Wu, Nikolay Yakovets, Da Yan, and Eiko Yoneki. The future is big graphs: a community view on graph processing systems. Communications of the ACM, 64(9):62–71, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3434642.

Idell:2021:MIP

[518] Kelly Idell, David Gefen, and Arik Ragowsky. Managing IT professional turnover. Communications of the ACM, 64(9):72–77, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3434641.

Kress-Gazit:2021:FGH

[519] Hadas Kress-Gazit, Kerstin Eder, Guy Hoffman, Henny Admoni, Brenna Argall, Rüdiger Ehlers, Christoffer Heckman, Nils Jansen, Ross Knepper, Jan Kretínský, Shelly Levy-Tzedek, Jamy Li, Todd Murphey, Laurel Riek, and Dorsa Sadigh. Formalizing and guaranteeing human-robot interaction. Communications of the ACM, 64(9):78–84, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3433637.

Bouguettaya:2021:ITS

[520] Athman Bouguettaya, Quan Z. Sheng, Boualem Benatallah, Azadeh Ghari Neiat, Sajib Mistry, Aditya Ghose, Surya Nepal, and Lina Yao. An Internet of Things service roadmap. Communications of the ACM, 64(9): 86–95, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3464960.

Morgenstern:2021:IWT

[521] Leora Morgenstern. The importance of WINOGRANDE: technical perspective. Communications of the ACM, 64(9):98, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3474378.

Sakaguchi:2021:WAW

[522] Keisuke Sakaguchi, Ronan Le Bras, Chandra Bhagavatula, and Yejin Choi. WinoGrande: an adversarial Winograd schema challenge at scale. Communications of the ACM, 64(9):99–106, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3474381.

Balietti:2021:DYE

[523] Stefano Balietti. Does your experiment smell?: technical perspective. Communications of the ACM, 64(9):107, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3474383.

Tosch:2021:PAT

[524] Emma Tosch, Eytan Bakshy, Emery D. Berger, David D. Jensen, and J. Eliot B. Moss. PlanAlyzer: assessing threats to the validity of online experiments. Communications of the ACM, 64(9):108–116, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3474385.

Hoffmann:2021:LBP

[525] Leah Hoffmann. Last byte: Playing with, and against, computers. Communications of the ACM, 64(9): 120–ff, September 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3474361.

Cerf:2021:CFT

[526] Vinton G. Cerf. Cerf’s up: The future of text redux. Communications of the ACM, 64(10):5, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3483539.

Sukhov:2021:NLC

[527] Andrei Sukhov, Igor Sorokin, and Doug Meil. New life for cordless communication, old regrets for software projects. Communications of the ACM, 64(10):6–7, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3479972.

Monroe:2021:NFC

[528] Don Monroe. News: Flatter chips. Communications of the ACM, 64 (10):8–10, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3479974.

Kirkpatrick:2021:NAP

[529] Keith Kirkpatrick. News: Algorithmic poverty. Communications of the ACM, 64(10):11–12, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3479977.

Edwards:2021:NST

[530] Chris Edwards. News: A switch in time. Communications of the ACM, 64(10):13–15, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3479979.

Cusumano:2021:TSMb

[531] Michael A. Cusumano. Technology strategy and management: section 230 and a tragedy of the commons. Communications of the ACM, 64(10):16–18, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481354.

Walther:2021:BPB

[532] Kendra Walther and Richard E. Ladner. Broadening participation: Broadening participation by teaching accessibility. Communications of the ACM, 64(10):19–21, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481356.

Thinyane:2021:GCR

[533] Hannah Thinyane. Global computing: Remaining connected throughout design. Communications of the ACM, 64(10):22–24, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481429.

Neville-Neil:2021:KVD

[534] George V. Neville-Neil. Kode vicious: Divide and conquer. Communications of the ACM, 64(10):25, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3481431.

Doctorow:2021:VCC

[535] Cory Doctorow. Viewpoint: Competitive compatibility: let’s fix the Internet, not the tech giants. Communications of the ACM, 64(10):26–29, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3446789.

Dubhashi:2021:EBD

[536] Devdatt Dubhashi. Economic and business dimensions: AI futures: fact and fantasy. Communications of the ACM, 64(10):30–31, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481581.

Varajao:2021:SDD

[537] João Varajão. Software development in disruptive times. Communications of the ACM, 64(10):32–35, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453932.

Frazelle:2021:NEM

[538] Jessie Frazelle. A new era for mechanical CAD. Communications of the ACM, 64(10):36–39, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3464909.

Groh:2021:HDM

[539] Matthew Groh, Ziv Epstein, Nick Obradovich, Manuel Cebrian, and Iyad Rahwan. Human detection of machinemanipulated media. Communications of the ACM, 64(10):40–47, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3445972.

Torro:2021:SRW

[540] Osku Torro, Henri Jalo, and Henri Pirkkalainen. Six reasons why virtual reality is a game-changing computing and communication platform for organizations. Communications of the ACM, 64(10):48–55, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3440868.

Tannian:2021:RPC

[541] Mark Tannian and Willie Coston. The role of professional certifications in computer occupations. Communications of the ACM, 64(10):56–63, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3474359.

Wing:2021:TA

[542] Jeannette M. Wing. Trustworthy AI. Communications of the ACM, 64 (10):64–71, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3448248.

Vu:2021:LTU

[543] Tam Vu. Liquid testing using builtin phone sensors: technical perspective. Communications of the ACM, 64(10):74, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481036.

Yue:2021:LTY

[544] Shichao Yue and Dina Katabi. Liquid testing with your smartphone. Communications of the ACM, 64(10):75– 83, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3481038.

Sadeghi:2021:RWD

[545] Ahmad-Reza Sadeghi. The real-world dilemma of security and privacy by design: technical perspective. Communications of the ACM, 64(10):84, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3481040.

Lee:2021:SWE

[546] Jihoon Lee, Gyuhong Lee, Jinsung Lee, Youngbin Im, Max Hollingsworth, Eric Wustrow, Dirk Grunwald, and Sangtae Ha. Securing the wireless emergency alerts system. Communications of the ACM, 64(10):85–93, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3481042.

Shasha:2021:LBRb

[547] Dennis Shasha. Last byte: Randomower. Communications of the ACM, 64(10):96–ff, October 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3477385.

Vardi:2021:DPC

[548] Moshe Y. Vardi. Departments: The paradox of choice in computingresearch conferences. Communications of the ACM, 64(11):5, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3488554.

Moffatt:2021:CPC

[549] Ann Moffatt. Career paths in computing: Grace Hopper, minicomputers, and megabytes: it’s a fun career. Communications of the ACM, 64(11):7, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3485446.

Sukhov:2021:BAI

[550] Andrei Sukhov and Vivek S. Buzruk. BLOG@CACM: Assessing Internet software engineering, encouraging competitions. Communications of the ACM, 64(11):8–9, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3484986.

Greengard:2021:NQD

[551] Samuel Greengard. News: Qubit devices inch toward reality. Communications of the ACM, 64(11):11–13, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3484988.

Edwards:2021:NHH

[552] Chris Edwards. News: Holograms on the horizon? Communications of the ACM, 64(11):14–16, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3484998.

Shein:2021:NFB

[553] Esther Shein. News: Filtering for beauty. Communications of the ACM, 64(11):17–19, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3484997.

Samuelson:2021:LST

[554] Pamela Samuelson. Legally speaking: Text and data mining of in-copyright works: is it legal? Communications of the ACM, 64(11):20–22, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3486628.

Kissner:2021:PPE

[555] Lea Kissner and Lorrie Cranor. Privacy: Privacy engineering superheroes. Communications of the ACM, 64(11): 23–25, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3486631.

Shilton:2021:CES

[556] Katie Shilton, Megan Finn, and Quinn DuPont. Computing ethics: Shaping ethical computing cultures. Communications of the ACM, 64(11):26–29, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3486639.

Repenning:2021:EEP

[557] Alexander Repenning and Ashok Basawapatna. Education: Explicative programming. Communications of the ACM, 64(11):30–33, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3486642.

Stoger:2021:VMA

[558] Karl Stöger, David Schneeberger, and Andreas Holzinger. Viewpoint: Medical artificial intelligence: the European legal perspective. Communications of the ACM, 64(11):34–36, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3458652.

Reich:2021:VWU

[559] Yoram Reich and Eswaran Subrahmanian. Viewpoint: We are not users: gaining control over new technologies. Communications of the ACM, 64(11):37–39, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3486595.

Jin:2021:CRS

[560] Hai Jin, Yuanchun Shi, and Dahua Lin. China region special section: Welcome back! Communications of the ACM, 64 (11):40–42, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481736.

Su:2021:CRS

[561] Fang Su, Xiao-Peng An, and Ji-Ye Mao. China region special section: Hot topics: Innovations and trends in China’s digital economy. Communications of the ACM, 64(11):44–47, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3481604.

Yuan:2021:CRS

[562] Chunfeng Yuan, Xiaopeng Gao, Yu Chen, and Yungang Bao. China region special section: Hot topics: Teaching undergraduates to build real computer systems. Communications of the ACM, 64(11):48–49, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3481606.

Han:2021:CRS

[563] Xu Han, Zhengyan Zhang, and Zhiyuan Liu. China region special section: Hot topics: Knowledgeable machine learning for natural language processing. Communications of the ACM, 64(11):50–51, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3481608.

Wu:2021:CRS

[564] Fei Wu, Qinming He, and Chao Wu. China region special section: Hot topics: AI+X micro-program fosters interdisciplinary skills in China. Communications of the ACM, 64(11):52–54, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3481612.

Yang:2021:CRS

[565] Jing Yang. China region special section: Hot topics: AI start-ups in China. Communications of the ACM, 64(11):55–56, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3481610.

Tian:2021:CRS

[566] Feng Tian, Yuntao Wang, and Yicheng Zhu. China region special section: Hot topics: Natural interactive techniques for the detection and assessment of neurological diseases. Communications of the ACM, 64(11):57–59, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3481616.

Zhang:2021:CRSa

[567] Yiming Zhang, Kai Lu, and Wenguang Chen. China region special section: Hot topics: Processing extremescale graphs on China’s supercomputers. Communications of the ACM, 64 (11):60–63, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481614.

Zhang:2021:CRSb

[568] Guofeng Zhang, Xiaowei Zhou, Feng Tian, Hongbin Zha, Yongtian Wang, and Hujun Bao. China region special section: Big trends: The present and future of mixed reality in China. Communications of the ACM, 64(11): 64–69, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3481619.

Yu:2021:CRSa

[569] Chun Yu and Jiajun Bu. China region special section: Big trends: The practice of applying AI to benefit visually impaired people in China. Communications of the ACM, 64(11): 70–75, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3481623.

Yu:2021:CRSb

[570] Zhiwen Yu, Huadong Ma, Bin Guo, and Zheng Yang. China region special section: Big trends: Crowdsensing 2.0. Communications of the ACM, 64 (11):76–80, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3481621.

Jia:2021:CRS

[571] Jia Jia, Wei Chen, Kai Yu, Xiaodong He, Jun Du, and Heung-Yeung Shum. China region special section: Big trends: The practice of speech and language processing in China. Communications of the ACM, 64(11):81–87, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3481625.

Cai:2021:CRS

[572] Liang Cai, Yi Sun, Zibin Zheng, Jiang Xiao, and Weiwei Qiu. China region special section: Big trends: Blockchain in China. Communications of the ACM, 64(11):88–93, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3481627.

Kamp:2021:WWW

[573] Poul-Henning Kamp. What went wrong? Communications of the ACM, 64(11):94–96, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3475165.

Groger:2021:TNA

[574] Christoph Gröger. There is no AI without data. Communications of the ACM, 64(11):98–108, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3448247.

Bastani:2021:IIS

[575] Favyen Bastani, Songtao He, Satvat Jagwani, Edward Park, Sofiane Abbar, Mohammad Alizadeh, Hari Balakrishnan, Sanjay Chawla, Sam Madden, and Mohammad Amin Sadeghi. Inferring and improving street maps with datadriven automation. Communications of the ACM, 64(11):109–117, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3450351.

Hentenryck:2021:TPF

[576] Pascal Van Hentenryck. Technical perspective: Finding the sweet spot amid accuracy and performance. Communications of the ACM, 64(11):120, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3485594.

Cristian:2021:MIO

[577] Alexandru Cristian, Luke Marshall, Mihai Negrea, Flavius Stoichescu, Peiwei Cao, and Ishai Menache. Multiitinerary optimization as cloud service. Communications of the ACM, 64(11): 121–129, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3485626.

Aharonov:2021:TPP

[578] Dorit Aharonov and Michael Chapman. Technical perspective: On proofs, entanglement, and games. Communications of the ACM, 64(11):130, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3485596.

Ji:2021:MR

[579] Zhengfeng Ji, Anand Natarajan, Thomas Vidick, John Wright, and Henry Yuen. MIP* = RE. Communications of the ACM, 64(11):131–138, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3485628.

Bainbridge:2021:LBW

[580] William Sims Bainbridge. Last byte: World of hackcraft: an obsessive gamer’s quest for the absolutely most significant computer ever. Communications of the ACM, 64(11):144–ff, November 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3484694.

Hanson:2021:ACC

[581] Vicki L. Hanson. From the ACM CEO: Communicating ACM priorities. Communications of the ACM, 64(12):5, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3492320.

Chien:2021:ELG

[582] Andrew A. Chien. Editor’s letter: Good, better, best: how sustainable should computing be? Communications of the ACM, 64(12):6–7, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3492903.

Cerf:2021:CHC

[583] Vinton G. Cerf. Cerf’s up: On heterogeneous computing. Communications of the ACM, 64(12):9, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3492896.

Staff:2021:LEC

[584] CACM Staff. Letters to the editor: Common ails. Communications of the ACM, 64(12):10–12, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3485271.

Hill:2021:SCB

[585] Robin K. Hill and Carlos Baquero. Seeking out Camille, and being open to others. Communications of the ACM, 64(12):14–15, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3490151.

Monroe:2021:NTS

[586] Don Monroe. News: Trouble at the source. Communications of the ACM, 64(12):17–19, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3490155.

Kirkpatrick:2021:NRA

[587] Keith Kirkpatrick. News: The road ahead for augmented reality. Communications of the ACM, 64(12):20–22, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3490317.

Greengard:2021:NWC

[588] Samuel Greengard. News: What is the cost of living online? Communications of the ACM, 64(12):23–25, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3490165.

Davidson:2021:APF

[589] Jack W. Davidson, Joseph A. Konstan, and Scott E. Delman. ACM publications finances: ACM publications finances for 2020. Communications of the ACM, 64(12):26–27, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3488236.

Denning:2021:PIC

[590] Peter J. Denning and Matti Tedre. The profession of IT: Computational thinking for professionals. Communications of the ACM, 64(12):30–33, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3491268.

Alstyne:2021:EBD

[591] Marshall W. Van Alstyne, Georgios Petropoulos, Geoffrey Parker, and Bertin Martens. Economic and business dimensions: ’in situ’ data rights. Communications of the ACM, 64(12): 34–35, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3491270.

Hill:2021:VAL

[592] Mark D. Hill and Vijay Janapa Reddi. Viewpoint: Accelerator-level parallelism. Communications of the ACM, 64(12):36–38, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3460970.

Neville-Neil:2021:KVPb

[593] George V. Neville-Neil. Kode vicious: Patent absurdity. Communications of the ACM, 64(12):39, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3491297.

Picard:2021:VWE

[594] Rosalind Picard. Viewpoint: What every engineer and computer scientist should know: the biggest contributor to happiness. Communications of the ACM, 64(12):40–42, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3465999.

Leqi:2021:WCB

[595] Liu Leqi, Dylan Hadfield-Menell, and Zachary C. Lipton. When curation becomes creation. Communications of the ACM, 64(12):44–47, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3475161.

Gehani:2021:DBP

[596] Ashish Gehani, Raza Ahmad, Hassaan Irshad, Jianqiao Zhu, and Jignesh Patel. Digging into big provenance (with SPADE). Communications of the ACM, 64(12):48–56, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3475358.

Hooker:2021:HL

[597] Sara Hooker. The hardware lottery. Communications of the ACM, 64 (12):58–65, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3467017.

Xu:2021:ACA

[598] Xiaowei Xu, Hailong Qiu, Qianjun Jia, Yuhao Dong, Zeyang Yao, Wen Xie, Huiming Guo, Haiyun Yuan, Jian Zhuang, Meiping Huang, and Yiyu Shi. AI-CHD: an AI-based framework for cost-effective surgical telementoring of congenital heart disease. Communications of the ACM, 64(12):66–74, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3450409.

Chandra:2021:DAS

[599] Ranveer Chandra and Stewart Collis. Digital agriculture for small-scale producers: challenges and opportunities. Communications of the ACM, 64 (12):75–84, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3454008.

Gebru:2021:DD

[600] Timnit Gebru, Jamie Morgenstern, Briana Vecchione, Jennifer Wortman Vaughan, Hanna Wallach, Hal Daumé III, and Kate Crawford. Datasheets for datasets. Communications of the ACM, 64(12):86–92, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3458723.

Adib:2021:TPC

[601] Fadel Adib. Technical perspective: Cooking up a solution to microwave heat distribution. Communications of the ACM, 64(12):94, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3491194.

Jin:2021:SDC

[602] Haojian Jin, Jingxian Wang, Swarun Kumar, and Jason Hong. Softwaredefined cooking using a microwave oven. Communications of the ACM, 64 (12):95–103, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3491196.

Sherwood:2021:TPR

[603] Timothy Sherwood. Technical perspective: a recipe for protecting against speculation attacks. Communications of the ACM, 64(12):104, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3491198.

Yu:2021:STT

[604] Jiyong Yu, Mengjia Yan, Artem Khyzha, Adam Morrison, Josep Torrellas, and Christopher W. Fletcher. Speculative taint tracking (STT): a comprehensive protection for speculatively accessed data. Communications of the ACM, 64(12):105–112, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3491201.

Hoffmann:2021:LBE

[605] Leah Hoffmann. Last byte: Exploring the promise of quantum computing. Communications of the ACM, 64 (12):120–ff, December 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3490319.

Chien:2022:ELG

[606] Andrew A. Chien. Editor’s letter: Is the global computing community irrevocably divided? Communications of the ACM, 65(1):5, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3502906.

Vardi:2022:DWA

[607] Moshe Y. Vardi. Departments: Will AI destroy education? Communications of the ACM, 65(1):7, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3501359.

Small:2022:CPC

[608] Peter M. Small. Career paths in computing: How computing empowered me to tackle big problems in medicine. Communications of the ACM, 65(1):9, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3495544.

Guzdial:2022:HEE

[609] Mark Guzdial. Is higher education an engine of opportunity, or a perpetuator of privilege? Communications of the ACM, 65(1):10–11, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3495561.

Edwards:2022:NSA

[610] Chris Edwards. News: Shrinking artificial intelligence. Communications of the ACM, 65(1):12–14, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3495562.

Shein:2022:NCL

[611] Esther Shein. News: Converting laws to programs. Communications of the ACM, 65(1):15–16, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3495564.

Kirkpatrick:2022:NMY

[612] Keith Kirkpatrick. News: Monetizing your personal data. Communications of the ACM, 65(1):17–19, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3495563.

Cusumano:2022:TSMa

[613] Michael A. Cusumano. Technology strategy and management: Epic versus Apple and the future of app stores. Communications of the ACM, 65(1):22–24, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3498659.

Kim:2022:LTA

[614] Pauline T. Kim. Law and technology: Addressing algorithmic discrimination. Communications of the ACM, 65(1):25–27, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3498660.

LaMacchia:2022:SLR

[615] Brian LaMacchia. Security: The long road ahead to transition to postquantum cryptography. Communications of the ACM, 65(1):28–30, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3498706.

Hellendoorn:2022:VGC

[616] Vincent J. Hellendoorn and Anand Ashok Sawant. Viewpoint: The growing cost of deep learning for source code. Communications of the ACM, 65(1):31–33, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3501261.

Fink:2022:VAI

[617] Olga Fink, Torbjørn Netland, and Stefan Feuerriegel. Viewpoint: Artificial intelligence across company borders. Communications of the ACM, 65(1):34–36, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3470449.

Matyas:2022:VEI

[618] Vashek Matyas, Kamil Malinka, Lydia Kraus, Lenka Knapova, and Agata Kruzikova. Viewpoint: Even if users do not read security directives, their behavior is not so catastrophic. Communications of the ACM, 65(1):37–40, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3471928.

Molino:2022:DML

[619] Piero Molino and Christopher Ré. Declarative machine learning systems. Communications of the ACM, 65 (1):42–49, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3475167.

Thomson:2022:SA

[620] Patrick Thomson. Static analysis. Communications of the ACM, 65(1):50–54, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3486592.

Kindt:2022:HRS

[621] Philipp H. Kindt, Trinad Chakraborty, and Samarjit Chakraborty. How reliable is smartphone-based electronic contact tracing for COVID-19? Communications of the ACM, 65(1):56–67, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3471933.

Lee:2022:DRD

[622] Gwanhoo Lee and Jaeho Kim. Delivering a rapid digital response to the COVID-19 pandemic. Communications of the ACM, 65(1):68–75, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3484732.

Fortnow:2022:FYP

[623] Lance Fortnow. Fifty years of P vs. NP and the possibility of the impossible. Communications of the ACM, 65(1):76–85, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3460351.

Pradel:2022:NSA

[624] Michael Pradel and Satish Chandra. Neural software analysis. Communications of the ACM, 65(1):86–96, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3460348.

Dellaert:2022:NRF

[625] Frank Dellaert. Neural radiance fields explode on the scene: technical perspective. Communications of the ACM, 65(1):98, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3498368.

Mildenhall:2022:NRS

[626] Ben Mildenhall, Pratul P. Srinivasan, Matthew Tancik, Jonathan T. Barron, Ravi Ramamoorthi, and Ren Ng. NeRF: representing scenes as neural radiance fields for view synthesis. Communications of the ACM, 65 (1):99–106, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3503250.

Guerreiro:2022:EGE

[627] Tiago Guerreiro. Eyelid gestures enhance mobile interaction: technical perspective. Communications of the ACM, 65(1):107, January 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3498365.

Fan:2022:EGP

[628] Mingming Fan, Zhen Li, and Franklin Mingzhe Li. Eyelid gestures for people with motor impairments. Communications of the ACM, 65(1):108–115, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3498367.

Shasha:2022:LBO

[629] Dennis Shasha. Last byte: Orbit design. Communications of the ACM, 65(1):120–ff, January 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3498585.

Hanson:2022:ACE

[630] Vicki L. Hanson. From the ACM CEO: Ethics and values in the ACM awards program. Communications of the ACM, 65(2):5, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3507697.

Cerf:2022:CHD

[631] Vinton G. Cerf. Cerf’s up: How do I rely on thee, smartphone? Communications of the ACM, 65(2):7, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3505589.

Hong:2022:TF

[632] Jason Hong. The trouble with Facebook. Communications of the ACM, 65(2):8–9, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3503775.

Krakovsky:2022:NSW

[633] Marina Krakovsky. News: Solving for why. Communications of the ACM, 65(2):11–13, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3503777.

Savage:2022:NVD

[634] Neil Savage. News: Virtual duplicates. Communications of the ACM, 65(2):14–16, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3503798.

Kugler:2022:NAI

[635] Logan Kugler. News: Artificial intelligence, machine learning, and the fight against world hunger. Communications of the ACM, 65(2):17–19, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3503779.

Williams:2022:ELH

[636] Tiffani L. Williams. Education: The lives of hidden figures matter in computer science education. Communications of the ACM, 65(2):20–22, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3506577.

Neville-Neil:2022:KVU

[637] George Neville-Neil. Kode vicious: I unplugged what? Communications of the ACM, 65(2):23–24, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3506579.

Haigh:2022:HRB

[638] Thomas Haigh. Historical reflections: Becoming universal. Communications of the ACM, 65(2):25–30, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3506578.

Tu:2022:VYA

[639] Yu-Ju (Tony) Tu, Shari S. Shang, and Junyi Wu. Viewpoint: Is your autonomous vehicle as smart as you expected? Communications of the ACM, 65(2):31–34, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3506574.

Hod:2022:VDS

[640] Shlomi Hod, Karni Chagal-Feferkorn, Niva Elkin-Koren, and Avidor Gal. Viewpoint: Data science meets law. Communications of the ACM, 65(2): 35–39, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3506575.

Prem:2022:VBN

[641] Erich Prem. Viewpoint: A brave new world of mediated online discourse. Communications of the ACM, 65(2):40–42, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3471929.

Clem:2022:SAG

[642] Timothy Clem and Patrick Thomson. Static analysis at GitHub. Communications of the ACM, 65(2):44–51, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3486594.

Tiganov:2022:DUS

[643] Daniil Tiganov, Lisa Nguyen Quang Do, and Karim Ali. Designing UIs for static-analysis tools. Communications of the ACM, 65(2):52–58, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3486600.

Tomasev:2022:RCA

[644] Nenad Tomasev, Ulrich Paquet, Demis Hassabis, and Vladimir Kramnik. Reimagining chess with AlphaZero.

Communications of the ACM, 65(2): 60–66, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3460349.

Nevo:2022:HTN

[645] Dorit Nevo and Benjamin D. Horne. How topic novelty impacts the effectiveness of news veracity interventions. Communications of the ACM, 65(2):68–75, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3460350.

Aho:2022:TLA

[646] Alfred Aho and Jeffrey Ullman. Turing lecture: Abstractions, their algorithms, and their compilers. Communications of the ACM, 65(2):76–91, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3490685.

Halevy:2022:PIO

[647] Alon Halevy, Cristian Canton-Ferrer, Hao Ma, Umut Ozertem, Patrick Pantel, Marzieh Saeidi, Fabrizio Silvestri, and Ves Stoyanov. Preserving integrity in online social networks. Communications of the ACM, 65(2):92–98, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3462671.

Burke:2022:PRP

[648] Robin Burke. Personalized recommendation of PoIs to people with autism: technical perspective. Communications of the ACM, 65(2):100, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3505266.

Mauro:2022:SPA

[649] Noemi Mauro, Liliana Ardissono, and Federica Cena. Supporting people with autism spectrum disorders in the exploration of PoIs: an inclusive recommender system. Communications of the ACM, 65(2):101–109, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3505267.

Mitra:2022:MST

[650] Sayan Mitra. Model structure takes guesswork out of state estimation: technical perspective. Communications of the ACM, 65(2):110, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3505268.

Berger:2022:WCT

[651] Guillaume O. Berger and Raphael M. Jungers. Worst-case topological entropy and minimal data rate for state estimation of switched linear systems. Communications of the ACM, 65(2): 111–118, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3505269.

Clegg:2022:LBAa

[652] Brian Clegg. Last byte: Agent invisible. Communications of the ACM, 65(2):120–ff, February 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3501765.

Vardi:2022:DAE

[653] Moshe Y. Vardi. Departments: ACM, ethics, and corporate behavior. Communications of the ACM, 65(3):5, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3516423.

Sari:2022:CPC

[654] Eunice Sari. Career paths in computing: Changing the world through HCI and UX. Communications of the ACM, 65(3):7, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3510541.

Staff:2022:LED

[655] CACM Staff. Letters to the editor: A division in computer science. Communications of the ACM, 65(3):8–9, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3511660.

Baquero:2022:PPT

[656] Carlos Baquero. Picking publication targets. Communications of the ACM, 65(3):10–11, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3510545.

Edwards:2022:NGE

[657] Chris Edwards. News: A group effort. Communications of the ACM, 65(3):12–14, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3510550.

Monroe:2022:NAA

[658] Don Monroe. News: Accelerating AI. Communications of the ACM, 65(3):15–16, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3510554.

Marks:2022:NAH

[659] Paul Marks. News: Algorithmic hiring needs a human face. Communications of the ACM, 65(3):17–19, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3510552.

Samuelson:2022:LSC

[660] Pamela Samuelson. Legally speaking: Copyright implications of emulation programs. Communications of the ACM, 65(3):20–22, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3512293.

Baecker:2022:CEC

[661] Ronald M. Baecker. Computing ethics: A call to action. Communications of the ACM, 65(3):23–25, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512297.

Skirpan:2022:PPC

[662] Michael Skirpan, Maggie Oates, Daragh Byrne, Robert Cunningham, and Lorrie Faith Cranor. Privacy: Is a privacy crisis experienced, a privacy crisis avoided? Communications of the ACM, 65(3):26–29, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3512325.

Gasser:2022:VFD

[663] Urs Gasser and Virgílio Almeida. Viewpoint: Futures of digital governance. Communications of the ACM, 65(3):30–32, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3477502.

Madnick:2022:VBF

[664] Stuart Madnick. Viewpoint: Bursting a few balloons regarding the famous DARPA red balloon challenge. Communications of the ACM, 65(3):33–34, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3517127.

Walsh:2022:VTF

[665] Toby Walsh. Viewpoint: The troubling future for facial recognition software. Communications of the ACM, 65(3):35–36, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3474096.

Nadeem:2022:HCA

[666] Ayman Nadeem. Human-centered approach to static-analysis-driven developer tools. Communications of the ACM, 65(3):38–45, March 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3486597.

McKusick:2022:CMS

[667] Kirk McKusick. A conversation with: Margo Seltzer and Mike Olson. Communications of the ACM, 65(3):46–51, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3498582.

Murphy-Hill:2022:PER

[668] Emerson Murphy-Hill, Ciera Jaspan, Carolyn Egelman, and Lan Cheng. The pushback effects of race, ethnicity, gender, and age in code review. Communications of the ACM, 65(3):52–57, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3474097.

Weiser:2022:CWO

[669] Orli Weiser, Yoram M. Kalman, Carmel Kent, and Gilad Ravid. 65 competencies: which ones should your data analytics experts have? Communications of the ACM, 65(3):58–66, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3467018.

Mangla:2022:TTD

[670] Tarun Mangla, Esther Showalter, Vivek Adarsh, Kipp Jones, Morgan Vigil-Hayes, Elizabeth Belding, and Ellen Zegura. A tale of three datasets: characterizing mobile broadband access in the U.S. Communications of the ACM, 65(3):67–74, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3462672.

DeBie:2022:ADS

[671] Tijl De Bie, Luc De Raedt, José Hernández-Orallo, Holger H. Hoos, Padhraic Smyth, and Christopher K. I. Williams. Automating data science. Communications of the ACM, 65(3):76–87, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3495256.

Aldrich:2022:TPH

[672] Jonathan Aldrich. Technical perspective: How do experts learn new programming languages? Communications of the ACM, 65(3):90, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3511061.

Shrestha:2022:HWG

[673] Nischal Shrestha, Colton Botta, Titus Barik, and Chris Parnin. Here we go again: why is it difficult for developers to learn another programming language? Communications of the ACM, 65(3):91–99, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3511062.

Chong:2022:TPA

[674] Frederic T. Chong. Technical perspective: Applying design-space exploration to quantum architectures. Communications of the ACM, 65(3):100, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3511063.

Murali:2022:TSA

[675] Prakash Murali, Dripto M. Debroy, Kenneth R. Brown, and Margaret Martonosi. Toward systematic architectural design of near-term trapped ion quantum computers. Communications of the ACM, 65(3):101–109, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3511064.

Hoffmann:2022:LBM

[676] Leah Hoffmann. Last byte: Mining energy from ’empty’ air. Communications of the ACM, 65(3):112–ff, March 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3510556.

Cerf:2022:CPI

[677] Vinton G. Cerf. Cerf’s up: Preserving the Internet. Communications of the ACM, 65(4):5, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3522782.

Guzdial:2022:BAC

[678] Mark Guzdial. BLOG@CACM: Achieving CS for all could take decades. Communications of the ACM, 65(4):6–7, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3516513.

Greengard:2022:NCA

[679] Samuel Greengard. News: Can AI learn to forget? Communications of the ACM, 65(4):9–11, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3516514.

Kirkpatrick:2022:NSW

[680] Keith Kirkpatrick. News: Still waiting for self-driving cars. Communications of the ACM, 65(4):12–14, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3516517.

Kugler:2022:NTI

[681] Logan Kugler. News: Technology’s impact on morality. Communications of the ACM, 65(4):15–16, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3516516.

Sako:2022:TSM

[682] Mari Sako. Technology strategy and management: Global supply chain disruption and resilience. Communications of the ACM, 65(4):18–21, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3517216.

Neville-Neil:2022:KVG

[683] George V. Neville-Neil. Kode vicious: Getting off the mad path. Communications of the ACM, 65(4):25–26, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3517219.

Storey:2022:VEA

[684] Veda C. Storey, Roman Lukyanenko, Wolfgang Maass, and Jeffrey Parsons.

Viewpoint: Explainable AI. Communications of the ACM, 65(4):27–29, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3490699.

Chien:2022:VCD

[685] Andrew A. Chien. Viewpoint: Communications’ digital initiative and its first digital event. Communications of the ACM, 65(4):30, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3517329.

Cauchard:2022:ERS

[686] Jessica R. Cauchard, Matthias Jarke, and Nuria Oliver. Europe Region special section: Welcome. Communications of the ACM, 65(4):32–34, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3514188.

Kagan:2022:ERS

[687] Dima Kagan, Michael Fire, and Galit Fuhrmann Alpert. Europe Region special section: Hot topics: Trends in computer science research within European countries. Communications of the ACM, 65(4):36–37, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3514089.

Steimle:2022:ERS

[688] Jürgen Steimle. Europe Region special section: Hot topics: On-skin computing. Communications of the ACM, 65(4):38–39, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3511668.

Smallwood:2022:ERS

[689] Thomas R. C. Smallwood, Véronique Lefebvre, and Linus Bengtsson. Europe Region special section: Hot topics: Mobile phone usage data for disaster response. Communications of the ACM, 65(4):40–41, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3512998.

Dobrica:2022:ERS

[690] Liliana Dobrica. Europe Region special section: Hot topics: Robotic process automation platform UiPath. Communications of the ACM, 65(4):42–43, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3511667.

Otto:2022:ERS

[691] Boris Otto. Europe Region special section: Hot topics: A federated infrastructure for European data spaces. Communications of the ACM, 65(4):44–45, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3512341.

Alam:2022:ERS

[692] Sadaf R. Alam, Javier Bartolome, Michele Carpene, Kalle Happonen, Jacques-Charles Lafoucriere, and Dirk Pleiter. Europe Region special section: Hot topics: Fenix: a Pan-European federation of supercomputing and cloud e-infrastructure services. Communications of the ACM, 65(4):46–47, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3511802.

Fettweis:2022:ERS

[693] Gerhard P. Fettweis and Holger Boche. Europe Region special section: Hot topics: On 6G and trustworthiness. Communications of the ACM, 65(4): 48–49, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3512996.

Kappel:2022:ERS

[694] Gertrude Kappel, Christian Brecher, Matthias Brockmann, and István Koren. Europe Region special section: Hot topics: Internet of Production: entering phase two of Industry 4.0. Communications of the ACM, 65(4):50–51, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3514093.

Perino:2022:ERS

[695] Diego Perino, Kleomenis Katevas, Andra Lutu, Eduard Marin, and Nicolas Kourtellis. Europe Region special section: Hot topics: Privacy-preserving AI for future networks. Communications of the ACM, 65(4):52–53, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512343.

Curry:2022:ERS

[696] Edward Curry, Fredrik Heintz, Morten Irgens, Arnold W. M. Smeulders, and Stefano Stramigioli. Europe Region special section: Hot topics: Partnership on AI, data, and robotics. Communications of the ACM, 65(4):54–55, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3513000.

Hochreiter:2022:ERS

[697] Sepp Hochreiter. Europe Region special section: Hot topics: Toward a broad AI. Communications of the ACM, 65(4):56–57, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512715.

Schmeck:2022:ERS

[698] Hartmut Schmeck, Antonello Monti, and Veit Hagenmeyer. Europe Region special section: Big trends: Energy informatics: key elements for tomorrow’s energy system. Communications of the ACM, 65(4):58–63, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3511666.

Middleton:2022:ERS

[699] Stuart E. Middleton, Emmanuel Letouzé, Ali Hossaini, and Adriane Chapman. Europe Region special section: Big trends: Trust, regulation, and human-in-the-loop AI: within the European region. Communications of the ACM, 65(4):64–68, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3511597.

Noia:2022:ERS

[700] Tommaso Di Noia, Nava Tintarev, Panagiota Fatourou, and Markus Schedl. Europe Region special section: Big trends: Recommender systems under European AI regulations. Communications of the ACM, 65(4):69–73, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512728.

Peleg:2022:ERS

[701] Mor Peleg, Yuval Shahar, and Silvana Quaglini. Europe Region special section: Big trends: MobiGuide: guiding clinicians and chronic patients anytime, anywhere. Communications of the ACM, 65(4):74–79, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3511596.

vanderAalst:2022:ERS

[702] Wil van der Aalst. Europe Region special section: Big trends: European leadership in process management. Communications of the ACM, 65(4):80–83, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3511595.

Ali:2022:ERS

[703] Shaukat Ali, Tao Yue, and Rui Abreu. Europe Region special section: Big trends: When software engineering meets quantum computing. Communications of the ACM, 65(4):84–88, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512340.

Bonawitz:2022:FLP

[704] Kallista Bonawitz, Peter Kairouz, Brendan Mcmahan, and Daniel Ramage. Federated learning and privacy. Communications of the ACM, 65(4):90–97, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3500240.

Furini:2022:DTA

[705] Marco Furini, Ombretta Gaggi, Silvia Mirri, Manuela Montangero, Elvira Pelle, Francesco Poggi, and Catia Prandi. Digital twins and artificial intelligence: as pillars of personalized learning models. Communications of the ACM, 65(4):98–104, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3478281.

Fresneda:2022:MSE

[706] Jorge Fresneda, Jeremy Hui, and Chelsey Hill. Market segmentation in the emoji era. Communications of the ACM, 65(4):105–112, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3478282.

Petre:2022:ECB

[707] Marian Petre. Exploring cognitive bias ‘in the wild’: technical perspective. Communications of the ACM, 65(4): 114, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3517215.

Chattopadhyay:2022:CBS

[708] Souti Chattopadhyay, Nicholas Nelson, Audrey Au, Natalia Morales, Christopher Sanchez, Rahul Pandita, and Anita Sarma. Cognitive biases in software development. Communications of the ACM, 65(4):115–122, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3517217.

Cudre-Mauroux:2022:LSC

[709] Philippe Cudré-Mauroux. Leveraging social context for fake news detection: technical perspective. Communications of the ACM, 65(4):123, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3517213.

Nguyen:2022:FLS

[710] Van-Hoang Nguyen, Kazunari Sugiyama, Preslav Nakov, and Min-Yen Kan. FANG: leveraging social context for fake news detection using graph representation. Communications of the ACM, 65(4):124–132, April 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3517214.

Shasha:2022:LBB

[711] Dennis Shasha. Last byte: Beating the house. Communications of the ACM, 65(4):136–ff, April 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3517220.

Chien:2022:ELFa

[712] Andrew A. Chien. Editor’s letter: Four aspirations for ACM 2032. Communications of the ACM, 65(5):5, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529373.

Hanson:2022:EA

[713] Vicki L. Hanson and Jennifer T. Chayes. Editorial: ACM at 75. Communications of the ACM, 65(5):6–7, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3527143.

Vardi:2022:DWT

[714] Moshe Y. Vardi. Departments: War and tech (and ACM). Communications of the ACM, 65(5):9, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3528570.

Munawar:2022:CPC

[715] Asim Munawar. Career paths in computing: How a shopping mall trip inspired me to work in neurosymbolic AI. Communications of the ACM, 65(5):11, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3528571.

Delman:2022:AOT

[716] Scott E. Delman. ACM Open: Thanks for the progress to date! Communications of the ACM, 65(5):12–13, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529750.

Staff:2022:AEA

[717] CACM Staff. ACM’s election: ACM’s 2022 general election: please take this opportunity to vote. Communications of the ACM, 65(5):15–23, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3527144.

Hong:2022:MTC

[718] Jason Hong. Modern tech can’t shield your secret identity. Communications of the ACM, 65(5):24–25, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3524013.

Edwards:2022:NNN

[719] Chris Edwards. News: Neural networks learn to speed up simulations. Communications of the ACM, 65(5):27–29, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3524015.

Mone:2022:NRR

[720] Gregory Mone. News: Raising robovoices. Communications of the ACM, 65(5):30–31, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3524019.

Kirkpatrick:2022:NAI

[721] Keith Kirkpatrick. News: Artificial intelligence and mental health. Communications of the ACM, 65(5):32–34, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3524017.

Reid:2022:LTT

[722] Blake Reid. Law and technology: Two paths for digital disability law. Communications of the ACM, 65(5):36–38, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3527201.

Wroclawski:2022:SCI

[723] John Wroclawski and Terry Benzel. Security: Cybersecurity as illuminator for the future of computing research. Communications of the ACM, 65(5):39–41, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3527202.

Yadav:2022:ETJ

[724] Aman Yadav, Marie Heath, and Anne Drew Hu. Education: Toward justice in computer science through community, criticality, and citizenship. Communications of the ACM, 65(5):42–44, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3527203.

vonHanxleden:2022:VIV

[725] Reinhard von Hanxleden. Viewpoint: Information: I’ vs. ‘we’ vs. ‘they’. Communications of the ACM, 65(5):45–47, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3491205.

Xin:2022:ITC

[726] Reynold Xin, Wes McKinney, Alan Gates, and Chris McCubbin. It takes a community — : the open source challenge. Communications of the ACM, 65(5):48–55, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3498568.

Videla:2022:MCC

[727] Alvaro Videla. Meaning and context in computer programs. Communications of the ACM, 65(5):56–58, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3498850.

Seshia:2022:ECP

[728] Sanjit A. Seshia. Explorations in cyberphysical systems education. Communications of the ACM, 65(5):60–69, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3490442.

Cox:2022:GPL

[729] Russ Cox, Robert Griesemer, Rob Pike, Ian Lance Taylor, and Ken Thompson. The Go programming language and environment. Communications of the ACM, 65(5):70–78, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3488716.

Park:2022:DDE

[730] Jiyong Park and Jongho Kim. A data-driven exploration of the race between human labor and machines in the 21st century. Communications of the ACM, 65(5):79–87, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3488376.

Eslami:2022:AIS

[731] Mohammed Eslami, Aaron Adler, Rajmonda S. Caceres, Joshua G. Dunn, Nancy Kelley-Loughnane, Vanessa A. Varaljay, and Hector Garcia Martin. Artificial intelligence for synthetic biology. Communications of the ACM, 65(5):88–97, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3500922.

Lohmann:2022:WIO

[732] Daniel Lohmann. ’What is the ideal operating system?’: technical perspective. Communications of the ACM, 65 (5):100, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3524299.

Kuo:2022:SCH

[733] Hsuan-Chi Kuo, Jianyan Chen, Sibin Mohan, and Tianyin Xu. Set the configuration for the heart of the OS: on the practicality of operating system kernel debloating. Communications of the ACM, 65(5):101–109, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3524301.

Soljanin:2022:BAL

[734] Emina Soljanin. Balancing at all loads: technical perspective. Communications of the ACM, 65(5):110, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3524292.

Mallick:2022:RCN

[735] Ankur Mallick, Malhar Chaudhari, Utsav Sheth, Ganesh Palanikumar, and Gauri Joshi. Rateless codes for nearperfect load balancing in distributed matrix–vector multiplication. Communications of the ACM, 65(5):111–118, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3524298.

Clegg:2022:LBAb

[736] Brian Clegg. Last byte: Agent provocateur. Communications of the ACM, 65(5):120–ff, May 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3517334.

Chien:2022:ELFb

[737] Andrew A. Chien. Editor’s letter: Five years as Editor-in-Chief of Communications. Communications of the ACM, 65(6):5, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3533672.

Kotsis:2022:EOA

[738] Gabriele Kotsis and Vicki L. Hanson. Editorial: Our ACM community. Communications of the ACM, 65(6):7, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3533677.

Cerf:2022:CDS

[739] Vinton G. Cerf. Cerf’s up: Digital synergy. Communications of the ACM, 65(6):9, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3534934.

Staff:2022:LEM

[740] CACM Staff. Letters to the editor: More on computing’s divided future. Communications of the ACM, 65(6):10–11, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3532630.

Sukhov:2022:RMI

[741] Andrei Sukhov. The role of math in IT education. Communications of the ACM, 65(6):14–15, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3530685.

Savage:2022:NAI

[742] Neil Savage. News: Always improving performance. Communications of the ACM, 65(6):16–18, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3530689.

Monroe:2022:NDU

[743] Don Monroe. News: A deeper understanding of deep learning. Communications of the ACM, 65(6):19–20, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3530686.

Kugler:2022:NAL

[744] Logan Kugler. News: Addressing labor shortages with automation. Communications of the ACM, 65(6):21–23, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3530687.

Greengard:2022:NIC

[745] Samuel Greengard. News: Immersion cooling heats up. Communications of the ACM, 65(6):24–26, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3530688.

Denning:2022:PII

[746] Peter J. Denning. The profession of IT: Involvement and detachment. Communications of the ACM, 65(6):28–31, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3532632.

Neumann:2022:IRT

[747] Peter G. Neumann. Inside risks: Toward total-system trustworthiness. Communications of the ACM, 65(6): 32–35, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3532631.

Neville-Neil:2022:KVP

[748] George Neville-Neil. Kode vicious: The planning and care of data. Communications of the ACM, 65(6):36–37, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3532633.

Knowles:2022:VOH

[749] Bran Knowles, Kelly Widdicks, Gordon Blair, Mike Berners-Lee, and Adrian Friday. Viewpoint: Our house is on fire. Communications of the ACM, 65(6):38–40, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3503916.

Kamp:2022:SIS

[750] Poul-Henning Kamp. The software industry is still the problem. Communications of the ACM, 65(6):42–43, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3486590.

Shortridge:2022:LAN

[751] Kelly Shortridge and Ryan Petrich. Lamboozling attackers: a new generation of deception. Communications of the ACM, 65(6):44–53, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3498578.

Crusoe:2022:MI

[752] Michael R. Crusoe, Sanne Abeln, Alexandru Iosup, Peter Amstutz, John Chilton, Nebojsa Tijanić, Hervé Ménager, Stian Soiland-Reyes, Bogdan Gavrilović, Carole Goble, and The Cwl Community. Methods included. Communications of the ACM, 65(6):54–63, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3486897.

Stoyanovich:2022:RDM

[753] Julia Stoyanovich, Serge Abiteboul, Bill Howe, H. V. Jagadish, and Sebastian Schelter. Responsible data management. Communications of the ACM, 65(6):64–74, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3488717.

Shah:2022:CEC

[754] Nihar B. Shah. Challenges, experiments, and computational solutions in peer review. Communications of the ACM, 65(6):76–87, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3528086.

Navarro:2022:CPB

[755] Gonzalo Navarro. The compression power of the BWT: technical perspective. Communications of the ACM, 65(6):90, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3531443.

Kempa:2022:RBW

[756] Dominik Kempa and Tomasz Kociumaka. Resolution of the Burrows– Wheeler Transform conjecture. Communications of the ACM, 65(6):91–98, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3531445.

Bigham:2022:CWI

[757] Jeffrey P. Bigham. Computation where the (inter)action is: technical perspective. Communications of the ACM, 65(6):99, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3531446.

Jain:2022:SDL

[758] Dhruv Jain, Hung Ngo, Pratyush Patel, Steven Goodman, Khoa Nguyen, Rachel Grossman-Kahn, Leah Findlater, and Jon Froehlich. Sound-Watch: deep learning for sound accessibility on smartwatches. Communications of the ACM, 65(6):100–108, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3531447.

Hoffmann:2022:LBL

[759] Leah Hoffmann. Last byte: Learning new things and avoiding obstacles. Communications of the ACM, 65(6):112–ff, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3530690.

Larus:2022:ELC

[760] James Larus. Editor’s letter: CACM community. Communications of the ACM, 65(7):5, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535269.

Matthews:2022:EEC

[761] Jeanna Matthews. Editorial: Embracing critical voices. Communications of the ACM, 65(7):7, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535268.

Fincher:2022:EEQ

[762] Sally Fincher and Kathi Fisler. Editorial: Exploring questions and answers in computer science education. Communications of the ACM, 65(7):9, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3535270.

Vardi:2022:DAM

[763] Moshe Y. Vardi. Departments: An association of the members, by the members, for the members. Communications of the ACM, 65(7):11, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3539619.

Baquero:2022:DRF

[764] Carlos Baquero and Rosa Cabecinhas. The dynamics of remembering and forgetting. Communications of the ACM, 65(7):12–13, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3535190.

Monroe:2022:NBP

[765] Don Monroe. News: Building a practical quantum computer. Communications of the ACM, 65(7):15–17, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3535191.

Greengard:2022:NBI

[766] Samuel Greengard. News: Brain implants get real. Communications of the ACM, 65(7):18–20, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535193.

Shein:2022:NUM

[767] Esther Shein. News: Using makeup to block surveillance. Communications of the ACM, 65(7):21–23, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3535192.

Samuelson:2022:LSA

[768] Pamela Samuelson. Legally speaking: Apple’s challenge to virtualization software. Communications of the ACM, 65(7):24–26, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3538636.

Yang:2022:EBD

[769] Longqi Yang, David Holtz, Sonia Jaffe, and Siddharth Suri. Economic and business dimensions: The future of information work. Communications of the ACM, 65(7):27–29, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3538638.

Cranor:2022:PCM

[770] Lorrie Faith Cranor. Privacy: Cookie monster. Communications of the ACM, 65(7):30–32, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3538639.

Mitzenmacher:2022:VAP

[771] Michael Mitzenmacher and Sergei Vassilvitskii. Viewpoint: Algorithms with predictions. Communications of the ACM, 65(7):33–35, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3528087.

Gilbert:2022:VWM

[772] Juan E. Gilbert, Stephanie Ludi, David A. Patterson, and Lisa M. Smith. Viewpoint: Words matter. Communications of the ACM, 65(7):36, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3539488.

Vachon:2022:KK

[773] Phil Vachon. The keys to the kingdom. Communications of the ACM, 65(7):38–41, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3511664.

Kamp:2022:STC

[774] Poul-Henning Kamp. Surveillance too cheap to meter. Communications of the ACM, 65(7):42–44, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3511661.

Seshia:2022:TVA

[775] Sanjit A. Seshia, Dorsa Sadigh, and S. Shankar Sastry. Toward verified artificial intelligence. Communications of the ACM, 65(7):46–55, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3503914.

Li:2022:LMP

[776] Hang Li. Language models: past, present, and future. Communications of the ACM, 65(7):56–63, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3490443.

Bright:2022:WSS

[777] Curtis Bright, Ilias Kotsireas, and Vijay Ganesh. When satisfiability solving meets symbolic computation. Communications of the ACM, 65(7):64–72, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3500921.

Vandin:2022:TPE

[778] Fabio Vandin. Technical perspective: Evaluating sampled metrics is challenging. Communications of the ACM, 65(7):74, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535334.

Krichene:2022:SMI

[779] Walid Krichene and Steffen Rendle. On sampled metrics for item recommendation. Communications of the ACM, 65(7):75–83, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535335.

Howe:2022:TPV

[780] Bill Howe. Technical perspective: Visualization search: from sketching to natural language. Communications of the ACM, 65(7):84, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535336.

Siddiqui:2022:EQA

[781] Tarique Siddiqui, Paul Luh, Zesheng Wang, Karrie Karahalios, and Aditya G. Parameswaran. Expressive querying for accelerating visual analytics. Communications of the ACM, 65(7):85–94, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3535337.

Shasha:2022:LBE

[782] Dennis Shasha. Last byte: Exclusivity probes. Communications of the ACM, 65(7):96–ff, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3538640.

Ioannidis:2022:ELM

[783] Yannis Ioannidis. Editor’s letter: To the members of ACM. Communications of the ACM, 65(8):5, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3546595.

Monaco:2022:CCB

[784] Federico Monaco. Cerf’s up: Color blind accessibility manifesto. Communications of the ACM, 65(8):7, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543881.

Meil:2022:ATH

[785] Doug Meil. Advancing in the technical hierarchy. Communications of the ACM, 65(8):8–9, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3542813.

Krakovsky:2022:NFF

[786] Marina Krakovsky. News: Formalizing fairness. Communications of the ACM, 65(8):11–13, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3542815.

Kugler:2022:NCU

[787] Logan Kugler. News: Crossing the uncanny valley. Communications of the ACM, 65(8):14–15, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542817.

Shein:2022:NNL

[788] Esther Shein. News: Neurotechnology and the law. Communications of the ACM, 65(8):16–18, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542816.

Brodley:2022:EWU

[789] Carla E. Brodley. Education: Why universities must resist GPA-based enrollment caps in the face of surging enrollments. Communications of the ACM, 65(8):20–22, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3544547.

Neville-Neil:2022:KVW

[790] George V. Neville-Neil. Kode Vicious: When should a black box be transparent? Communications of the ACM, 65(8):23–24, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3544550.

Shustek:2022:IID

[791] Len Shustek. Interview: An interview with Dana Scott. Communications of the ACM, 65(8):25–29, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3544551.

Parashar:2022:VTS

[792] Manish Parashar, Amy Friedlander, Erwin Gianchandani, and Margaret Martonosi. Viewpoint: Transforming science through cyberinfrastructure. Communications of the ACM, 65(8):30–32, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3507694.

Mike:2022:VCT

[793] Koby Mike, Noa Ragonis, Rinat B. Rosenberg-Kima, and Orit Hazzan. Viewpoint: Computational thinking in the era of data science. Communications of the ACM, 65(8):33–35, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3545109.

Mattioli:2022:FCC

[794] Michael Mattioli. FPGAs in client compute hardware. Communications of the ACM, 65(8):36–42, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3546035.

Chen:2022:IML

[795] Valerie Chen, Jeffrey Li, Joon Sik Kim, Gregory Plumb, and Ameet Talwalkar. Interpretable machine learning: moving from mythos to diagnostics. Communications of the ACM, 65(8):43–50, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3546036.

Gyongyosi:2022:AQI

[796] Laszlo Gyongyosi and Sandor Imre. Advances in the quantum Internet. Communications of the ACM, 65 (8):52–63, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3524455.

Wang:2022:DC

[797] Yihong Wang, Konstantinos Papangelis, Ioanna Lykourentzou, Vassilis-Javed Khan, Michael Saker, Yong Yue, and Jonathan Grudin. The dawn of crowdfarms. Communications of the ACM, 65(8):64–70, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3490698.

Abadi:2022:SRD

[798] Daniel Abadi, Anastasia Ailamaki, David Andersen, Peter Bailis, Magdalena Balazinska, Philip A. Bernstein, Peter Boncz, Surajit Chaudhuri, Alvin Cheung, Anhai Doan, Luna Dong, Michael J. Franklin, Juliana Freire, Alon Halevy, Joseph M. Hellerstein, Stratos Idreos, Donald Kossmann, Tim Kraska, Sailesh Krishnamurthy, Volker Markl, Sergey Melnik, Tova Milo, C. Mohan, Thomas Neumann, Beng Chin Ooi, Fatma Ozcan, Jignesh Patel, Andrew Pavlo, Raluca Popa, Raghu Ramakrishnan, Christopher Re, Michael Stonebraker, and Dan Suciu. The Seattle report on database research. Communications of the ACM, 65(8):72–79, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3524284.

Zhang:2022:CDS

[799] Qin Zhang. Can data structures treat us fairly?: technical perspective. Communications of the ACM, 65(8):82, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543843.

Aumuller:2022:SNN

[800] Martin Aumüller, Sariel Har-Peled, Sepideh Mahabadi, Rasmus Pagh, and Francesco Silvestri. Sampling near neighbors in search for fairness. Communications of the ACM, 65(8):83–90, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543667.

Zilberman:2022:HLE

[801] Noa Zilberman. hXDP: light and efficient packet processing offload: technical perspective. Communications of the ACM, 65(8):91, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543844.

Brunella:2022:HES

[802] Marco Spaziani Brunella, Giacomo Belocchi, Marco Bonola, Salvatore Pontarelli, Giuseppe Siracusano, Giuseppe Bianchi, Aniello Cammarano, Alessandro Palumbo, Luca Petrucci, and Roberto Bifulco. hXDP: efficient software packet processing on FPGA NICs. Communications of the ACM, 65(8):92–100, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3543668.

Logan:2022:LBL

[803] E. C. Logan. Last byte: The Luce Goose: a space-traveling bail-bond agent and the ship’s resourceful AI chase fugitives across the galaxy. Communications of the ACM, 65(8):104–ff, August 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542853.

Vardi:2022:DTD

[804] Moshe Y. Vardi. Departments: Technology and democracy. Communications of the ACM, 65(9):5, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3550062.

Sorensen:2022:CPC

[805] Kristian Sørensen. Career paths in computing: The making of an IT strategy consultant. Communications of the ACM, 65(9):7, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3546937.

Kambhampati:2022:CNA

[806] Subbarao Kambhampati. Changing the nature of AI research. Communications of the ACM, 65(9):8–9, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3546954.

Edwards:2022:NCM

[807] Chris Edwards. News: Competition makes big datasets the winners. Communications of the ACM, 65(9): 11–13, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3546955.

Kirkpatrick:2022:NR

[808] Keith Kirkpatrick. News: The road to 6G. Communications of the ACM, 65 (9):14–16, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3546959.

Kugler:2022:NHA

[809] Logan Kugler. News: How AI is driving the esports boom. Communications of the ACM, 65(9):17–18, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3546956.

Guadamuz:2022:LTA

[810] Andres Guadamuz. Law and technology: These are not the apes you are looking for. Communications of the ACM, 65(9):20–22, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3548761.

Kuehn:2022:SSL

[811] Andreas Kuehn. Security: Security by labeling. Communications of the ACM, 65(9):23–25, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3548762.

Denning:2022:PIA

[812] Peter Denning and Roland Ibbett. The profession of IT: The Atlas milestone. Communications of the ACM, 65(9): 26–29, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3548781.

Dally:2022:PCM

[813] William Dally. Point/counterpoint: On the model of computation: point. Communications of the ACM, 65(9): 30–32, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3548783.

Vishkin:2022:PCM

[814] Uzi Vishkin. Point/counterpoint: On the model of computation: counterpoint. Communications of the ACM, 65(9):32–34, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3548784.

Maraninchi:2022:VLU

[815] Florence Maraninchi. Viewpoint: Let us not put all our eggs in one basket. Communications of the ACM, 65 (9):35–37, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3528088.

Gazis:2022:M

[816] Alexandros Gazis and Eleftheria Katsiri. Middleware 101. Communications of the ACM, 65(9):38–42, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3546958.

Cobbs:2022:PP

[817] Archie L. Cobbs. Persistence programming. Communications of the ACM, 65(9):43–47, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3546957.

Troncoso:2022:DDP

[818] Carmela Troncoso, Dan Bogdanov, Edouard Bugnion, Sylvain Chatel, Cas Cremers, Seda Gürses, Jean-Pierre Hubaux, Dennis Jackson, James R. Larus, Wouter Lueks, Rui Oliveira, Mathias Payer, Bart Preneel, Apostolos Pyrgelis, Marcel Salathé, Theresa Stadler, and Michael Veale. Deploying decentralized, privacy-preserving proximity tracing. Communications of the ACM, 65(9):48–57, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3524107.

Lamport:2022:DBB

[819] Leslie Lamport. Deconstructing the bakery to build a distributed state machine. Communications of the ACM, 65(9):58–66, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3513259.

Latah:2022:WSB

[820] Majd Latah and Kubra Kalkan. When SDN and blockchain shake hands. Communications of the ACM, 65(9): 68–78, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3500920.

Dressler:2022:PLR

[821] Falko Dressler. Physical layer resilience through deep learning in software radios: technical perspective. Communications of the ACM, 65(9):82, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3547130.

Restuccia:2022:PWR

[822] Francesco Restuccia and Tommaso Melodia. Polymorphic wireless receivers. Communications of the ACM, 65(9):83–91, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3547131.

Christin:2022:ESM

[823] Nicolas Christin. The effectiveness of security measures: technical perspective. Communications of the ACM, 65(9):92, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3547132.

DeKoven:2022:MSP

[824] Louis F. DeKoven, Audrey Randall, Ariana Mirian, Gautam Akiwate, Ansel Blume, Lawrence K. Saul, Aaron Schulman, Geoffrey M. Voelker, and Stefan Savage. Measuring security practices. Communications of the ACM, 65(9):93–102, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3547133.

Hoffmann:2022:LBA

[825] Leah Hoffmann. Last byte: Advancing the ability of robots to help. Communications of the ACM, 65(9): 104–ff, September 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3546960.

Chien:2022:ELC

[826] Andrew A. Chien. Editor’s letter: Computing’s grand challenge for sustainability. Communications of the ACM, 65(10):5, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3559163.

Cerf:2022:CQC

[827] Vinton G. Cerf. Cerf’s up: Quis custodiet ipsos custodes? Communications of the ACM, 65(10):7, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3558590.

Romanov:2022:DIC

[828] Aleksandr Romanov. Driving an innovation contest into crisis. Communications of the ACM, 65(10):8–9, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3554917.

Monroe:2022:NNA

[829] Don Monroe. News: Neurosymbolic AI. Communications of the ACM, 65 (10):11–13, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3554918.

Garfinkel:2022:NMJ

[830] Simson Garfinkel and Eugene H. Spafford. News: In memoriam: Juris Hartmanis 1928–2022. Communications of the ACM, 65(10):14–15, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3559705.

Greengard:2022:NHM

[831] Samuel Greengard. News: Hidden malware ratchets up cybersecurity risks. Communications of the ACM, 65(10):16–18, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3554925.

Shein:2022:NAA

[832] Esther Shein. News: Applied AI teaches handwriting. Communications of the ACM, 65(10):19–20, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3554919.

Cusumano:2022:TSMb

[833] Michael A. Cusumano. Technology strategy and management: Data platforms and network effects. Communications of the ACM, 65(10):22–24, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3555833.

Neville-Neil:2022:KVS

[834] George Neville-Neil. Kode vicious: Securing the company jewels. Communications of the ACM, 65(10):25–26, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3555844.

Barik:2022:VSS

[835] Titus Barik, Sumit Gulwani, and Mario Juarez. Viewpoint: Storytelling and science. Communications of the ACM, 65(10):27–30, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3526100.

Bockmon:2022:VWY

[836] Ryan Bockmon and Stephen Cooper. Viewpoint: What’s your placebo? Communications of the ACM, 65 (10):31–33, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3528085.

Butler:2022:WMT

[837] Jenna Butler and Catherine Yeh. Walk a mile in their shoes. Communications of the ACM, 65(10):34–41, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561989.

Kamp:2022:LAS

[838] Poul-Henning Kamp. Linear address spaces. Communications of the ACM, 65(10):42–44, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3561991.

Ramachandran:2022:CRA

[839] Umakishore Ramachandran and Zvi Galil. Creating a revolutionary academic program. Communications of the ACM, 65(10):46–56, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3503915.

Deshpande:2022:AQC

[840] Advait Deshpande. Assessing the quantum-computing landscape. Communications of the ACM, 65(10):57– 65, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3524109.

Rosenfeld:2022:SYC

[841] Ariel Rosenfeld and Oleg Maksimov. Should young computer scientists stop collaborating with their doctoral advisors? Communications of the ACM, 65(10):66–72, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3529089.

Markopoulou:2022:TPT

[842] Athina Markopoulou. Technical perspective: Traffic classification in the era of deep learning. Communications of the ACM, 65(10):74, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3556640.

Akbari:2022:TCI

[843] Iman Akbari, Mohammad A. Salahuddin, Leni Ven, Noura Limam, Raouf Boutaba, Bertrand Mathieu, Stephanie Moteau, and Stephane Tuffin. Traffic classification in an increasingly encrypted web. Communications of the ACM, 65(10):75–83, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3559439.

Zhang:2022:TPA

[844] Yang Zhang. Technical perspective: Achieve big with devices that track small things. Communications of the ACM, 65(10):84, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3556638.

Parizi:2022:APE

[845] Farshid Salemi Parizi, Eric Whitmire, and Shwetak Patel. AuraRing: precise electromagnetic finger tracking. Communications of the ACM, 65 (10):85–92, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3556639.

Shasha:2022:LBC

[846] Dennis Shasha. Last byte: Card Nim. Communications of the ACM, 65(10):96, October 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3555723.

Vardi:2022:DAL

[847] Moshe Y. Vardi. Departments: Accountability and liability in computing. Communications of the ACM, 65(11):5, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3565237.

Danielescu:2022:CPC

[848] Andreea Danielescu. Career paths in computing: The many shapes of a computer science career. Communications of the ACM, 65(11):7, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3559771.

Staff:2022:LEB

[849] CACM Staff. Letters to the editor: The blood price of unrestricted privacy. Communications of the ACM, 65(11):8–9, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3563965.

Hazzan:2022:BDS

[850] Orit Hazzan, Koby Mike, and Herbert Bruderer. BLOG@CACM: Up with data science, and the first programmer. Communications of the ACM, 65 (11):10–11, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3563107.

Edwards:2022:NGB

[851] Chris Edwards. News: Go backward, save energy. Communications of the ACM, 65(11):13–15, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563108.

Kirkpatrick:2022:NAM

[852] Keith Kirkpatrick. News: Applying the metaverse. Communications of the ACM, 65(11):16–18, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3565470.

Kugler:2022:NIV

[853] Logan Kugler. News: The impact of virtual meetings. Communications of the ACM, 65(11):19–21, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563109.

Samuelson:2022:LSE

[854] Pamela Samuelson. Legally speaking: An emergent legal right to repair electronic devices. Communications of the ACM, 65(11):22–25, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563966.

Cranor:2022:PMA

[855] Lorrie Faith Cranor. Privacy: Mobileapp privacy nutrition labels missing key ingredients for success. Communications of the ACM, 65(11):26–28, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3563967.

Barr:2022:EWM

[856] Valerie Barr. Education: What must all post-secondary students learn about computing? Communications of the ACM, 65(11):29–31, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563968.

Markoff:2022:VRS

[857] John Markoff. Viewpoint: Rethinking Silicon Valley: Stewart Brand’s lost journal. Communications of the ACM, 65(11):32–34, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3530988.

White:2022:VIF

[858] Ryen W. White. Viewpoint: Intelligent futures in task assistance. Communications of the ACM, 65(11): 35–39, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3528083.

Kapoor:2022:IRS

[859] Hemangee Kalpesh Kapoor, Mausam, and Venkatesh Raman. India region special section: Welcome back! Communications of the ACM, 65(11): 40–42, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3564603.

Balakrishnan:2022:IRS

[860] M. Balakrishnan. India region special section: Big trends: Computing and assistive technology solutions for the visually impaired. Communications of the ACM, 65(11):44–47, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3551634.

Veezhinathan:2022:IRS

[861] Kamakoti Veezhinathan. India region special section: Big trends: Building the SHAKTI microprocessor. Communications of the ACM, 65(11):48–51, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3556632.

Darke:2022:IRS

[862] Priyanka Darke, Ravindra Metta, Raveendra Kumar Medicherla, and R. Venkatesh. India region special section: Hot topics: Impactful research and tooling for program correctness. Communications of the ACM, 65 (11):52–53, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3551665.

Warriem:2022:IRS

[863] Jayakrishnan Warriem, Andrew Thangaraj, Madhavan Mukund, and Bharathi Balaji. India region special section: Big trends: Digital interventions at scale: lessons from NPTEL and IIT Madras B.S. degree program. Communications of the ACM, 65(11):54–57, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3550473.

Bhowmick:2022:IRS

[864] Plaban Kumar Bhowmick, Partha Pratim Das, Partha Pratim Chakrabarti, and Debarshi Kumar Sanyal. India region special section: Big trends: National digital library of India: democratizing education in India. Communications of the ACM, 65(11):58–61, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3550480.

Karkare:2022:IRS

[865] Amey Karkare and Purushottam Kar. India region special section: Hot topics: Prutor: an intelligent learning and management system for programming courses. Communications of the ACM, 65(11):62–64, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3550481.

Chaudhuri:2022:IRS

[866] Bidisha Chaudhuri, Meenakshi D’Souza, and Janaki Srinivasan. India region special section: Big trends: Bringing the missing women back: CS education for women in India’s engineering institutions. Communications of the ACM, 65(11):65–67, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3550490.

Balasubramanian:2022:IRS

[867] Vineeth N. Balasubramanian. India region special section: Hot topics: Toward explainable deep learning. Communications of the ACM, 65 (11):68–69, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3550491.

Bhatti:2022:IRS

[868] Iram Tariq Bhatti, Mahum Naseer, Muhammad Shafique, and Osman Hasan. India region special section: Hot topics: A formal approach to identifying the impact of noise on neural networks. Communications of the ACM, 65(11):70–73, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3550492.

Belsare:2022:IRS

[869] Hemant Belsare, Milind Sohoni, Rahul Gokhale, Pooja Prasad, Asim R. P., Chirag Marathe, Parth Gupta, Shubhada Sali, Swapnil Patil, and Sudhanshu Deshmukh. India region special section: Big trends: Computing for climate resilience in agriculture. Communications of the ACM, 65 (11):74–79, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3554928.

Singhal:2022:IRS

[870] Shivangi Singhal, Rishabh Kaushal, Rajiv Ratn Shah, and Ponnurangam Kumaraguru. India region special section: Hot topics: Fake news in India: scale, diversity, solution, and opportunities. Communications of the ACM, 65 (11):80–81, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3550493.

Athreya:2022:IRS

[871] Siva Athreya, Gautam I. Menon, and Rajesh Sundaresan. India region special section: Big trends: COVID-19 modeling for India and a roadmap for the future. Communications of the ACM, 65(11):82–87, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3557798.

Aiswarya:2022:IRS

[872] C. Aiswarya, V. Arvind, and Saket Saurabh. India region special section: Big trends: Theory research in India: 2019–2022. Communications of the ACM, 65(11):88–93, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3550496.

Ravindran:2022:IRS

[873] B. Ravindran, Sunita Sarawagi, and Aditi Jain. India region special section: Big trends: AI and data science centers in top Indian academic institutions. Communications of the ACM, 65 (11):94–97, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3556634.

Mukhopadhyay:2022:IRS

[874] Debdeep Mukhopadhyay. India region special section: Big trends: Cybersecurity in India. Communications of the ACM, 65(11):98–102, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3554910.

Kleppmann:2022:RPC

[875] Mark Kleppmann and Peter Alvaro. Research for practice: convergence. Communications of the ACM, 65(11): 104–106, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3563901.

Falkner:2022:IBN

[876] Matthias Falkner and John Apostolopoulos. Intent-based networking for the enterprise: a modern network architecture. Communications of the ACM, 65(11):108–117, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3538513.

Moskovitch:2022:RMS

[877] Yuval Moskovitch and H. V. Jagadish. Reliability at multiple stages in a data analysis pipeline. Communications of the ACM, 65(11):118–128, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3500923.

Dantu:2022:PLP

[878] Karthik Dantu. The power of lowpower GPS receivers for nanosats: technical perspective. Communications of the ACM, 65(11):132, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3559769.

Narayana:2022:HEE

[879] Sujay Narayana, R. Venkatesha Prasad, Vijay S. Rao, Luca Mottola, and T. V. Prabhakar. Hummingbird: an energy-efficient GPS receiver for small satellites. Communications of the ACM, 65(11):133–140, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3559770.

Clegg:2022:LBN

[880] Brian Clegg. Last byte: The new wizard of Menlo Park. Communications of the ACM, 65(11):144–ff, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3470741.

Cerf:2022:CTB

[881] Vinton G. Cerf. Cerf’s up: On truth and belief. Communications of the ACM, 65(12):5, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3569720.

Baquero:2022:HAG

[882] Carlos Baquero. Is having AI generate text cheating? Communications of the ACM, 65(12):6–7, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3565976.

Monroe:2022:NQC

[883] Don Monroe. News: Quantum computers and the universe. Communications of the ACM, 65(12):9–11, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3565977.

Greengard:2022:NSR

[884] Samuel Greengard. News: Swarm robotics moves forward. Communications of the ACM, 65(12):12–14, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3565979.

Shein:2022:NPP

[885] Esther Shein. News: Preserving the past with immersive technologies. Communications of the ACM, 65 (12):15–17, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3565978.

Denning:2022:PIC

[886] Peter J. Denning and John Arquilla. The profession of IT: The context problem in artificial intelligence. Communications of the ACM, 65(12):18–21, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3567605.

Neville-Neil:2022:KVF

[887] George Neville-Neil. Kode vicious: The four horsemen of an ailing software project. Communications of the ACM, 65(12):22–23, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3567606.

Dubhashi:2022:VCU

[888] Devdatt Dubhashi. Viewpoint: Can universities combat the ‘wrong kind of AI’? Communications of the ACM, 65 (12):24–26, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3522710.

Pentland:2022:VBN

[889] Alex Pentland. Viewpoint: Building a new economy: data, AI, and Web3. Communications of the ACM, 65(12):27–29, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3547659.

Barroso:2022:VGP

[890] L. A. Barroso, T. Choudhury, M. Gupta, O. Olukotun, R. A. Popa, D. Song, and David A. Patterson. Viewpoint: Global perspectives of diversity, equity, and inclusion. Communications of the ACM, 65(12):30–31, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3548454.

Staff:2022:FRF

[891] CACM Staff. FHIR: reducing friction in the exchange of healthcare data. Communications of the ACM, 65 (12):34–41, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3565861.

Kamp:2022:COO

[892] Poul-Henning Kamp. CSRB’s opus one. Communications of the ACM, 65 (12):42–44, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3565860.

Laberge:2022:SPG

[893] Nicholas Laberge, K. Hunter Wapman, Allison C. Morgan, Sam Zhang, Daniel B. Larremore, and Aaron Clauset. Subfield prestige and gender inequality among U.S. computing faculty. Communications of the ACM, 65 (12):46–55, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3535510.

Chen:2022:PCH

[894] Juan Chen, Sheikh Ghafoor, and John Impagliazzo. Producing competent HPC graduates. Communications of the ACM, 65(12):56–65, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3538878.

Dongarra:2022:TLE

[895] Jack J. Dongarra. Turing lecture: The evolution of mathematical software. Communications of the ACM, 65 (12):66–72, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3554977.

Amalfitano:2022:HDJ

[896] Domenico Amalfitano, Ana C. R. Paiva, Alexis Inquel, Luís Pinto, Anna Rita Fasolino, and René Just. How do Java mutation tools differ? Communications of the ACM, 65(12): 74–89, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3526099.

Nowara:2022:SBS

[897] Ewa Nowara, Daniel Mcduff, Ashutosh Sabharwal, and Ashok Veeraraghavan. Seeing beneath the skin with computational photography. Communications of the ACM, 65(12):90–100, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3536421.

Boncz:2022:TPA

[898] Peter Boncz. Technical perspective: The ‘art’ of automatic benchmark extraction. Communications of the ACM, 65(12):104, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3567465.

Deep:2022:DBQ

[899] Shaleen Deep, Anja Gruenheid, Kruthi Nagaraj, Hiro Naito, Jeff Naughton, and Stratis Viglas. DIAMETRICS: benchmarking query engines at scale. Communications of the ACM, 65(12): 105–112, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3567464.

Hoffmann:2022:LBF

[900] Leah Hoffmann. Last byte: Finding an alternate way forward with AI. Communications of the ACM, 65 (12):120–ff, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3565980.

Chien:2023:ELC

[901] Andrew A. Chien. Editor’s letter: Computing divided: How wide the chasm? Communications of the ACM, 66(1):5, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3572994.

Vardi:2023:DHW

[902] Moshe Y. Vardi. Departments: How not to win a tech war. Communications of the ACM, 66(1):7, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3571077.

Zhou:2023:CPC

[903] Michelle Zhou. Career paths in computing: a computer scientist with a biologist’s ambition: Advance humanity. Communications of the ACM, 66(1): 9, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3571064.

Rotenberg:2023:MAF

[904] Marc Rotenberg and Jeremy Roschelle. Making AI fair, and how to use it. Communications of the ACM, 66 (1):10–11, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3570517.

Edwards:2023:NEC

[905] Chris Edwards. News: Error control begins to shape quantum architectures. Communications of the ACM, 66(1):13–15, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3570518.

Savage:2023:NOC

[906] Neil Savage. News: The outlook for crypto. Communications of the ACM, 66(1):16–18, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3570520.

Kugler:2023:NMT

[907] Logan Kugler. News: Making traffic a thing of the past. Communications of the ACM, 66(1):19–20, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3570519.

Garfinkel:2023:NMF

[908] Simson Garfinkel and Eugene H. Spafford. News: In memoriam: Frederick P. Brooks, Jr. 1931–2022. Communications of the ACM, 66(1):21–22, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3572995.

Delaney:2023:NRV

[909] John Delaney. News: Remembering Valérie Issarny. Communications of the ACM, 66(1):23, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3573217.

Cusumano:2023:TSMa

[910] Michael A. Cusumano. Technology strategy and management: From quantum computing to quantum communications. Communications of the ACM, 66(1):24–27, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3571450.

Roberts:2023:LTG

[911] Alexandra J. Roberts. Law and technology: Getting a handle on handles. Communications of the ACM, 66(1):28–30, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3571451.

Massacci:2023:SSU

[912] Fabio Massacci and Giorgio di Tizio. Security: Are software updates useless against advanced persistent threats? Communications of the ACM, 66 (1):31–33, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3571452.

Welsh:2023:VEP

[913] Matt Welsh. Viewpoint: The end of programming. Communications of the ACM, 66(1):34–35, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3570220.

Noy:2023:VWC

[914] Natasha Noy and Carole Goble. Viewpoint: Are we cobblers without shoes?: Making computer science data FAIR.

Communications of the ACM, 66 (1):36–38, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3528574.

Saetra:2023:VAE

[915] Henrik Skaug Sætra, Mark Coeckelbergh, and John Danaher. Viewpoint: The AI ethicist’s dirty hands problem. Communications of the ACM, 66(1):39–41, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3529751.

Eaton:2023:DLP

[916] Brian Eaton, Jeff Stewart, Jon Tedesco, and N. Cihan Tas. Distributed latency profiling through critical path tracing. Communications of the ACM, 66(1):44–51, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3570522.

Alagappan:2023:RPC

[917] Ramnatthan Alagappan and Peter Alvaro. Research for practice: Crash consistency. Communications of the ACM, 66(1):52–54, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3570521.

Lewis:2023:CAM

[918] Ted G. Lewis. The many faces of resilience. Communications of the ACM, 66(1):56–61, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3519262.

Wang:2023:CAA

[919] Luhui Wang, Cong Zhao, Shusen Yang, Xinyu Yang, and Julie Mccann. ACE: Toward applicationcentric, edge-cloud, collaborative intelligence. Communications of the ACM, 66(1):62–73, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3529087.

Chi:2023:CAD

[920] Yuze Chi, Weikang Qiao, Atefeh Sohrabizadeh, Jie Wang, and Jason Cong. Democratizing domain-specific computing. Communications of the ACM, 66(1):74–85, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3524108.

Castaneda:2023:RAL

[921] Armando Castañeda, Sergio Rajsbaum, and Michel Raynal. A linearizability-based hierarchy for concurrent specifications. Communications of the ACM, 66(1):86–97, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3546826.

Conitzer:2023:RHI

[922] Vincent Conitzer, Gillian K. Hadfield, and Shannon Vallor. Research highlights: The impact of auditing for algorithmic bias: Technical perspective. Communications of the ACM, 66(1):100, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3571152.

Raji:2023:RHA

[923] Inioluwa Deborah Raji and Joy Buolamwini. Research highlights: Actionable auditing revisited: Investigating the impact of publicly naming biased performance results of commercial AI products. Communications of the ACM, 66(1):101–108, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3571151.

Shasha:2023:LBM

[924] Dennis Shasha. Last byte: Maximal cocktails. Communications of the ACM, 66(1):112, January 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3571275.

Ioannidis:2023:PMA

[925] Yannis Ioannidis. From the President: To the members of ACM. Communications of the ACM, 66(2):5, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3578649.

Cerf:2023:CQC

[926] Vinton G. Cerf. Cerf’s up: On QR codes and safety. Communications of the ACM, 66(2):7, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3578891.

Staff:2023:LEN

[927] CACM Staff. Letters to the Editor: Neighborhood watch. Communications of the ACM, 66(2):8–11, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3576929.

Mike:2023:WDS

[928] Koby Mike and Orit Hazzan. What is data science? Communications of the ACM, 66(2):12–13, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3575663.

Monroe:2023:NPQ

[929] Don Monroe. News: Post-quantum cryptography. Communications of the ACM, 66(2):15–17, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3575664.

Greengard:2023:NCL

[930] Samuel Greengard. News: Computational linguistics finds its voice. Communications of the ACM, 66(2):18–20, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3575666.

Kirkpatrick:2023:NCA

[931] Keith Kirkpatrick. News: Can AI demonstrate creativity? Communications of the ACM, 66(2):21–23, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3575665.

Ericson:2023:EFW

[932] Barbara Ericson. Education: Four ways to add active learning to computing courses. Communications of the ACM, 66(2):26–29, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3576930.

Neville-Neil:2023:KVE

[933] George V. Neville-Neil. Kode vicious: The elephant in the room. Communications of the ACM, 66(2):30–31, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3576931.

Johnson:2023:CEE

[934] Deborah G. Johnson and Mario Verdicchio. Computing ethics: Ethical AI is not about AI. Communications of the ACM, 66(2):32–34, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3576932.

Isbell:2023:VSE

[935] Charles Isbell, Michael L. Littman, and Peter Norvig. Viewpoint: Software engineering of machine learning systems. Communications of the ACM, 66(2):35–37, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3539783.

Raffel:2023:VBM

[936] Colin Raffel. Viewpoint: Building machine learning models like open source software. Communications of the ACM, 66(2):38–40, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3545111.

Yellin:2023:VPO

[937] Daniel M. Yellin. Viewpoint: The premature obituary of programming. Communications of the ACM, 66(2): 41–44, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3555367.

Gilbert:2023:VAB

[938] Juan E. Gilbert, Jeremy A. Magruder Waisome, and Simone Smarr. Viewpoint: An analysis of black faculty in CS research departments. Communications of the ACM, 66(2):45–47, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3571279.

Bush:2023:Z

[939] Matthew Bush and Atefeh Mashatan. From zero to 100. Communications of the ACM, 66(2):48–55, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3573127.

Loftus:2023:AZT

[940] Michael Loftus, Andrew Vezina, Rick Doten, and Atefeh Mashatan. The arrival of zero trust: What does it mean? Communications of the ACM, 66(2):56–62, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3573129.

Mccauley:2023:CAE

[941] James Mccauley, Scott Shenker, and George Varghese. Contributed articles: Extracting the essential simplicity of the Internet. Communications of the ACM, 66(2):64–74, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3547137.

Baldassarre:2023:CAR

[942] Maria Teresa Baldassarre, Neil Ernst, Ben Hermann, Tim Menzies, and Rahul Yedida. Contributed articles: (Re)Use of research results (is rampant). Communications of the ACM, 66(2):75–81, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3554976.

Reed:2023:CAH

[943] Daniel Reed, Dennis Gannon, and Jack Dongarra. Contributed articles: HPC forecast: Cloudy and uncertain. Communications of the ACM, 66(2):82–90, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3552309.

Shani:2023:RAL

[944] Chen Shani, Jonathan Zarecki, and Dafna Shahaf. Review articles: The lean data scientist: Recent advances toward overcoming the data bottleneck. Communications of the ACM, 66(2):92–102, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3551635.

Vechev:2023:RHT

[945] Martin Vechev. Research highlights: Technical perspective: Beautiful symbolic abstractions for safe and secure machine learning. Communications of the ACM, 66(2):104, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3576893.

Drews:2023:RHP

[946] Samuel Drews, Aws Albarghouthi, and Loris D’Antoni. Research highlights: Proving data-poisoning robustness in decision trees. Communications of the ACM, 66(2):105–113, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3576894.

Bainbridge:2023:LBA

[947] William Sims Bainbridge. Last byte: Aftermath impact. Communications of the ACM, 66(2):116–ff, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3578622.

Vardi:2023:DWR

[948] Moshe Y. Vardi. Departments: Who is responsible around here? Communications of the ACM, 66(3):5, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3580584.

Wolf:2023:CDF

[949] Marty J. Wolf, Joseph A. Konstan, Helena Mentis, Jane Prey, and Harald Störrle. Committee on Disclosure of Findings: After the complaint: What should ACM disclose? Communications of the ACM, 66(3):6–11, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3581640.

Oliver:2023:CPC

[950] Nuria Oliver. Career paths in computing: A philosopher’s daughter navigates a career in AI. Communications of the ACM, 66(3):13, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579857.

Baquero:2023:WEH

[951] Carlos Baquero. What ever happened to peer-to-peer systems? Communications of the ACM, 66(3):14–15, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579633.

Monroe:2023:NNN

[952] Don Monroe. News: Neural networks for drug discovery and design. Communications of the ACM, 66(3):17–19, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579634.

Kirkpatrick:2023:NAS

[953] Keith Kirkpatrick. News: Adding smarts to vending machines drives convenience, efficiency. Communications of the ACM, 66(3):20–22, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3579651.

Kugler:2023:NVI

[954] Logan Kugler. News: Virtual influencers in the real world. Communications of the ACM, 66(3):23–25, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3579635.

Denning:2023:PIM

[955] Peter J. Denning. The profession of IT: A map for innovation. Communications of the ACM, 66(3):28–31, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3581762.

Samuelson:2023:LSLa

[956] Pamela Samuelson. Legally speaking: a legal challenge to algorithmic recommendations. Communications of the ACM, 66(3):32–34, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3581763.

Cranor:2023:PMS

[957] Lorrie Faith Cranor and Hana Habib. Privacy: Metrics for success: Why and how to evaluate privacy choice usability. Communications of the ACM, 66(3):35–37, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3581764.

Gillespie:2023:VDE

[958] Alexandra Gillespie. Viewpoint: Designing an ethical tech developer. Communications of the ACM, 66(3):38–40, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3545110.

Woods:2023:VTP

[959] Daniel W. Woods. Viewpoint: a turning point for cyber insurance. Communications of the ACM, 66(3):41–44, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3545795.

Auer:2023:MPL

[960] Raphael Auer, Rainer Böhme, Jeremy Clark, and Didem Demirag. Mapping the privacy landscape for central bank digital currencies. Communications of the ACM, 66(3):46–53, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579316.

Limoncelli:2023:SYO

[961] Thomas A. Limoncelli. Split your overwhelmed teams. Communications of the ACM, 66(3):54–56, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579315.

Berger:2023:CAA

[962] Sara Berger and Francesca Rossi. Contributed articles: AI and neurotechnology: Learning from AI ethics to address an expanded ethics landscape. Communications of the ACM, 66(3):58–68, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529088.

Tsaih:2023:CAA

[963] Rua-Huan Tsaih, Hsin-Lu Chang, Chih-Chun Hsu, and David C. Yen. Contributed articles: The AI techstack model. Communications of the ACM, 66(3):69–77, March 2023. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3568026.

Chancellor:2023:CAT

[964] Stevie Chancellor. Contributed articles: Toward practices for humancentered machine learning. Communications of the ACM, 66(3):78–85, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3530987.

Ragan-Kelley:2023:RHR

[965] Jonathan Ragan-Kelley. Research highlights: Reconsidering the design of user-schedulable languages: Technical perspective. Communications of the ACM, 66(3):88, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3580370.

Hagedorn:2023:RHA

[966] Bastian Hagedorn, Johannes Lenfers, Thomas Koehler, Xueying Qin, Sergei Gorlatch, and Michel Steuwer. Research highlights: Achieving high performance the functional way: Expressing high-performance optimizations as rewrite strategies. Communications of the ACM, 66(3):89–97, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3580371.

Hoffmann:2023:LBC

[967] Leah Hoffmann. Last byte: Considering the impact of technology on society. Communications of the ACM, 66(3):100–ff, March 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3579652.

Cerf:2023:CSC

[968] Vinton G. Cerf. Cerf’s up: Scaffolding as catalyst. Communications of the ACM, 66(4):5, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3586159.

Marcus:2023:HBA

[969] Gary Marcus. Hoping for the best as AI evolves. Communications of the ACM, 66(4):6–7, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3583078.

Monroe:2023:NAI

[970] Don Monroe. News: Artificial intelligence for materials discovery. Communications of the ACM, 66(4):9–11, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3583080.

Greengard:2023:NAR

[971] Samuel Greengard. News: AI rewrites coding. Communications of the ACM, 66(4):12–14, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3583083.

Shein:2023:NLC

[972] Esther Shein. News: … and the computer plays along. Communications of the ACM, 66(4):15–17, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3583082.

Hall:2023:APF

[973] Wendy Hall, Divesh Srivastava, and Scott E. Delman. ACM publications finances: ACM publications finances for 2021. Communications of the ACM, 66(4):18–19, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3586184.

Sako:2023:TSM

[974] Mari Sako. Technology strategy and management: Contracting for artificial intelligence. Communications of the ACM, 66(4):20–23, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3584707.

Neville-Neil:2023:KVA

[975] George V. Neville-Neil. Kode vicious: All sliders to the right. Communications of the ACM, 66(4):24–25, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3584711.

VanAlstyne:2023:EBD

[976] Marshall Van Alstyne, Michael D. Smith, and Herbert Lin. Economic and business dimensions: Improving section 230, preserving democracy, and protecting free speech. Communications of the ACM, 66(4):26–28, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3584710.

Ott:2023:VWR

[977] David Ott, Kenny Paterson, and Dennis Moreau. Viewpoint: Where is the research on cryptographic transition and agility? Communications of the ACM, 66(4):29–32, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3567825.

Linkov:2023:VCE

[978] Igor Linkov, Alexandre Ligo, Kelsey Stoddard, Beatrice Perez, Andrew Strelzoffx, Emanuele Bellini, and Alexander Kott. Viewpoint: Cyber efficiency and cyber resilience. Communications of the ACM, 66(4):33–37, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3549073.

Patterson:2023:VFC

[979] David A. Patterson. Viewpoint: Faculty can save lives of the elderly. Communications of the ACM, 66(4):38, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3555368.

Navas:2023:OVC

[980] Jorge A. Navas and Ashish Gehani. OCCAM-v2: Combining static and dynamic analysis for effective and efficient whole-program specialization. Communications of the ACM, 66(4):40–47, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3583112.

Kaczorowski:2023:OSC

[981] Maya Kaczorowski, Falcon Momot, George Neville-Neil, and Chris Mc-Cubbin. OSS supply-chain security: What will it take? Communications of the ACM, 66(4):48–54, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3583119.

Nassi:2023:CAP

[982] Ben Nassi, Yisroel Mirsky, Jacob Shams, Raz Ben-Netanel, Dudi Nassi, and Yuval Elovici. Contributed articles: Protecting autonomous cars from phantom attacks. Communications of the ACM, 66(4):56–69, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3552308.

Bernal:2023:CAE

[983] Sergio López Bernal, Alberto Huertas Celdrán, and Gregorio Martínez Pérez. Contributed articles: Eight reasons to prioritize brain–computer interface cybersecurity. Communications of the ACM, 66(4):68–78, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3535509.

Bentkamp:2023:RAM

[984] Alexander Bentkamp, Jasmin Blanchette, Visa Nummelin, Sophie Tourret, Petar Vukmirović, and Uwe Waldmann. Review articles: Mechanical mathematicians. Communications of the ACM, 66(4):80–90, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3557998.

Mottola:2023:RHT

[985] Luca Mottola. Research highlights: Technical perspective: On abstractions and embedded networks. Communications of the ACM, 66(4):92, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3583753.

Oostvogels:2023:RHS

[986] Jonathan Oostvogels, Fan Yang, Sam Michiels, and Danny Hughes. Research highlights: Symbol-synchronous buses: Deterministic, low-latency wireless mesh networking with LEDs. Communications of the ACM, 66(4):93–101, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3583762.

Shasha:2023:LBC

[987] Dennis Shasha. Last byte: Cooperation in the commons. Communications of the ACM, 66(4):104–ff, April 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3584697.

Vardi:2023:DAP

[988] Moshe Y. Vardi. Departments: ACM for the public good. Communications of the ACM, 66(5):5, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3589609.

Maitland:2023:CPC

[989] Jules Maitland. Career paths in computing: A career built on using technology to help others. Communications of the ACM, 66(5):7, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3587170.

Tray:2023:SDB

[990] Alex Tray. Securing data for business telephones. Communications of the ACM, 66(5):8–9, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3586581.

Edwards:2023:NAS

[991] Chris Edwards. News: AlphaFold spreads through protein science. Communications of the ACM, 66(5):10–12, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3586582.

Savage:2023:NLS

[992] Neil Savage. News: Locking down secure open source software. Communications of the ACM, 66(5):13–14, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3586584.

Shein:2023:NWC

[993] Esther Shein. News: Women in computer science are making strides. Communications of the ACM, 66(5):15–17, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3586583.

Albert:2023:LTD

[994] Kendra Albert and James Grimmelmann. Law and technology: Do the right thing. Communications of the ACM, 66(5):18–20, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3587825.

Lipner:2023:SUT

[995] Steve Lipner and John Pescatore. Security: Updates, threats, and risk management. Communications of the ACM, 66(5):21–23, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3587826.

Guzdial:2023:EPT

[996] Mark Guzdial, Emma Dodoo, Bahare Naimpour, Tamara Nelson-Fromm, and Aadarsh Padiyath. Education: Putting a teaspoon of programming into other subjects. Communications of the ACM, 66(5):24–26, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3587926.

Steen:2023:VEP

[997] Marc Steen. Viewpoint: Ethics as a participatory and iterative process. Communications of the ACM, 66(5): 27–29, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3550069.

Sevilla:2023:VPR

[998] Jaime Sevilla, Anson Ho, and Tamay Besiroglu. Viewpoint: Please report your compute. Communications of the ACM, 66(5):30–32, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3563035.

Ruppel:2023:VLT

[999] Emily Ruppel, Sihang Liu, Elba Garza, Sukyoung Ryu, Alexandra Silva, and Talia Ringer. Viewpoint: Long-term mentoring for computer science researchers. Communications of the ACM, 66(5):33–35, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3564287.

Amla:2023:VNC

[1000] Nina Amla, Dilma Da Silva, Michael Littman, and Manish Parashar. Viewpoint: NSF on Chien’s grand challenge for sustainability. Communications of the ACM, 66(5):36–37, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3586992.

Ward:2023:RBS

[1001] Peter Ward, Paul Wankadia, and Kavita Guliani. Reinventing backend subsetting at Google. Communications of the ACM, 66(5):40–47, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3588863.

Nagy:2023:RPF

[1002] Stefan Nagy. Research for practice:

The fun in fuzzing. Communications of the ACM, 66(5):48–50, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3588045.

Feitelson:2023:CCM

[1003] Dror G. Feitelson. From code complexity metrics to program comprehension. Communications of the ACM, 66(5):52–61, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3546576.

Broy:2023:DUC

[1004] Manfred Broy and Bernhard Rumpe.

Development use cases for semanticsdriven modeling languages. Communications of the ACM, 66(5):62–71, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3569927.

Gorantala:2023:UPF

[1005] Shruthi Gorantala, Rob Springer, and Bryant Gipson. Unlocking the potential of fully homomorphic encryption. Communications of the ACM, 66(5):72–81, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3572832.

Hoefler:2023:DHP

[1006] Torsten Hoefler, Thomas Häner, and Matthias Troyer. Disentangling hype from practicality: On realistically achieving quantum advantage. Communications of the ACM, 66(5):82–87, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3571725.

Segev:2023:RHT

[1007] Gil Segev. Technical perspective: Finding connections between one-way functions and Kolmogorov complexity. Communications of the ACM, 66(5):90, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3587166.

Liu:2023:TBC

[1008] Yanyi Liu and Rafael Pass. Toward basing cryptography on the hardness of EXP. Communications of the ACM, 66(5):91–99, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3587167.

Raiola:2023:LBC

[1009] Ralph Raiola. Last byte: ChatGPT, can you tell me a story? An exercise in challenging the true creativity of generative AI. Communications of the ACM, 66(5):104–ff, May 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3587998.

Cerf:2023:CCE

[1010] Vinton G. Cerf. Cerf’s up: Connectivity is everything. Communications of the ACM, 66(6):5, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3595116.

Meil:2023:AE

[1011] Doug Meil. AI in the enterprise. Communications of the ACM, 66(6):6–7, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3591212.

Savage:2023:NMC

[1012] Neil Savage. News: Making connections. Communications of the ACM, 66(6):8–10, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3591216.

Greengard:2023:NBA

[1013] Samuel Greengard. News: Better algorithms through faster math. Communications of the ACM, 66(6):11–13, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3591213.

Kugler:2023:NSS

[1014] Logan Kugler. News: Standards to secure the sensors that power IoT. Communications of the ACM, 66(6):14–16, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3591215.

Kirkpatrick:2023:NUA

[1015] Keith Kirkpatrick. News: Using algorithms to deliver disaster aid. Communications of the ACM, 66(6):17–19, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3591214.

Garfinkel:2023:NMW

[1016] Simson L. Garfinkel and Eugene H. Spafford. News: In memoriam: William A. Wulf. Communications of the ACM, 66(6):20–23, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3594711.

Denning:2023:PICa

[1017] Peter J. Denning. The profession of IT: Can generative AI bots be trusted? Communications of the ACM, 66(6):24–27, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3592981.

Neumann:2023:IRC

[1018] Peter G. Neumann. Inside risks: Computer-related risks and remediation challenges. Communications of the ACM, 66(6):28–30, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3593005.

Neville-Neil:2023:KVP

[1019] George V. Neville-Neil. Kode vicious: The parchment path? Communications of the ACM, 66(6):31–32, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3593006.

Haigh:2023:HRC

[1020] Thomas Haigh. Historical reflections:

Conjoined twins: Artificial intelligence and the invention of computer science. Communications of the ACM, 66(6):33–37, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3593007.

Beck:2023:VBD

[1021] Micah D. Beck and Terry R. Moore.

Viewpoint: Breaking up a digital monopoly. Communications of the ACM, 66(6):38–41, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3564239.

Moss:2023:VTE

[1022] Emanuel Moss. Viewpoint: Trust is not enough: Accuracy, error, randomness, and accountability in an algorithmic society. Communications of the ACM, 66(6):42–44, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3565479.

Franceschet:2023:VDA

[1023] Massimo Franceschet. Viewpoint: Decentraland: The alleged decentralization of blockchain applications. Communications of the ACM, 66(6):45–47, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563942.

Falcao:2023:PP

[1024] Gabriel Falcao and João Dinis Ferreira. To PiM or not to PiM. Communications of the ACM, 66(6):48–55, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3589995.

Bird:2023:TFC

[1025] Christian Bird, Denae Ford, Thomas Zimmermann, Nicole Forsgren, Eirini Kalliamvakou, Travis Lowdermilk, and Idan Gazit. Taking flight with Copilot. Communications of the ACM, 66(6):56–62, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3589996.

Fichte:2023:CAS

[1026] Johannes K. Fichte, Daniel Le Berre, Markus Hecher, and Stefan Szeider. Contributed articles: The silent (r)evolution of SAT. Communications of the ACM, 66(6):64–72, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3560469.

Harrison:2023:CAB

[1027] Philippa Harrison. Contributed articles: Buffer-bloated router? How to prevent it and improve performance. Communications of the ACM, 66(6):73–77, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3564262.

Kim:2023:RAS

[1028] Tae Wan Kim and Alan Strudler. Review articles: Should robots have rights or rites? Communications of the ACM, 66(6):78–85, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3571721.

Vardi:2023:RAI

[1029] Gal Vardi. Review articles: On the implicit bias in deep-learning algorithms. Communications of the ACM, 66(6):86–93, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3571070.

Kemper:2023:RHT

[1030] Alfons Kemper. Research highlights:

Technical perspective: FoundationDB performs balancing act. Communications of the ACM, 66(6):96, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592837.

Zhou:2023:RHF

[1031] Jingyu Zhou, Meng Xu, Alexander Shraer, Bala Namasivayam, Alex Miller, Evan Tschannen, Steve Atherton, Andrew J. Beamon, Rusty Sears, John Leach, Dave Rosenthal, Xin Dong, Will Wilson, Ben Collins, David Scherer, Alec Grieser, Yang Liu, Alvin Moore, Bhaskar Muppana, Xiaoge Su, and Vishesh Yadav. Research highlights: FoundationDB: a distributed key–value store. Communications of the ACM, 66(6):97–105, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592838.

Pointcheval:2023:RHT

[1032] David Pointcheval. Research highlights: Technical perspective: What does provable security mean for cryptographic schemes? Communications of the ACM, 66(6):106, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592836.

DeFeo:2023:RHS

[1033] Luca De Feo, Bertram Poettering, and Alessandro Sorniotti. Research highlights: On the (in)Security of El-Gamal in OpenPGP. Communications of the ACM, 66(6):107–115, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592835.

Hoffmann:2023:LBS

[1034] Leah Hoffmann. Last byte: Still networking. Communications of the ACM, 66(6):120–ff, June 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3591217.

Vardi:2023:DRA

[1035] Moshe Y. Vardi. Departments: Revisiting ACM’s open-conference principle. Communications of the ACM, 66(7):5, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3596801.

Elton:2023:CPC

[1036] Hina Elton. Career paths in computing: From UX to product: an intentional career route. Communications of the ACM, 66(7):7, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3596212.

Staff:2023:LEE

[1037] CACM Staff. Letters to the Editor: The end is not clear. Communications of the ACM, 66(7):9, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3596710.

Sze:2023:BUP

[1038] Vivienne Sze. BLOG@CACM: Unleashing the power of deep learning. Communications of the ACM, 66(7):10–11, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3595956.

Shasha:2023:LBS

[1039] Dennis Shasha. Last byte: Sampling the goods. Communications of the ACM, 66(7):140–ff, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3595917.

Edwards:2023:NAO

[1040] Chris Edwards. News: Accelerating optical communications with AI. Communications of the ACM, 66(7):13–15, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595957.

Savage:2023:NRC

[1041] Neil Savage. News: The rise of the chatbots. Communications of the ACM, 66(7):16–17, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3596206.

Mone:2023:NOD

[1042] Gregory Mone. News: Outsmarting deepfake video. Communications of the ACM, 66(7):18–19, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3595958.

Samuelson:2023:LSLb

[1043] Pamela Samuelson. Legally speaking: Legal challenges to generative AI, Part I. Communications of the ACM, 66(7):20–23, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3597151.

Acquisti:2023:PLL

[1044] Alessandro Acquisti and Ryan Steed. Privacy: Learning to live with privacypreserving analytics. Communications of the ACM, 66(7):24–27, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3597173.

Johnson:2023:VFA

[1045] Jeff A. Johnson and Daniel H. Bullock. Viewpoint: Fragility in AIs using artificial neural networks. Communications of the ACM, 66(7):28–31, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3571280.

Zaman:2023:VLA

[1046] Tariq Zaman. Viewpoint: Lost in Afghanistan: Can the world take ICT4D seriously? Communications of the ACM, 66(7):32–34, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3574156.

Salim:2023:EAO

[1047] Flora Salim, Bingsheng He, and Ken ichi Kawarabayashi. Welcome. Communications of the ACM, 66(7):36–38, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3598580. East Asia and

Oceania Region Special Section.

Shortridge:2023:OCM

[1048] Kelly Shortridge and Josiah Dykstra.

Opportunity cost and missed chances in optimizing cybersecurity. Communications of the ACM, 66(7):96–104, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3597464.

Ozsu:2023:CAD

[1049] M. Tamer Özsu. Contributed articles:

Data science — a systematic treatment. Communications of the ACM, 66(7):106–116, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3582491.

Medvedev:2023:RAT

[1050] Paul Medvedev. Review articles: Theoretical analysis of sequencing bioinformatics algorithms and beyond. Communications of the ACM, 66(7):118– 125, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3571723.

Sitaraman:2023:RHO

[1051] Ramesh K. Sitaraman. Research highlights: Opening the door to SSD algorithmics: Technical perspective. Communications of the ACM, 66(7):128, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3596204.

Lange:2023:RHO

[1052] Tomer Lange, Joseph (Seffi) Naor, and Gala Yadgar. Research highlights: Offline and online algorithms for SSD management. Communications of the ACM, 66(7):129–137, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3596205.

Spina:2023:BTH

[1053] Damiano Spina, Mark Sanderson, Daniel Angus, Gianluca Demartini, Dana Mckay, Lauren L. Saling, and Ryen W. White. Big trends: Human– AI cooperation to tackle misinformation and polarization. Communications of the ACM, 66(7):40–45, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/10.

1145/3588431. East Asia and Oceania Region Special Section.

Kim:2023:BTS

[1054] Ji-Hoon Kim, Sungyeob Yoo, and Joo-Young Kim. Big trends: South Korea’s nationwide effort for AI semiconductor industry. Communications of the ACM, 66(7):46–51, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3587264. East Asia and

Oceania Region Special Section.

Luo:2023:BTA

[1055] Tao Luo, Weng-Fai Wong, Rick Siow Mong Goh, Anh Tuan Do, Zhixian Chen, Haizhou Li, Wenyu Jiang, and Weiyun Yau. Big trends: Achieving green AI with energy-efficient deep learning using neuromorphic computing. Communications of the ACM, 66(7):52–57, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3588591.

East Asia and Oceania Region Special Section.

Kitsuregawa:2023:BTA

[1056] Masaru Kitsuregawa, Shigeo Urushidani, Kazutsuna Yamaji, Hiroki Takakura, Ichiro Hasuo, Imari Sato, Fuyuki Ishikawa, Isao Echizen, and Kensaku Mori. Big trends: Activities of national institute of informatics in Japan. Communications of the ACM, 66(7):58–63, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3589736.

East Asia and Oceania Region Special Section.

Lu:2023:HTO

[1057] Qinghua Lu, Liming Zhu, Xiwei Xu, Jon Whittle, Didar Zowghi, and Aurelie Jacquet. Hot topics: Operationalizing responsible AI at scale: CSIRO Data61’s pattern-oriented responsible AI engineering approach. Communications of the ACM, 66(7):64–66, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3589946. East Asia and

Oceania Region Special Section.

Luczak-Roesch:2023:HTV

[1058] Markus Luczak-Roesch, Matthias Galster, and Kevin Shedlock. Hot topics: The veracity grand challenge in computing: a perspective from Aotearoa New Zealand. Communications of the ACM, 66(7):67–69, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3589154. East Asia and

Oceania Region Special Section.

Bhowmick:2023:HTC

[1059] Sourav S. Bhowmick. Hot topics: CLOSET: Data-driven COI detection and management in peer-review venues. Communications of the ACM, 66(7):70–71, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3589151.

East Asia and Oceania Region Special Section.

Ogata:2023:HTL

[1060] Hiroaki Ogata, Rwitajit Majumdar, and Brendan Flanagan. Hot topics: Learning and evidence analytics framework bridges research and practice for educational data science. Communications of the ACM, 66(7):72–74, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3589660. East Asia and

Oceania Region Special Section.

Bui:2023:HTB

[1061] Hung Bui, Minh Hoai Nguyen, Dat Quoc Nguyen, Linh Pham, and Dinh Phung. Hot topics: Building and nurturing AI development in Vietnam. Communications of the ACM, 66(7): 75–76, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3589948. East Asia

and Oceania Region Special Section.

Zhang:2023:HTP

[1062] Hao Zhang, Jun Yang, Cheng Chen, Siqi Wang, Jiashu Li, and Mian Lu. Hot topics: Principles and practices of real-time feature computing platforms for ML. Communications of the ACM, 66(7):77–78, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3589224. East Asia and

Oceania Region Special Section.

Gramoli:2023:HTF

[1063] Vincent Gramoli and Qiang Tang.

Hot topics: The future of blockchain consensus. Communications of the ACM, 66(7):79–80, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3589225.

East Asia and Oceania Region Special Section.

Putra:2023:HTC

[1064] Guntur Dharma Putra, Sidra Malik, Volkan Dedeoglu, Raja Jurda, and Salil S. Kanhere. Hot topics: Challenges in designing blockchain for cyber-physical systems. Communications of the ACM, 66(7):81–82, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3589648. East Asia and

Oceania Region Special Section.

Bhowmick:2023:HTD

[1065] Sourav S. Bhowmick and Byron Choi. Hot topics: To draw is human: Toward no-code subgraph search. Communications of the ACM, 66(7):83–84, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3589153. East Asia and

Oceania Region Special Section.

Ikoma:2023:HTD

[1066] Eiji Ikoma and Masaru Kitsuregawa. Hot topics: DIAS — Earth environment Data Integration and Analysis System. Communications of the ACM, 66(7):85–86, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3589233.

East Asia and Oceania Region Special Section.

Georgakopoulos:2023:HTD

[1067] Dimitrios Georgakopoulos and Prem Prakash Jayaraman. Hot topics: Digital twins and dependency/constraint-aware AI for digital manufacturing. Communications of the ACM, 66(7):87–88, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/10.

1145/3589662. East Asia and Oceania Region Special Section.

Withana:2023:HTC

[1068] Anusha Withana. Hot topics: Codesigning personalized assistive devices using personal fabrication. Communications of the ACM, 66(7):89–90, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3589150. East Asia and

Oceania Region Special Section.

Inami:2023:HTJ

[1069] Masahiko Inami. Hot topics: JIZAI body: Human-machine integration based on Asian thought. Communications of the ACM, 66(7):91, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592539. East Asia and

Oceania Region Special Section.

Kawahara:2023:HTF

[1070] Yoshihiro Kawahara. Hot topics: Fusing creativity and innovation in Asia’s manufacturing industry. Communications of the ACM, 66(7):92–93, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3589949. East Asia and

Oceania Region Special Section.

Higashino:2023:HTM

[1071] Teruo Higashino, Hidetoshi Kotera, and Yasushi Yagi. Hot topics: MEXT Society 5.0 realization research support project. Communications of the ACM, 66(7):94–95, July 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3589232. East Asia and

Oceania Region Special Section.

Chien:2023:ELG

[1072] Andrew A. Chien. Editor’s letter: GenAI: Giga$, TeraWatt-hours, and GigaTons of CO_{2}. Communications of the ACM, 66(8):5, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3606254.

Cerf:2023:CLL

[1073] Vinton G. Cerf. Cerf’s up: Large language models. Communications of the ACM, 66(8):7, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3606337.

Hong:2023:TFC

[1074] Jason I. Hong. Teaching the FATE community about privacy. Communications of the ACM, 66(8):10–11, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3603718.

Monroe:2023:NCA

[1075] Don Monroe. News: Computerassisted proofs take on fluid flow. Communications of the ACM, 66(8):12–14, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3603745.

Greengard:2023:NMD

[1076] Samuel Greengard. News: 3D modeling draws on AI. Communications of the ACM, 66(8):15–16, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3603748.

Kirkpatrick:2023:NCF

[1077] Keith Kirkpatrick. News: The carbon footprint of artificial intelligence. Communications of the ACM, 66(8):17–19, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3603746.

Williams:2023:CEV

[1078] Tom Williams, Cynthia Matuszek, Kristiina Jokinen, Raj Korpan, James Pustejovsky, and Brian Scassellati. Computing ethics: Voice in the machine: Ethical considerations for language-capable robots. Communications of the ACM, 66(8):20–23, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3604632.

Shapiro:2023:EIC

[1079] R. Benjamin Shapiro, Kayla DesPortes, and Betsy DiSalvo. Education: Improving computing education research through valuing design. Communications of the ACM, 66(8):24–26, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604633.

Neville-Neil:2023:KVHa

[1080] George Neville-Neil. Kode vicious:

The human touch. Communications of the ACM, 66(8):27–28, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604634.

Temiz:2023:VSB

[1081] Hakan Temiz. Viewpoint: SI and binary prefixes: Clearing the confusion. Communications of the ACM, 66(8):29–31, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3572027.

Izquierdo:2023:VMT

[1082] Javier Luis Cánovas Izquierdo and Jordi Cabot. Viewpoint: For a more transparent governance of open source. Communications of the ACM, 66(8):32–34, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3570635.

Sheh:2023:VWY

[1083] Wesley Sheh and Joel Nishimura. Viewpoint: Why you should be able to make your own individualized, digital nano-currency. Communications of the ACM, 66(8):35–38, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3583987.

Kuipers:2023:VAS

[1084] Benjamin Kuipers. Viewpoint: AI and society: Ethics, trust, and cooperation. Communications of the ACM, 66(8):39–42, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3583134.

Garfinkel:2023:SYT

[1085] Simson Garfinkel and Jon Stewart. Sharpening your tools. Communications of the ACM, 66(8):44–52, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3600098.

Norvig:2023:MTJ

[1086] Peter Norvig, Alfred Spector, Chris Wiggins, Jeannette M. Wing, Ben Fried, and Michael Tingley. More than just algorithms. Communications of the ACM, 66(8):53–61, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3600099.

Hope:2023:CIS

[1087] Tom Hope, Doug Downey, Daniel S.

Weld, Oren Etzioni, and Eric Horvitz. A computational inflection for scientific discovery. Communications of the ACM, 66(8):62–73, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3576896.

Oosthoek:2023:TTM

[1088] Kris Oosthoek, Jack Cable, and Georgios Smaragdakis. A tale of two markets: Investigating the ransomware payments economy. Communications of the ACM, 66(8):74–83, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3582489.

Jarrahi:2023:PDC

[1089] Mohammad Hossein Jarrahi, Ali Memariani, and Shion Guha. The principles of data-centric AI. Communications of the ACM, 66(8):84–92, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3571724.

Alouf:2023:CWU

[1090] Sara Alouf. Can we uncover private backbone infrastructures? Communications of the ACM, 66(8):94, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604621.

Salamatian:2023:MVC

[1091] Loqman Salamatian, Scott Anderson, Joshua Mathews, Paul Barford, Walter Willinger, and Mark Crovella. A manifold view of connectivity in the private backbone networks of hyperscalers. Communications of the ACM, 66(8):95–103, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3604620.

Richards:2023:PRW

[1092] John Richards. Pandemics, remote work, and accessibility. Communications of the ACM, 66(8):104, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604623.

Mack:2023:MAV

[1093] Kelly Mack, Maitraye Das, Dhruv Jain, Danielle Bragg, John Tang, Andrew Begel, Erin Beneteau, Josh Urban Davis, Abraham Glasser, Joon Sung Park, and Venkatesh Potluri. Mixed abilities and varied experiences: a group autoethnography of a virtual summer internship. Communications of the ACM, 66(8):105–113, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604622.

Kaye:2023:LBS

[1094] Jofish Kaye. Last byte: L-space and large language models. Communications of the ACM, 66(8):116–ff, August 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3596900.

Vardi:2023:DRT

[1095] Moshe Y. Vardi. Departments: To regulate tech, nullify click-through contracts. Communications of the ACM, 66(9):5, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3609981.

Rivera:2023:CPC

[1096] Fiona Rivera. Career paths in computing: Finding the right mix of creativity and technology. Communications of the ACM, 66(9):7, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3607823.

Baquero:2023:BYE

[1097] Carlos Baquero. BLOG@CACM: A year embedded in the Crypto-NFT space. Communications of the ACM, 66(9):8–9, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3607876.

Edwards:2023:NHA

[1098] Chris Edwards. News: Historic algorithms help unlock shortest-path problem breakthrough. Communications of the ACM, 66(9):10–12, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3607866.

Shein:2023:NSL

[1099] Esther Shein. News: Shining a light on the Dark Web. Communications of the ACM, 66(9):13–14, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3607867.

Greengard:2023:NDW

[1100] Samuel Greengard. News: Distilling what we know. Communications of the ACM, 66(9):15–17, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3607891.

Narechania:2023:LTC

[1101] Tejas N. Narechania. Law and technology: Closing our digital divides. Communications of the ACM, 66(9): 18–20, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3608964.

Zanero:2023:SCA

[1102] Stefano Zanero. Security: Coming of age. Communications of the ACM, 66 (9):21–23, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3608965.

Denning:2023:PIS

[1103] Peter J. Denning. The profession of IT:

The smallness of large language models. Communications of the ACM, 66 (9):24–27, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3608966.

Wu:2023:ODA

[1104] Weiyue Wu and Shaoshan Liu. Opinion: Dilemma of the artificial intelligence regulatory landscape. Communications of the ACM, 66(9):28–31, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3584665.

Ligett:2023:OWN

[1105] Katrina Ligett and Kobbi Nissim.

Opinion: We need to focus on how our data is used, not just how it is shared. Communications of the ACM, 66(9):32–34, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3572912.

Dershowitz:2023:ORR

[1106] Nachum Dershowitz and Rakesh M.

Verma. Opinion: Rebutting rebuttals. Communications of the ACM, 66 (9):35–41, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3584664.

Yasseri:2023:OCC

[1107] Taha Yasseri and Filippo Menczer. Opinion: Can crowdsourcing rescue the social marketplace of ideas? Communications of the ACM, 66(9):42–45, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3578645.

Levine:2023:CCA

[1108] Edlyn V. Levine. Cargo cult AI. Communications of the ACM, 66(9): 46–51, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606946.

Kaffes:2023:RPS

[1109] Kostis Kaffes and Peter Alvaro. Research for practice: OS scheduling. Communications of the ACM, 66(9): 52–54, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606945.

Zhao:2023:RHC

[1110] Shengdong Zhao, Felicia Tan, and Katherine Fennedy. Research: Headsup computing moving beyond the device-centered paradigm. Communications of the ACM, 66(9):56–63, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3571722.

Bakirtzis:2023:RDC

[1111] Georgios Bakirtzis, Steven Carr, David Danks, and Ufuk Topcu. Research: Dynamic certification for autonomous systems. Communications of the ACM, 66 (9):64–72, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3574133.

Knowles:2023:RHA

[1112] Bran Knowles, Jason D’Cruz, John T.

Richards, and Kush R. Varshney. Research: Humble AI. Communications of the ACM, 66(9):73–79, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3587035.

Kounev:2023:RSC

[1113] Samuel Kounev, Nikolas Herbst, Cristina L. Abad, Alexandru Iosup, Ian Foster, Prashant Shenoy, Omer Rana, and Andrew A. Chien. Research: Serverless computing: What it is, and what it is not? Communications of the ACM, 66(9):80–92, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3587249.

Woodruff:2023:RHT

[1114] David P. Woodruff. Research highlights: Tapping the link between algorithmic model counting and streaming: Technical perspective. Communications of the ACM, 66(9):94, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3607825.

Pavan:2023:RHM

[1115] A. Pavan, N. V. Vinodchandran, Arnab Bhattacharyya, and Kuldeep S. Meel. Research highlights: Model counting meets distinct elements. Communications of the ACM, 66(9):95–102, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3607824.

Hoffmann:2023:LBH

[1116] Leah Hoffmann. Last byte: How many ways can you teach a robot? Communications of the ACM, 66(9): 104–ff, September 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3608109.

Cerf:2023:CDS

[1117] Vinton G. Cerf. Cerf’s up: The dilemma of scale. Communications of the ACM, 66(10):5, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3617385.

Marcus:2023:BCA

[1118] Gary Marcus. BLOG@CACM: Controlling AI. Communications of the ACM, 66(10):6–7, October 2023. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3613250.

Mone:2023:NBD

[1119] Gregory Mone. News: Biomedical digital twins. Communications of the ACM, 66(10):9–11, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3613251.

Kugler:2023:NFR

[1120] Logan Kugler. News: The fight to repair. Communications of the ACM, 66(10):12–14, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3613252.

Geer:2023:NUP

[1121] David Geer. News: Using psychology to bolster cybersecurity. Communications of the ACM, 66(10):15–17, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3613253.

Cusumano:2023:TSMb

[1122] Michael A. Cusumano. Technology strategy and management: Generative AI as a new innovation platform. Communications of the ACM, 66(10):18–21, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3615859.

Leveson:2023:IRC

[1123] Nancy G. Leveson and John P.

Thomas. Inside risks: Certification of safety-critical systems. Communications of the ACM, 66(10):22–26, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3615860.

Neville-Neil:2023:KVS

[1124] George V. Neville-Neil. Kode vicious:

Stone knives and bear skins. Communications of the ACM, 66(10):27–29, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3615861.

Kuyenga:2023:OBC

[1125] Madison C. Allen Kuyenga, Eleanor R. Glover Gladney, Michael Lachney, Marwin McKnight, Theodore S. Ransaw, Dominick Sanders, and Aman Yadav. Opinion: Barbershop computing. Communications of the ACM, 66 (10):30–33, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3586990.

Bessen:2023:OHS

[1126] James Bessen. Opinion: How software stifles competition and innovation. Communications of the ACM, 66(10):34–36, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3588898.

Bar:2023:ONT

[1127] Dominik Bär, Nicolas Pröllochs, and Stefan Feuerriegel. Opinion: New threats to society from free-speech social media platforms. Communications of the ACM, 66(10):37–40, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3587094.

Skarbez:2023:OIT

[1128] Rick Skarbez, Missie Smith, and Mary Whitton. Opinion: It is time to let go of ‘virtual reality’. Communications of the ACM, 66(10):41–43, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3590959.

Tellman:2023:DFC

[1129] Zachary Tellman. Designing a framework for conversational interfaces. Communications of the ACM, 66 (10):44–49, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3605572.

Casari:2023:BR

[1130] Amanda Casari, Julia Ferraioli, and Juniper Lovato. Beyond the repository. Communications of the ACM, 66(10):50–55, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3605160.

Seymour:2023:RBD

[1131] Michael Seymour, Kai Riemer, Lingyao Yuan, and Alan R. Dennis. Research: Beyond deep fakes. Communications of the ACM, 66(10):56–67, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3584973.

Hearst:2023:RSI

[1132] Marti A. Hearst. Research: Show it or tell it? Text, visualization, and their combination. Communications of the ACM, 66(10):68–75, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3593580.

Hirzel:2023:RLC

[1133] Martin Hirzel. Research: Low-code programming models. Communications of the ACM, 66(10):76–85, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3587691.

Barbosa:2023:RGE

[1134] Haniel Barbosa, Clark Barrett, Byron Cook, Bruno Dutertre, Gereon Kremer, Hanna Lachnitt, Aina Niemetz, Andres Nötzli, Alex Ozdemir, Mathias Preiner, Andrew Reynolds, Cesare Tinelli, and Yoni Zohar. Research: Generating and exploiting automated reasoning proof certificates. Communications of the ACM, 66(10):86–95, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3587692.

Sekar:2023:RHR

[1135] Shreyas Sekar. Research highlights: A rare glimpse of tracking fake reviews: Technical perspective. Communications of the ACM, 66(10):97, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3615429.

He:2023:RHL

[1136] Sherry He, Brett Hollenbeck, and Davide Proserpio. Research highlights: Leveraging social media to buy fake reviews. Communications of the ACM, 66(10):98–105, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3615428.

Huang:2023:RHS

[1137] Polly Huang. Research highlights: Sensing interaction with everyday objects using near-field communication in textiles: Technical perspective. Communications of the ACM, 66(10):106, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3615451.

Wang:2023:RHL

[1138] Jingxian Wang, Junbo Zhang, Ke Li, Chengfeng Pan, Carmel Majidi, and Swarun Kumar. Research highlights: Locating everyday objects using NFC textiles. Communications of the ACM, 66(10):107–114, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3615450.

Shasha:2023:LBT

[1139] Dennis Shasha. Last byte: Tidy towers. Communications of the ACM, 66 (10):116–ff, October 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3614559.

Vardi:2023:DWC

[1140] Moshe Y. Vardi. Departments: What came first, math or computing? Communications of the ACM, 66(11):5, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3625734.

Meil:2023:BLP

[1141] Doug Meil. BLOG@CACM: Lessons from PL/I: a most ambitious programming language. Communications of the ACM, 66(11):6–7, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3623564.

Monroe:2023:NQS

[1142] Don Monroe. News: Quantum speedup for the Fast Fourier Transform? Communications of the ACM, 66(11):8–10, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3623641.

Kugler:2023:NWO

[1143] Logan Kugler. News: What’s old is new again. Communications of the ACM, 66 (11):11–12, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3624009.

Kirkpatrick:2023:NAI

[1144] Keith Kirkpatrick. News: Accessibility and inclusion through technology. Communications of the ACM, 66 (11):13–15, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3623642.

Samuelson:2023:LSL

[1145] Pamela Samuelson. Legally speaking: Legal challenges to generative AI, Part II. Communications of the ACM, 66 (11):16–19, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3625251.

Jacobson:2023:OBC

[1146] Sheldon H. Jacobson. Opinion: On being a computer science communicator. Communications of the ACM, 66 (11):20–22, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3588899.

Flaherty:2023:EBD

[1147] Sean Flaherty and Gregg Gordon. Economic and business dimensions: The gamification of academia. Communications of the ACM, 66(11):23–24, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3625253.

Warden:2023:OML

[1148] Pete Warden, Matthew Stewart, Brian Plancher, Sachin Katti, and Vijay Janapa Reddi. Opinion: Machine learning sensors. Communications of the ACM, 66(11):25–28, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3586991.

Das:2023:PPA

[1149] Sauvik Das, Hao-Ping (Hank) Lee, and Jodi Forlizzi. Privacy: Privacy in the age of AI. Communications of the ACM, 66(11):29–31, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3625254.

Mendonca:2023:OIA

[1150] Ricardo Fabrino Mendonça, Fernando Filgueiras, and Virgilio Almeida. Opinion: The infrapolitics of algorithmic resistance. Communications of the ACM, 66(11):32–34, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3573008.

Forster:2023:ORC

[1151] Anna Förster. Opinion: Rethinking conference formats. Communications of the ACM, 66(11):35–37, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3583679.

Videla:2023:EI

[1152] Alvaro Videla. Echoes of intelligence. Communications of the ACM, 66(11):38–43, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3610283.

Noda:2023:DWA

[1153] Abi Noda, Margaret-Anne Storey, Nicole Forsgren, and Michaela Greiler. DevEX: What actually drives productivity? Communications of the ACM, 66(11):44–49, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3610285.

Cazzola:2023:RSF

[1154] Walter Cazzola and Luca Favalli. Research: Scrambled features for breakfast: Concepts of agile language development. Communications of the ACM, 66(11):50–60, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3596217.

DiNitto:2023:RIH

[1155] Elisabetta Di Nitto, Ismael García-Varea, Mehdi Jazayeri, Damian A.

Tamburri, and Svetlana Tikhonenko. Research: Informatics higher education in Europe: a data portal and case study. Communications of the ACM, 66 (11):61–67, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3591107.

Prakash:2023:RTS

[1156] Shvetank Prakash, Matthew Stewart, Colby Banbury, Mark Mazumder, Pete Warden, Brian Plancher, and Vijay Janapa Reddi. Research: Is TinyML sustainable? Communications of the ACM, 66(11):68–77, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3608473.

Sattath:2023:RUC

[1157] Or Sattath. Research: Uncloneable cryptography. Communications of the ACM, 66(11):78–86, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3576897.

Fraser:2023:RHT

[1158] Gordon Fraser. Research highlights:

Technical perspective: What’s all the fuss about fuzzing? Communications of the ACM, 66(11):88, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3617381.

Bohme:2023:RHB

[1159] Marcel Böhme, Valentin J. M. Manès, and Sang Kil Cha. Research highlights:

Boosting fuzzer efficiency: an information theoretic perspective. Communications of the ACM, 66(11):89–97, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3611019.

Koningstein:2023:LBH

[1160] Ross Koningstein. Last byte: How climate change was won. Communications of the ACM, 66(11):100–ff, November 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3623643.

Cerf:2023:CET

[1161] Vinton G. Cerf. Cerf’s up: On expert testimony. Communications of the ACM, 66(12):5, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3631277.

Chayka:2023:BCC

[1162] Antony Chayka and Andrei Sukhov. BLOG@CACM: Comparing chatbots trained in different languages. Communications of the ACM, 66(12):6–7, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3626835.

Edwards:2023:NTT

[1163] Chris Edwards. News: Tales of topological qubits. Communications of the ACM, 66(12):8–10, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624436.

Shein:2023:NSD

[1164] Esther Shein. News: Saving digital libraries and the Internet Archive. Communications of the ACM, 66(12): 11–12, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3625392.

Savage:2023:NRP

[1165] Neil Savage. News: Revamping Python for an AI world. Communications of the ACM, 66(12):13–14, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624987.

Dutta:2023:OTS

[1166] Subhabrata Dutta and Tanmoy Chakraborty.

Opinion: Thus spake ChatGPT. Communications of the ACM, 66(12):16–19, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3616863.

Malizia:2023:OWC

[1167] Alessio Malizia and Fabio Paternò.

Opinion: Why is the current XAI not meeting the expectations? Communications of the ACM, 66(12):20–23, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3588313.

Denning:2023:PICb

[1168] Peter J. Denning and Todd W. Lyons.

The profession of IT: A clash of civilizations. Communications of the ACM, 66 (12):24–26, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3625832.

Neville-Neil:2023:KVHb

[1169] George V. Neville-Neil. Kode vicious: Halfway around the world. Communications of the ACM, 66(12):27–28, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3625551.

Berndtsson:2023:OSS

[1170] Mikael Berndtsson, Anna-Carin Jonsson, Magnus Carlsson, and Thomas Svahn. Opinion: A strategy for scaling advanced analytics. Communications of the ACM, 66(12):29–31, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3582075.

Andriole:2023:OFW

[1171] Stephen Andriole. Opinion: Five ways executives misunderstand technology. Communications of the ACM, 66(12):32–34, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3595291.

Haigh:2023:HRT

[1172] Thomas Haigh. Historical reflections: There was no ‘First AI Winter’. Communications of the ACM, 66 (12):35–39, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3625833.

Miller:2023:PFM

[1173] Ethan L. Miller, George Neville-Neil, Achilles Benetopoulos, Pankaj Mehra, and Daniel Bittman. Pointers in far memory. Communications of the ACM, 66(12):40–45, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3617581.

Stark:2023:HFC

[1174] Samuel W. Stark, A. Theodore Markettos, and Simon W. Moore. How flexible is CXL’s memory protection? Communications of the ACM, 66(12): 46–51, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3617580.

Elmqvist:2023:RDA

[1175] Niklas Elmqvist. Research: Data analytics anywhere and everywhere. Communications of the ACM, 66(12): 52–63, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3584858.

Medvedev:2023:RTA

[1176] Paul Medvedev. Research: Theoretical analysis of edit distance algorithms. Communications of the ACM, 66(12):64–71, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3582490.

Lee:2023:RWT

[1177] Jongho Lee, Mohit Gupta, Bhuvana Krishnaswamy, and Suman Banerjee. Research: When two cameras are a crowd. Communications of the ACM, 66(12):72–82, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3588998.

Teng:2023:RHT

[1178] Shang-Hua Teng. Research highlights: Technical perspective: Maximum flow through a network: a storied problem and a groundbreaking solution. Communications of the ACM, 66(12):84, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3623277.

Chen:2023:RHA

[1179] Li Chen, Rasmus Kyng, Yang P. Liu, Richard Peng, Maximilian Probst Gutenberg, and Sushant Sachdeva. Research highlights: Almost-linear-time algorithms for maximum flow and minimum-cost flow. Communications of the ACM, 66(12):85–92, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3610940.

Hoffmann:2023:LBU

[1180] Leah Hoffmann. Last byte: Ulterior motives. Communications of the ACM, 66(12):96–ff, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3625241.

Vardi:2024:DCY

[1181] Moshe Y. Vardi. Departments: Computing, you have blood on your hands! Communications of the ACM, 67(1): 5, January 2024. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3632963.

Zheng:2024:CPC

[1182] Sam Zheng. Career paths in computing: From eye tracking to AI-powered learning. Communications of the ACM, 67(1):7, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3632182.

Vakulov:2024:BPL

[1183] Alex Vakulov. BLOG@CACM: Protecting life-saving medical devices from cyberattack. Communications of the ACM, 67(1):8–9, January 2024. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

https://dl.acm.org/doi/10.1145/
3631535.

Monroe:2024:NEN

[1184] Don Monroe. News: Epigenomics now. Communications of the ACM, 67(1):11–13, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3631934.

Goth:2024:NWL

[1185] Gregory Goth. News: Why are lawyers afraid of AI? Communications of the ACM, 67(1):14–16, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3631935.

Greengard:2024:NWG

[1186] Samuel Greengard. News: Wayfinding without GPS. Communications of the ACM, 67(1):17–19, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3627104.

Hall:2024:APF

[1187] Wendy Hall, Divesh Srivastava, and Scott Delman. ACM publications finances: ACM publications finances for 2022: Striving for balance between mission, sustainability, and growth. Communications of the ACM, 67(1):22–26, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3635166.

Demartini:2024:ODB

[1188] Gianluca Demartini, Kevin Roitero, and Stefano Mizzaro. Opinion: Data bias management. Communications of the ACM, 67(1):28–32, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3611641.

Cusumano:2024:TSM

[1189] Michael Cusumano. Technology strategy and management: NVIDIA at the center of the generative AI ecosystem — for now. Communications of the ACM, 67(1):33–35, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3631537.

Langvardt:2024:LTB

[1190] Kyle Langvardt and Alan Z. Rozenshtein. Law and technology: Beyond the editorial analogy: The future of the First Amendment on the Internet. Communications of the ACM, 67(1):36–38, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3631538.

Suhail:2024:SPL

[1191] Sabah Suhail, Mubashar Iqbal, and Raja Jurdak. Security: The perils of leveraging evil digital twins as securityenhancing enablers. Communications of the ACM, 67(1):39–42, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3631539.

Stefanija:2024:OTS

[1192] Ana Pop Stefanija, Bart Buelens, Elfi Goesaert, Tom Lenaerts, Jo Pierson, and Jan Van den Bussche. Opinion: Toward a solid acceptance of the decentralized Web of Personal Data: Societal and technological convergence. Communications of the ACM, 67(1):43–46, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3624555.

Yao:2024:ORH

[1193] Danfeng (Daphne) Yao. Opinion: Rebuttal how-to: Strategies, tactics, and the big picture in research. Communications of the ACM, 67(1):47–49, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624722.

Hoffmann:2024:QAM

[1194] Leah Hoffmann. Q&A: Achievement in microarchitecture. Communications of the ACM, 67(1):50–51, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3627709.

Russinovich:2024:CCE

[1195] Mark Russinovich. Confidential computing: Elevating cloud security and privacy. Communications of the ACM, 67(1):52–53, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3624577.

Kaplan:2024:HVI

[1196] David Kaplan. Hardware VM isolation in the cloud. Communications of the ACM, 67(1):54–59, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3624576.

Dhanuskodi:2024:CFC

[1197] Gobikrishna Dhanuskodi, Sudeshna Guha, Vidhya Krishnan, Aruna Manjunatha, Rob Nertney, Michael O’Connor, and Phil Rogers. Creating the first confidential GPUs. Communications of the ACM, 67(1):60–67, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626827.

Delignat-Lavaud:2024:WST

[1198] Antoine Delignat-Lavaud, Cédric Fournet, Kapil Vaswani, Sylvan Clebsch, Maik Riechert, Manuel Costa, and Mark Russinovich. Why should I trust your code? Communications of the ACM, 67(1):68–76, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624578.

Brown:2024:RTS

[1199] Neil C. C. Brown, Felienne F. J. Hermans, and Lauren E. Margulieux. Research: 10 things software developers should learn about learning. Communications of the ACM, 67(1):78–87, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3584859.

Raghavan:2024:RWS

[1200] Manish Raghavan. Research: What should we do when our ideas of fairness conflict? Communications of the ACM, 67(1):88–97, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3587930.

Abeywickrama:2024:RST

[1201] Dhaminda B. Abeywickrama, Amel Bennaceur, Greg Chance, Yiannis Demiris, Anastasia Kordoni, Mark Levine, Luke Moffat, Luc Moreau, Mohammad Reza Mousavi, Bashar Nuseibeh, Subramanian Ramamoorthy, Jan Oliver Ringert, James Wilson, Shane Windsor, and Kerstin Eder. Research: On specifying for trustworthiness. Communications of the ACM, 67(1):98–109, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624699.

Du:2024:RSL

[1202] Mengnan Du, Fengxiang He, Na Zou, Dacheng Tao, and Xia Hu. Research: Shortcut learning of large language models in natural language understanding. Communications of the ACM, 67(1):110–120, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3596490.

Waller:2024:RHU

[1203] Annalu Waller. Research highlights: The user as a key ingredient in AAC design: Technical perspective. Communications of the ACM, 67(1):122, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3629475.

Valencia:2024:RHN

[1204] Stephanie Valencia, Mark Steidl, Michael L. Rivera, Cynthia L. Bennett, Jeffrey P. Bigham, and Henny Admoni. Research highlights: Nonverbal communication through expressive objects. Communications of the ACM, 67 (1):123–131, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/3610939.

Shasha:2024:LBS

[1205] Dennis Shasha. Last byte: Seesaw gold. Communications of the ACM, 67(1):136–ff, January 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3630113.

Cerf:2024:CVF

[1206] Vinton G. Cerf. Cerf’s up: Validating factual personal information. Communications of the ACM, 67(2):5, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638528.

Tray:2024:BWB

[1207] Alex Tray. BLOG@CACM: Why bother localizing information technology products? Communications of the ACM, 67(2):6–7, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638537.

Klarreich:2024:NAA

[1208] Erica Klarreich. News: Algorithmic advance: The group isomorphism problem. Communications of the ACM, 67(2):9–11, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3637207.

Edwards:2024:NTT

[1209] Chris Edwards. News: Teaching transformed. Communications of the ACM, 67(2):12–13, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3637208.

Mone:2024:NVR

[1210] Gregory Mone. News: Virtual reality as therapy. Communications of the ACM, 67(2):14–15, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638536.

Hoadley:2024:OAP

[1211] Christopher Hoadley and Sara Vogel.

Opinion: Autocorrect is not: People are multilingual and computer science should be too. Communications of the ACM, 67(2):16–18, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3633585.

Diakopoulos:2024:CEL

[1212] N. Diakopoulos, C. Trattner, D. Jannach, I. Costera Meijer, and E. Motta. Computing ethics: Leveraging professional ethics for responsible AI. Communications of the ACM, 67(2):19–21, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3625252.

Haigh:2024:HRH

[1213] Thomas Haigh. Historical reflections: How the AI boom went bust. Communications of the ACM, 67(2):22–26, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3634901.

Neville-Neil:2024:KVD

[1214] George V. Neville-Neil. Kode vicious: Dear diary. Communications of the ACM, 67(2):27–28, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3633288.

Simha:2024:OUC

[1215] Rahul Simha, Amruth N. Kumar, and Rajendra K. Raj. Opinion: Undergraduate computer science curricula. Communications of the ACM, 67 (2):29–31, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624729.

Fraser:2024:OVF

[1216] Steven Fraser and Dennis Mancl. Opinion: Virtual and the future of conferences. Communications of the ACM, 67(2):32–34, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624638.

Spector:2024:OGB

[1217] Alfred Z. Spector. Opinion: Gaining benefit from artificial intelligence and data science: a three-part framework. Communications of the ACM, 67(2):35–38, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624726.

Vachon:2024:SM

[1218] Phil Vachon. Security mismatch. Communications of the ACM, 67(2):40–41, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3630746.

Limoncelli:2024:KWY

[1219] Thomas A. Limoncelli. Knowing what you need to know. Communications of the ACM, 67(2):42–46, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3630745.

Buyl:2024:RIL

[1220] Maarten Buyl and Tijl De Bie. Research: Inherent limitations of AI fairness. Communications of the ACM, 67(2):48–55, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624700.

Denny:2024:RCE

[1221] Paul Denny, James Prather, Brett A. Becker, James Finnie-Ansley, Arto Hellas, Juho Leinonen, Andrew Luxton-Reilly, Brent N. Reeves, Eddie Antonio Santos, and Sami Sarsa. Research: Computing education in the era of generative AI. Communications of the ACM, 67(2):56–67, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624720.

Shanahan:2024:RTA

[1222] Murray Shanahan. Research: Talking about large language models. Communications of the ACM, 67(2):68–79, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3624724.

Kim:2024:RAH

[1223] Rae Yule Kim. Research: Anthropomorphism and human-robot interaction. Communications of the ACM, 67(2):80–85, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624716.

Patterson:2024:REE

[1224] David Patterson, Jeffrey M. Gilbert, Marco Gruteser, Efren Robles, Krishna Sekar, Yong Wei, and Tenghui Zhu. Research: Energy and emissions of machine learning on smartphones vs. the cloud. Communications of the ACM, 67(2):86–97, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624719. See corrections

[1231].

Saxena:2024:RHH

[1225] Nitin Saxena. Research highlights:

How easy is it to describe hard polynomials?: Technical perspective. Communications of the ACM, 67(2):100, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3631338.

Limaye:2024:RHS

[1226] Nutan Limaye, Srikanth Srinivasan, and Sébastien Tavenas. Research highlights: Superpolynomial lower bounds against low-depth algebraic circuits. Communications of the ACM, 67(2): 101–108, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3611094.

Buttazzo:2024:RHB

[1227] Giorgio Buttazzo. Research highlights:

Bridging AI with real-time systems: Technical perspective. Communications of the ACM, 67(2):109, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3631339.

Liu:2024:RHT

[1228] Shengzhong Liu, Shuochao Yao, Xinzhe Fu, Rohan Tabish, Simon Yu, Ayoosh Bansal, Heechul Yun, Lui Sha, and Tarek Abdelzaher. Research highlights: Taming algorithmic priority inversion in mission-critical perception pipelines. Communications of the ACM, 67(2):110–117, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3610801.

Clegg:2024:LBH

[1229] Brian Clegg. Last byte: The human touch. Communications of the ACM, 67(2):120–ff, February 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3587273.

Vardi:2024:DWD

[1230] Moshe Y. Vardi. Departments: What do computing and economics have to say to each other? Communications of the ACM, 67(3):5, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639465.

Patterson:2024:ELC

[1231] David Patterson, Jeffrey M. Gilbert, Marco Gruteser, Efren Robles, Krishna Sekar, Yong Wei, and Tenghui Zhu. Editor’s letter: Corrigendum for February 2024 article. Communications of the ACM, 67(3):6, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3643802. See [1224].

Morton:2024:CPC

[1232] Thomas Morton. Career paths in computing: Computations for this world and out of this world. Communications of the ACM, 67(3):7, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638140.

Hill:2024:BPU

[1233] Robin K. Hill and Carlos Baquero.

BLOG@CACM: Pondering the ugly underbelly, and whether images are real. Communications of the ACM, 67(3):8–10, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3639711.

Greengard:2024:NSR

[1234] Samuel Greengard. News: Service robots roll forward. Communications of the ACM, 67(3):11–13, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3637863.

Kugler:2024:NBM

[1235] Logan Kugler. News: The battle to mitigate e-waste. Communications of the ACM, 67(3):14–16, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638041.

Shein:2024:NWS

[1236] Esther Shein. News: What’s that smell? Communications of the ACM, 67(3):17–19, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3638042.

Garfinkel:2024:NMN

[1237] Simson Garfinkel and Eugene H. Spafford. News: In memoriam: Niklaus Wirth. Communications of the ACM, 67(3):20, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3641309.

Blancato:2024:OIO

[1238] Filippo Gualtiero Blancato. Opinion: Innovation is overrated: a provocation. Communications of the ACM, 67(3):22–24, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3624727.

Samuelson:2024:LSU

[1239] Pamela Samuelson. Legally speaking: U.S. Copyright Office’s questions about generative AI. Communications of the ACM, 67(3):25–28, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3637627.

Cranor:2024:PIT

[1240] Lorrie Faith Cranor, Yuvraj Agarwal, and Pardis Emami-Naeini. Privacy: Internet of Things security and privacy labels should empower consumers. Communications of the ACM, 67(3):29–31, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3637630.

Denning:2024:PIC

[1241] Peter Denning and B. Scot Rousse.

The profession of IT: Can machines be in language? Communications of the ACM, 67(3):32–35, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3637629.

Mazurczyk:2024:ODA

[1242] Wojciech Mazurczyk, Dongwon Lee, and Andreas Vlachos. Opinion: Disinformation 2.0 in the age of AI: a cybersecurity perspective. Communications of the ACM, 67(3):36–39, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624721.

Burtch:2024:OGA

[1243] Gordon Burtch, Dokyun Lee, and Zhichen Chen. Opinion: Generative AI degrades online communities. Communications of the ACM, 67(3):40–42, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624732.

Deokuliar:2024:ITD

[1244] Harsh Deokuliar, Raghvinder S. Sangwan, Yoaukim Badr, and Satish M. Srinivasan. Improving testing of deep-learning systems. Communications of the ACM, 67(3):44–48, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3633311.

Matsudaira:2024:SMG

[1245] Kate Matsudaira. Software managers’ guide to operational excellence. Communications of the ACM, 67(3):49–51, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3633289.

Ziegler:2024:RMG

[1246] Albert Ziegler, Eirini Kalliamvakou, X. Alice Li, Andrew Rice, Devon Rifkin, Shawn Simister, Ganesh Sittampalam, and Edward Aftandilian. Research: Measuring GitHub Copilot’s impact on productivity. Communications of the ACM, 67(3):54–63, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3633453.

Ahmed:2024:RIB

[1247] Saad Ahmed, Bashima Islam, Kasim Sinan Yildirim, Marco Zimmerling, Przemysław Pawełczak, Muhammad Hamad Alizai, Brandon Lucia, Luca Mottola, Jacob Sorber, and Josiah Hester. Research: The Internet of Batteryless Things. Communications of the ACM, 67(3):64–73, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/3624718.

Kimelfeld:2024:RUF

[1248] Benny Kimelfeld and Phokion G. Kolaitis. Research: A unifying framework for incompleteness, inconsistency, and uncertainty in databases. Communications of the ACM, 67(3):74–83, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624717.

Chapman:2024:RCD

[1249] Roderick Chapman, Claire Dross, Stuart Matthews, and Yannick Moy. Research: Co-developing programs and their proof of correctness. Communications of the ACM, 67(3):84–94, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624728.

Wichs:2024:RHT

[1250] Daniel Wichs. Research highlights:

Technical perspective: Hiding secrets in programs. Communications of the ACM, 67(3):96, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3632568.

Jain:2024:RHI

[1251] Aayush Jain, Huijia Lin, and Amit Sahai. Research highlights: Indistinguishability obfuscation from wellfounded assumptions. Communications of the ACM, 67(3):97–105, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3611095.

Hoffmann:2024:LBV

[1252] Leah Hoffmann. Last byte: Verifying correctness. Communications of the ACM, 67(3):108, 107, March 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639575.

Cerf:2024:CTA

[1253] Vinton G. Cerf. Cerf’s up: Thoughts on AI interoperability. Communications of the ACM, 67(4):5, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3649475.

Hazzan:2024:BUD

[1254] Orit Hazzan and Avital Binah-Pollak. BLOG@CACM: Using data science to predict how rituals will evolve. Communications of the ACM, 67(4):6–8, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3640539.

Edwards:2024:NAB

[1255] Chris Edwards. News: Answers begin to surface to Deep Earth puzzles. Communications of the ACM, 67(4):9–11, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3640478.

Shein:2024:NGS

[1256] Esther Shein. News: Governments setting limits on AI. Communications of the ACM, 67(4):12–14, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3640506.

Kirkpatrick:2024:NAA

[1257] Keith Kirkpatrick. News: Abusing AI for advertising. Communications of the ACM, 67(4):15–17, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3640507.

Sako:2024:TSM

[1258] Mari Sako. Technology strategy and management: How generative AI fits into knowledge work. Communications of the ACM, 67(4):20–22, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3638567.

Johnson:2024:OGA

[1259] Maggie Johnson. Opinion: Generative AI and CS education. Communications of the ACM, 67(4):23–24, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3632523.

Farrell:2024:OBS

[1260] Henry Farrell and Cosma Shalizi. Opinion: Bias, skew, and search engines are sufficient to explain online toxicity. Communications of the ACM, 67(4):25–28, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3624715.

Neville-Neil:2024:KVT

[1261] George V. Neville-Neil. Kode Vicious: Is there another system? Communications of the ACM, 67(4):29–30, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3640316.

Faggioli:2024:OWD

[1262] Guglielmo Faggioli, Laura Dietz, Charles L. A. Clarke, Gianluca Demartini, Matthias Hagen, Claudia Hauff, Noriko Kando, Evangelos Kanoulas, Martin Potthast, Benno Stein, and Henning Wachsmuth. Opinion: Who determines what is relevant? Humans or AI? Why not both? Communications of the ACM, 67(4):31–34, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.
1145/3624730.

Kapoor:2024:OID

[1263] Hemangee K. Kapoor and David A. Patterson. Opinion: Increasing diversity, equity, and inclusion awareness: an example from India. Communications of the ACM, 67(4):35–36, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3630747.

Cooper:2024:DOU

[1264] Geoffrey H. Cooper. Device onboarding using FDO and the untrusted installer model. Communications of the ACM, 67(4):38–45, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3637830.

Vachon:2024:SJ

[1265] Phil Vachon. The security jawbreaker. Communications of the ACM, 67(4):46–48, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3637829.

Tang:2024:RSD

[1266] Ruixiang Tang, Yu-Neng Chuang, and Xia Hu. Research: The science of detecting LLM-generated text. Communications of the ACM, 67(4):50–59, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624725.

Lee:2024:RSC

[1267] Minhyung Lee, Jinyoung Min, Junyeong Lee, Chanhee Kwak, and Hanbyeol Stella Choi. Research: A small clue made of fragmented pieces. Communications of the ACM, 67(4):60–69, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624714.

Himmel:2024:RFP

[1268] Anne-Sophie Himmel, George B.

Mertzios, André Nichterlein, and Rolf Niedermeier. Research: Fast parameterized preprocessing for polynomialtime solvable graph problems. Communications of the ACM, 67(4):70–79, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3624713.

Steinhofel:2024:RLB

[1269] Dominic Steinhöfel and Andreas Zeller.

Research: Language-based software testing. Communications of the ACM, 67(4):80–84, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3631520.

Lea:2024:RHT

[1270] Doug Lea. Research highlights: Technical perspective: Learning-based memory allocation for C++ server workloads. Communications of the ACM, 67 (4):86, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3636500.

Maas:2024:RHC

[1271] Martin Maas, David G. Andersen, Michael Isard, Mohammad Mahdi Javanmard, Kathryn S. McKinley, and Colin Raffel. Research highlights: Combining machine learning and lifetime-based resource management for memory allocation and beyond. Communications of the ACM, 67(4):87–96, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3611018.

Shasha:2024:LBJ

[1272] Dennis Shasha. Last byte: Jump snatch. Communications of the ACM, 67(4):100–ff, April 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3638565.

Vardi:2024:DMD

[1273] Moshe Y. Vardi. Departments: More is different, in social media too. Communications of the ACM, 67(5):5, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3653683.

Kapoor:2024:AEA

[1274] Hemangee Kapoor, Cherri Pancake, and Gerald Segal. ACM election: ACM’s 2024 general election. Communications of the ACM, 67(5):7–21, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3652030.

Bruderer:2024:BDD

[1275] Herbert Bruderer and Alex Vakulov.

BLOG@CACM: The dangers of digitization, and the importance of data backup. Communications of the ACM, 67(5):23–25, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3643990.

Monroe:2024:NME

[1276] Don Monroe. News: More efficient fault-tolerant quantum computing. Communications of the ACM, 67(5):26–28, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3640350.

Savage:2024:NCC

[1277] Neil Savage. News: Can ChatGPT learn Chinese or Swahili? Communications of the ACM, 67(5):29–31, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3640351.

Mone:2024:NTF

[1278] Gregory Mone. News: Thirsty fabs.

Communications of the ACM, 67(5): 32–33, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.

org/doi/10.1145/3640352.

Grimmelmann:2024:LTR

[1279] James Grimmelmann. Law and technology: The return of age verification laws. Communications of the ACM, 67(5):34–36, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3651865.

Pochat:2024:SRR

[1280] Victor Le Pochat. Security: Reflecting on research practices. Communications of the ACM, 67(5):37–39, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3651965.

Clegg:2024:ECC

[1281] Tamara L. Clegg. Education: Community is critical too. Communications of the ACM, 67(5):40–42, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3651982.

Dunbar:2024:OVL

[1282] William Dunbar. Opinion: Vendorlocked DevOps strategies. Communications of the ACM, 67(5):43–44, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3633776.

Berente:2024:OTD

[1283] Nicholas Berente, Cameron Kormylo, and Christoph Rosenkranz. Opinion: Test-driven ethics for machine learning. Communications of the ACM, 67(5):45–47, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3633455.

Strelzoff:2024:OHI

[1284] Andrew Strelzoff, Benjamin D. Trump, Christopher L. Cummings, Madison Smith, Stephanie E. Galaitsi, Kelsey Stoddard, Jeff Keisler, Moshe Vardi, Nathaniel Bastian, Alexander Kott, and Igor Linkov. Opinion: Human intuition and algorithmic efficiency must be balanced to enhance data mesh resilience. Communications of the ACM, 67(5):48–51, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3632290.

Kaushik:2024:RHS

[1285] Divyansh Kaushik, Zachary C. Lipton, and Alex John London. Resolving the human-subjects status of ML’s crowdworkers. Communications of the ACM, 67(5):52–59, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3641858.

Chisnall:2024:HDI

[1286] David Chisnall. How to design an ISA. Communications of the ACM, 67(5):60–66, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3640538.

Stork:2024:RCV

[1287] David G. Stork. Research: Computer vision, ML, and AI in the study of fine art. Communications of the ACM, 67(5):68–75, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3633454.

Schocken:2024:RNT

[1288] Shimon Schocken. Research: Nand to Tetris: Building a modern computer system from first principles. Communications of the ACM, 67(5):76–85, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626513.

Ryu:2024:RJL

[1289] Sukyoung Ryu and Jihyeok Park. Research: JavaScript language design and implementation in tandem. Communications of the ACM, 67(5):86–95, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3624723.

Freund:2024:RHI

[1290] Daniel Freund. Research highlights: Improving refugees’ integration with online resource allocation: Technical perspective. Communications of the ACM, 67(5):9898, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3639539.

Ahani:2024:RHD

[1291] Narges Ahani, Paul Gölz, Ariel D. Procaccia, Alexander Teytelboym, and Andrew C. Trapp. Research highlights: Dynamic placement in refugee resettlement. Communications of the ACM, 67(5):99–106, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3611073.

Pitre:2024:LBF

[1292] Boisy Pitre. Last byte: A fit of anger. Communications of the ACM, 67(5):108–ff, May 2024. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649306.